### v1

# Global well-posedness of 2D Navier-Stokes with Dirichlet boundary fractional noise

Abstract.
In this paper, we prove the global well-posedness and interior regularity for the 2D Navier-Stokes equations driven by a fractional noise acting as an inhomogeneous Dirichlet-type boundary condition. The model describes a vertical slice of the ocean with a relative motion between the two surfaces and can be thought of as a stochastic variant of the Couette flow. The relative motion of the surfaces is modeled by a Gaussian noise which is coloured in space and fractional in time with Hurst parameter $\mathcal{H}>\frac{3}{4}$.
Acknowledgements.
The authors thank Robert Denk and Tim Seitz for several useful discussions.
Key words and phrases:
Navier-Stokes Equations, Stochastic Boundary Conditions, Maximal Regularity, Fractional Brownian Motion, Dirichlet Boundary Conditions, Infinite Energy Solutions, Couette Flow
1991 Mathematics Subject Classification:
60H15, 60H30, 76D03 (47A60, 35J25)
[…]
1. Introduction
In many situations occurring in applied sciences, noise can affect the evolution of a system only through the boundary of a region where the system evolves. Such phenomena can be modeled via partial differential equations with boundary noise, as introduced by Da Prato and Zabczyck in the seminal paper [20]. Such a description presents several issues from a mathematical viewpoint. Indeed, nowadays it is well-known that in the one dimensional case, the solution of the heat equation with white noise Dirichlet or Neumann boundary conditions has low (space) regularity compared to the case of noise diffused inside the domain. This is due to the large amplitude of the fluctuations of the solutions close to the boundary. In particular, in the case of Dirichlet boundary conditions the solution is only a distribution. This allowed us to treat only a restricted class of nonlinearities, exploiting specific properties of the heat semigroup and studying carefully the blow-up of the solution close to the boundary. For some results in this direction, the reader is referred to [2, 12, 23, 34]. On the contrary, in the last few years, maximal $L^{p}$ regularity techniques provided new ideas to treat partial differential equations with white noise Neumann boundary conditions with more severe nonlinearities such as those coming from fluid dynamical models. Indeed, some results on the global and local well-posedness of the 2D Navier-Stokes equations and the 3D primitive equations with boundary noise perturbations of Neumann time have been proven in [1] and [10], respectively. Besides the physical interests in studying the Navier-Stokes equations with boundary noise in virtue of its connection with the Couette flow (see also below for further motivations), the present manuscript also aims at (partially) filling the gap in the literature between Dirichlet and Neumann type boundary conditions for fluid dynamical models.
Throughout the manuscript, we fix a finite time horizon $T>0$ and we let $\mathcal{O}=\mathbb{T}\times(0,a)$ where $\mathbb{T}$ is the one-dimensional torus and $a>0$. Moreover, we denote by
(1.1) $\Gamma_{b}=\mathbb{T}\times\{0\}\quad\text{ and }\quad\Gamma_{u}=\mathbb{T}\times\{a\},$
the bottom and the upper part of the boundary of $\mathcal{O}$, respectively.
In this paper we are interested in the global well-posedness and the interior regularity of the 2D Navier-Stokes equations with boundary noise for the unknown velocity field $u(t,\omega,x,z)=(u_{1},u_{2}):(0,T)\times\Omega\times\mathcal{O}\to{\mathbb{R}}^{2}$ and pressure $P:(0,T)\times\Omega\times\mathcal{O}\to{\mathbb{R}}$, formally written as
(1.2) $\partial_{t}u&=\Delta u+\nabla P-(u\cdot\nabla)u\quad&\text{ on }&(0,T)\times\mathcal{O},\\ \operatorname{div}u&=0\quad&\text{ on }&(0,T)\times\mathcal{O},\\ u_{1}&=g\,\dot{W}^{\mathcal{H}}\quad&\text{ on }&(0,T)\times\Gamma_{u},\\ u_{2}&=0\quad&\text{ on }&(0,T)\times\Gamma_{u},\\ u&=0\quad&\text{ on }&(0,T)\times\Gamma_{b},\\ u(0)&=u_{\mathrm{in}}\quad&\text{ on }&\mathcal{O},\end{aligned}\right.$
where $(u_{\mathrm{in}},g)$ are given data and $W^{\mathcal{H}}$ is a fractional Brownian motion with Hurst parameter $\mathcal{H}>\frac{3}{4}$, respectively. The assumptions on $(u_{\mathrm{in}},g,W^{\mathcal{H}})$ are made precise below. Even if we consider a more regular noise in time than the one introduced in [20], the combination of the blow-up of the solution close to the boundary and the Navier-Stokes nonlinearity makes the global well-posedness and the interior regularity of (1.2) a non-trivial issue, which, indeed, cannot be treated simply by the techniques introduced in [1]. Indeed, to the best of our knowledge, this is the first instance of a global well-posedness result for a fluid dynamical system with non-homogeneous Dirichlet-type boundary conditions of a regularity class comparable with the time derivative of a fractional Brownian motion with Hurst parameter $\mathcal{H}>\frac{3}{4}$, see [14, 26, 35] and the references therein for some results in this direction. Moreover, the reader is referred to [24] for the analysis of some properties of (1.2) in the 3D case replacing $g\,\dot{W}^{\mathcal{H}}$ with an Ornstein Uhlenbeck process and to [2, 34] for some results on the existence and uniqueness of solutions for the heat equation with white noise Dirichlet type boundary conditions perturbed by some Lipschitz forcing. Finally, in [9, 54] the emphasis is on the non-penetration boundary conditions, namely it is studied the case $u_{2}=g(x,t)$ on $\Gamma_{u}\cup\Gamma_{b}$, with $g$ much more regular either in time and space than $g\,\dot{W}^{\mathcal{H}}$.
According to [33, 50, 51], see also the discussion in the introduction of [1], the geometry considered in (1.2) can be seen as an idealization of the ocean dynamics (more precisely, a vertical slice of the ocean). The model (1.2), describes a Couette flow, namely a viscous fluid in the space between two surfaces, one of which is moving tangentially relative to the other. The relative motion of the surfaces imposes a shear stress on the fluid and induces the flow. Let us recall that the onset of turbulence is often related to the randomness of background movement [45]. Moreover, according to [52, Chapter 3] in any turbulent flow there are unavoidably perturbations in boundary conditions and material properties. We model these features by the noise term $g\,\dot{W}^{\mathcal{H}}$. As introduced by Kolmogorov in [39], fractional Brownian motion can be thought of as a model for turbulence. Moreover, to describe turbulence in 3D fluids, models of random vortex filaments have been introduced in [27]. These have been analyzed for fractional Brownian motion with $\mathcal{H}>1/2$ in [49] and $\mathcal{H}<1/2$ in [28].
1.1. Main result
We begin by introducing some notation. Throughout this manuscript, we consider a complete filtered probability space $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\geq 0},\mathbf{P})$, a separable Hilbert space $U$. We say that a process $\Phi$ is $\mathcal{F}$-progressive measurable if $\Phi|_{(0,t)\times\Omega}$ is $\mathcal{F}_{t}\times\mathcal{B}((0,t))$-measurable for all $t>0$, where $\mathcal{B}$ denotes the Borel $\sigma$-algebra. We postpone to subsection 1.3 the relevant notation on function spaces. On the noise $W^{\mathcal{H}}$ we enforce the following
Assumption 1.1.
$W^{\mathcal{H}}$ is a $U$-cylindrical fractional Brownian motion with Hurst parameter $\mathcal{H}\in(\frac{3}{4},1)$ and $g\in\mathscr{L}_{2}(U,H^{-s}(\Gamma_{u}))$ with $s\in[0,\frac{1}{2})$ and $\mathcal{H}-\frac{s}{2}>\frac{3}{4}$.
Note that 1.1 is consistent with the results obtained in [22] for the stochastic heat equation with Dirichlet fractional noise. The reader is referred to 1.4 for the case of a time-dependent $g$. Following [21, Chapter 15] and [19], we look for solutions to (1.2) in the form
$u=w_{g}+v,$
where $w_{g}$ is a mild solution of the linear problem with non-homogeneous boundary conditions
(1.3) $\partial_{t}w_{g}&=\Delta w_{g}+\nabla P_{g}\ \ &\text{on }&(0,T)\times\mathcal{O},\\ \operatorname{div}w_{g}&=0\ \ &\text{ on }&(0,T)\times\mathcal{O},\\ w_{g,1}&=g\,\dot{W}^{\mathcal{H}}\ \ &\text{ on }&(0,T)\times\Gamma_{u},\\ {w_{g,2}}&=0\ \ &\text{ on }&(0,T)\times\Gamma_{u},\\ w_{g}&=0\ \ &\text{ on }&(0,T)\times\Gamma_{b},\\ w_{g}(0)&=0\ \ &\text{ on }&\mathcal{O}\end{aligned}\right.$
and $v$ is a weak solution of
(1.4) $\partial_{t}v&=\Delta v+\nabla(P-P_{g})\\ &\quad\quad\ \ -\operatorname{div}((v+w_{g})\otimes(v+w_{g}))\ \ &\text{ on }&(0,T)\times\mathcal{O},\\ \operatorname{div}v&=0\ \ &\text{ on }&(0,T)\times\mathcal{O},\\ v&=0\ \ &\text{ on }&(0,T)\times\left(\Gamma_{b}\cup\Gamma_{u}\right),\\ v(0)&=u_{\mathrm{in}}\ \ &\text{ on }&\mathcal{O}.\end{aligned}\right.$
In (1.4), due to the divergence-free of $v$ and $w_{g}$, we rewrote the Navier-Stokes nonlinearity in the conservative form to accommodate the weak (PDE) setting.
As discussed in [21, Chapter 13], if $g$, $u_{\mathrm{in}}$, $W^{\mathcal{H}}(t)$ were regular enough, then $u=v+w_{g}$ would be a classical solution of the Navier-Stokes equations with non-homogeneous boundary conditions (1.2).
Next, we introduce the class of solutions we are going to consider. To motivate them, let us first discuss the regularity of $w_{g}$. It is well-known that, in the case of Dirichlet-type boundary conditions, the solution of a linear problem with boundary noise and $\mathcal{H}=\frac{1}{2}$ is a distribution which blows-up close to the boundary, see [2, 20], and the same holds also in case of $\mathcal{H}\neq\frac{1}{2}$, see [12]. Therefore, we cannot expect that the mild solution of (1.3) has arbitrarily good integrability properties as in [1, 10]. This has drastic consequences in our analysis. As we will show in 3.1, we have, $\mathbf{P}-a.s.$,
(1.5) $w_{g}\in C([0,T];L^{2q}(\mathcal{O};{\mathbb{R}}^{2}))\ \text{ for all }q\in(1,q_{\mathcal{H}})$
where
(1.6) $q_{\mathcal{H}}:=\frac{2}{2s+5-4\mathcal{H}}\in(1,2).$
Let us stress that $q_{\mathcal{H}}<2$ and $\lim_{\mathcal{H}\downarrow 3/4}q_{\mathcal{H}}\leq 1$ even if $s=0$. As we will see below, this fact creates major difficulties in our analysis of the auxiliary Navier-Stokes equations (1.4). In particular, $w_{g}\otimes w_{g}\in C([0,\infty);L^{q}(\mathcal{O};{\mathbb{R}}^{2}))$ $\mathbf{P}-a.s.$ and, from parabolic regularity, the best regularity we can hope for is $v\in L^{p}([0,\infty);H^{1,q}(\mathcal{O};{\mathbb{R}}^{2}))$ $\mathbf{P}-a.s.$ for all $p<\infty$. Thus, in general,
$v\not\in L^{2}(0,T;H^{1}(\mathcal{O};{\mathbb{R}}^{2}))\ \ \mathbf{P}-a.s.\text{ for any }T<\infty.$
Therefore, $v$ is a solution of the Navier-Stokes equations with infinite energy and the argument used in [1] does not work. The case of infinite energy solutions of 2D Navier-Stokes equations already appeared in the literature [11, 31]. In [31] the unboundedness of the energy is due to a rough initial data $u_{0}\not\in L^{2}$ while in [11] to a rough forcing term $f\not\in L^{2}(0,T;H^{-1})$ acting on the bulk. Our case does not fit in any of the above situations due to the presence of transport-type terms depending on the $w_{g}$ in (1.4) and the fact that we are working on domains. For this reason, our proofs rely on different methods. For details, the reader is referred to the text before 1.4.
In light of the previous discussion, we are now ready to define solutions to (1.2). Below, we set $A:B=\sum_{i,j=1}^{2}A^{i,j}B^{i,j}$ for two matrices $A$ and $B$ and $\mathbb{L}^{q}$ the image of $L^{q}(\mathcal{O};{\mathbb{R}}^{2})$ via the Leray projection $\mathbb{P}$ defined rigorously in subsection 2.1.
Definition 1.2.
Let $T<\infty$, $u_{\mathrm{in}}\in L^{0}_{\mathscr{F}_{0}}(\Omega;\mathbb{L}^{2})$ and $q\in(1,q_{\mathcal{H}})$.
  • ($q$-solution) A progressively measurable process $u$ with $\mathbf{P}-a.s.$ paths in $L^{2q^{\prime}}(0,T;\mathbb{L}^{2q})$, is a pathwise weak $q$-solution of (1.2) if $u=v+w_{g}$ and $v$ satisfies for all divergence-free $\varphi\in C^{\infty}(\mathcal{O};{\mathbb{R}}^{2})$ such that $\varphi=0$ on $\Gamma_{b}\cup\Gamma_{u}$ and a.e. $t\in(0,T)$,
    $\int_{\mathcal{O}}v(x,t)\varphi(x)\,\mathrm{d}x-\int_{\mathcal{O}}u_{\mathrm{in}}(x)\varphi(x)\,\mathrm{d}x$
    $\qquad\qquad=\int_{0}^{t}\int_{\mathcal{O}}\big{(}v\cdot\Delta\varphi+[(v+w_{g})\otimes(v+w_{g})]:\nabla\varphi\big{)}\,\mathrm{d}x\,\mathrm{d}s.$
  • (unique $q$-solution) A $q$-solution $u$ to (1.2) is said to be a unique if for any other $q$-solution $\widetilde{u}$ we have $u=\widetilde{u}$ a.e. on $[0,T]\times\Omega$.
  • (unique solution) A $q$-solution $u$ is said to be a unique solution to (1.2) if it is also a $\widetilde{q}$-solution for all $\widetilde{q}\in(1,q_{\mathcal{H}})$.
Before stating our main result, let us first comment on the above definition. Due to the argument below (1.4), one cannot expect solutions to (1.2) with integrability in space larger or equal to $2q_{\mathcal{H}}$. Furthermore, unique solution of (1.2) are independent of the choice of $q\in(1,q_{\mathcal{H}})$. Such independence is expected from solutions to (1.2) in light of (1.5). Finally, let us discuss the regularity class chosen to define $q$-solutions. Since $\mathcal{O}$ is two-dimensional, the space $L^{2q^{\prime}}(0,T;L^{2q}(\mathcal{O};{\mathbb{R}}^{2}))$ has Sobolev index given by (keeping in mind the parabolic scaling)
$-\frac{2}{2q^{\prime}}-\frac{2}{2q}=-1.$
In particular, the regularity class chosen for $q$-solutions to (1.2) is critical for the Navier-Stokes equations in two dimensions and satisfies the classical Ladyzhenskaya–Prodi–Serrin condition. In light of the recent convex integration results [15, 16, 42] in absence of noise and with periodic boundary conditions in all directions, the regularity assumption in our definition is expected to be sharp for obtaining uniqueness and a-fortiori well-posedness.
The main result of the current work reads as follows.
Theorem 1.3.
Let 1.1 be satisfied and $u_{\mathrm{in}}\in L^{0}_{\mathscr{F}_{0}}(\Omega;\mathbb{L}^{2})$.
  1. (1)
    There exists a unique solution of (1.2) in the sense of 1.2 with paths in
    $u\in C([0,T];\mathbb{L}^{2})\ \ \mathbf{P}-a.s.$
  2. (2)
    The unique solution of (1.2) satisfies, for all $t_{0}\in(0,T)$ and $\mathcal{O}_{0}\subset\mathcal{O}$ such that $\mathrm{dist}({\mathcal{O}_{0}},\partial\mathcal{O})>0$,
    $u\in C([t_{0},T];C^{\infty}({\mathcal{O}_{0}};\mathbb{R}^{2}))\ \ \mathbf{P}-a.s.$
The proof of Theorem 1.3(1) and (2) are given subsection 3.3 and subsection 4.2, respectively. Routine extensions of the above are commented in 1.4 below.
Next, let us discuss the main ideas behind the proof of Theorem 1.3. As commented above, due to (1.5), we cannot deal with the techniques introduced in [1] to study (1.4). Indeed, contrary to [1, 19], the splitting introduced above is not enough to study the global well-posedness of (1.2) since (1.4) has no Leray solutions since $w_{g}\otimes w_{g}\not\in L^{2}(0,T;L^{2})$. Thus, we control the blow-up of the energy of $v$ introducing further splittings depending on the regularity of $w_{g}$. As discussed above we will show that $w_{g}\in C([0,T];L^{2q})$ for some $q\in(1,q_{\mathcal{H}})$. Since the space $C([0,T];L^{2q})$, $q>1$ is subcritical for 2D Navier-Stokes equations we have some hopes to exploit the strong time regularity of $w_{g}$ to circumvent its rough behaviour in space. The heuristic idea above is realized by writing
(1.7) $v=\sum_{i=0}^{N-1}v_{i}+\overline{v},$
where $N$ depends only on $q$. The terms $\{v_{i}\}_{i\in\{0,\dots,N-1\}}$ are defined inductively solving homogeneous Stokes equations with forcing having mixed regularity in space and time, such that the regularity in space increases in $i$ while the regularity in time decreases in $i$. On the contrary $\overline{v}$ is a Leray-type solution of the remainder equation. In particular, $N$ is chosen large enough such that the equation for $\overline{v}$ has a forcing in $L^{2}(0,T;H^{-1})$ and therefore is regular enough to prove the existence and uniqueness of Leray solutions, see Theorem 2.5 below. The reader is referred to subsection 3.2 for further discussions.
The interior regularity of $u$ in Theorem 1.3(1) is treated considering again the splitting $u=v+w_{g}$ introduced above. The interior regularity of $w_{g}$ can be proved similarly to the linear part of [1]. On the contrary, the low regularity of $v$ does not allow us to study directly its interior regularity by Serrin’s argument as in [1]. For this reason, we rely on the splitting (1.7) analysed in subsection 3.2 to study the well-posedness of (1.4). Combining maximal $L^{p}$ regularity techniques for studying the interior regularity of the $v_{i}$’s, an induction argument and a Serrin argument for treating the interior regularity of $\overline{v}$, we obtain the required regularity of $v$. According to [58] (see also [41, Section 13.1]), it seems not possible to gain higher-order interior time-regularity for the Navier-Stokes problem. This fact is in contrast to the case of the heat equation with white noise boundary conditions, see [12]. The reason behind this is the presence of the unknown pressure $P$ which, due to its non-local nature, provides a connection between the interior and the boundary regularity where the noise acts.
To conclude, let us note that, in contrast to [11, 31], we employ a different splitting scheme to prove existence due to the presence of the transport-type terms originated by $w_{g}$. Moreover, the number of splitting $N$ depends on how much the Sobolev index of the space $C([0,T];L^{2q}(\mathcal{O};{\mathbb{R}}^{2}))$, i.e. $-\frac{1}{q}$, is far from the critical threshold $-1$. In particular, $N\to\infty$ as $q\downarrow 1$. As commented above, such a splitting is also convenient when proving the interior regularity for $u$ which was not addressed in the above-mentioned works.
Remark 1.4 (Extensions).
One can readily check that Theorem 1.3 extends in the following cases:
  • (The case of bounded domains) If $\mathcal{O}$ is replaced by a smooth $C^{2}$-bounded domain in ${\mathbb{R}}^{2}$. However, we prefer to keep the same geometry of [1] for two reasons. Firstly, and more importantly, as discussed in section 1, the model considered has a clear physical interpretation. Secondly, in this way, we can easily compare our results, techniques and assumptions with those of [1].
  • (Fractional Volterra noise) If $W^{\mathcal{H}}$ is replaced by a $\alpha$-regular Volterra process with $\alpha>\frac{1}{4}$. Let us recall that a fractional Brownian motion with Hurst parameter $\mathcal{H}$ is an example of a $\alpha$-regular Volterra process with $\alpha=\mathcal{H}-\frac{1}{2}$. These are non-Markovian stochastic processes which can be represented as integrals of kernels with respect to the Brownian motion and include for example the fractional Liouville Brownian motion and the Rosenblatt process. Stochastic convolutions with respect to such processes were analyzed in [13, 17, 18].
  • (Time-dependent $g$) The term $g$ in the boundary noise $g\dot{W}^{\mathcal{H}}$ depends on time as long as it is progressively measurable and the corresponding process $w_{g}$ satisfies (1.5) holds.
1.2. Overview
The paper is organized as follows. In section 2 we introduce the functional setting to deal with problem (1.2). The proof of Theorem 1.3 is the object of section 3 and section 4. In particular, the global well-posedness, i.e. item 1, is addressed in section 3 considering the linear problem (1.3) in subsection 3.1 and the nonlinear one (1.4) in subsection 3.2. The interior regularity, i.e. item 2, is the object of section 4. In particular, in subsection 4.1 we study the interior regularity of the solution of the linear problem (1.3), while in subsection 4.2 we consider the nonlinear problem (1.4).
1.3. Notation
Here we collect some notation which will be used throughout the paper. Further notation will be introduced where needed. By $C$ we will denote several constants, perhaps changing value line by line. If we want to keep track of the dependence of $C$ from some parameter $\xi$ we will use the symbol $C(\xi)$. Sometimes we will use the notation $a\lesssim b$ (resp. $a\lesssim_{\xi}b$) if it exists a constant such that $a\leq Cb$ (resp. $a\leq C(\xi)b$).
Fix $q\in(1,\infty)$. For an integer $k\geq 1$, $W^{k,q}$ denotes the usual Sobolev spaces. In the non-integer case $s\in(0,\infty)\setminus{\mathbb{N}}$, we let $W^{s,q}=B^{s}_{q,q}$ where $B^{s}_{q,q}$ is the Besov space with smoothness $s$, and integrability $q$ and microscopic integrability $q$. Moreover, $H^{s,q}$ denotes the Bessel potential spaces. Both Besov and Bessel potential spaces can be defined using Littlewood-Paley decompositions and restrictions (see e.g. [56], [55, Section 6]) or using the interpolation methods starting with the standard Sobolev spaces $W^{k,q}$ (see e.g. [8, Chapter 6]). Finally, we set $\mathcal{A}^{s,q}(D;{\mathbb{R}}^{d})=(\mathcal{A}^{s,q}(D))^{d}$ for an integer $d\geq 1$, a domain $D$ and $\mathcal{A}\in\{W,H\}$. Let $\mathcal{K}_{1}$ and $\mathcal{K}_{2}$ be two Hilbert spaces. We denote by $\mathscr{L}_{2}(\mathcal{K}_{1},\mathcal{K}_{2})$ the set of Hilbert-Schmidt operators from $\mathcal{K}_{1}$ to $\mathcal{K}_{2}$. Below, we need the following Fubini-type result:
$H^{s}(D;\mathcal{K}_{1})=\mathscr{L}_{2}(\mathcal{K}_{1},H^{s}(D))\ \ \text{ for all }s\in{\mathbb{R}}.$
The above follows from [37, Theorem 9.3.6] and interpolation.
2. Preliminaries
2.1. The Stokes operator and its spectral properties
In this section, we introduce the functional analytic setup to define all the objects necessary in the following. Throughout this subsection, we let $r\in(1,\infty)$. Recall that $\mathcal{O}=\mathbb{T}\times(0,a)$ where $a>0$.
We begin by introducing the Helmholtz projection on $L^{r}(\mathcal{O};{\mathbb{R}}^{2})$, see e.g. [53, Subsection 7.4]. Let $f\in L^{r}(\mathcal{O};{\mathbb{R}}^{2})$ and let $\psi_{f}\in W^{1,r}(\mathcal{O})$ be the unique solution of the following elliptic problem
(2.1) $\Delta\psi_{f}&=\operatorname{div}f\quad&\text{ on }&\mathcal{O},\\ \partial_{n}\psi_{f}&=f\cdot\widehat{n}&\text{ on }&\Gamma_{u}\cup\Gamma_{b}.\end{aligned}\right.$
Here $n$ denotes the exterior normal vector field on $\partial\mathcal{O}$. Of course, the above elliptic problem is interpreted in its naturally weak formulation:
(2.2) $\int_{\mathcal{O}}\nabla\psi_{f}\cdot\nabla\varphi\,\mathrm{d}x\mathrm{d}z=\int_{\mathcal{O}}f\cdot\nabla\varphi\,\mathrm{d}x\mathrm{d}z\ \ \text{ for all }\varphi\in C^{\infty}(\mathcal{O}).$
By [53, Corollary 7.4.4] , we have $\psi_{f}\in W^{1,r}(\mathcal{O})$ and $\|\nabla\psi_{f}\|_{L^{r}(\mathcal{O};{\mathbb{R}}^{2})}\lesssim\|f\|_{L^{r}(\mathcal{O};{\mathbb{R}}^{2})}$. Then the Helmholtz projection is given by $\mathbb{P}_{r}$ is defined as
$\mathbb{P}_{r}f=f-\nabla\psi_{f},\quad f\in L^{r}(\mathcal{O};{\mathbb{R}}^{2}).$
Next, we define the Stokes operator on $L^{r}(\mathcal{O};{\mathbb{R}}^{2})$. For notational convenience, we define $A_{r}$ as minus the Stokes operator so that $A_{r}$ is a positive operator for $r=2$ (i.e. $\langle A_{2}u,u\rangle\geq 0$ for all $u\in\mathsf{D}(A_{2})$). Let
$\mathbb{L}^{r}:=\mathbb{P}(L^{r}(\mathcal{O};{\mathbb{R}}^{2})),\quad\mathbb{H}^{s,r}:=H^{s,r}(\mathcal{O};{\mathbb{R}}^{2})\cap\mathbb{L}^{r},\ \ s\in{\mathbb{R}}.$
Then, we define the operator $A_{r}:\mathsf{D}(A_{r})\subseteq\mathbb{L}^{r}\to\mathbb{L}^{r}$ where
$\mathsf{D}(A_{r})=\big{\{}f=(f_{1},f_{2})\in W^{2,r}(\mathcal{O};\mathbb{R}^{2})\cap\mathbb{L}^{r}\,:\,$ $f|_{\Gamma_{b}\cup\Gamma_{u}}=0\big{\}},$
and $A_{r}u=-\mathbb{P}\Delta u$ for $u\in\mathsf{D}(A_{r})$.
In the main arguments, we need stochastic and deterministic maximal $L^{r}$-regularity estimates for convolutions. By [46, 40], it is enough to provide the boundedness of the $H^{\infty}$-calculus for $A_{r}$. The reader is referred to [53, Chapters 3 and 4] and [37, Chapter 10] for the main notation and basic results on the $H^{\infty}$-calculus.
Contrary to [1], the boundary conditions we are interested in here are much more classical. Indeed, the Stokes operator with no-slip boundary conditions is well-studied. The reader is referred e.g. to [36, Section 2.8], [47], [32] and [38, Section 9] for the proof of this nowadays classical statement.
Lemma 2.1.
For all $r\in(1,\infty)$, the operator $A_{r}$ is invertible and has a bounded $H^{\infty}$-calculus of angle 00. Moreover, the domain of the fractional powers of $A_{r}$ is characterized as follows:
$\mathsf{D}(A_{r}^{\alpha})=\left\{\begin{aligned} &\mathbb{H}^{2\alpha,r}\ \ &\text{ if }&\ \alpha<\frac{1}{2r},\\ &\big{\{}u\in\mathbb{H}^{2\alpha,r}\,:\,u|_{\partial\mathcal{O}}=0\big{\}}\ \ &\text{ if }&\ \frac{1}{2r}<\alpha\leq 1.\end{aligned}\right.$
The above implies that $-A_{r}$ generates an analytic semigroup on $\mathbb{L}^{r}$ which admits stochastic and deterministic maximal $L^{p}$-regularity for all $p\in(1,+\infty)$, see [53, Chapter 3-4] and [46]. We denote such semigroup by $S_{r}(t)$. We continue introducing some known facts about the “Sobolev tower” of spaces associated with the operator $A_{r}$. We denote by
$X_{\alpha,A_{r}}$ $=\mathsf{D}(A_{r}^{\alpha})$ for $\alpha\geq 0,$
$X_{\alpha,A_{r}}$ $=(\mathbb{L}^{r},\|A_{r}^{{\alpha}}\cdot\|_{\mathbb{L}^{r}})^{\sim}$ for $\alpha<0,$
where $\sim$ denotes the completion. Indeed, since $0\in\rho(A_{r})$ by 2.1, we have that $f\mapsto\|A_{r}^{{\alpha}}f\|_{\mathbb{L}^{r}}$ is a norm for all $\alpha<0$. Since $(A_{r})^{*}=A_{r^{\prime}}$, it follows that (see e.g. [4, Chapter 5, Theorem 1.4.9])
(2.3) $(X_{\alpha,A_{r}})^{*}=X_{-\alpha,A_{r^{\prime}}}.$
For convenience of notation, we simply write $A,\ S(t)$ in place of $A_{2}$ and $S(t)$. Moreover we define
$H:=\mathbb{L}^{2},\quad V:=\mathsf{D}(A^{1/2}).$
[…]
2.2. The Dirichlet map
Now we are interested in $L^{2}$-estimates for the Dirichlet map, i.e. we are interested in studying the weak solutions of the elliptic problem
(2.4) $-\Delta u+\nabla\pi&=0,&\text{ on }&\mathcal{O},\\ \operatorname{div}u&=0\qquad&\text{ on }&\mathcal{O},\\ u(\cdot,0)&=0,&\text{ on }&\Gamma_{b},\\ u_{1}(\cdot,a)&=g,&\text{ on }&\Gamma_{u},\\ u_{2}&=0,&\text{ on }&\Gamma_{u}.\end{aligned}\right.$
To state the main result of this subsection, we formulate (2.4) in the very weak setting. To this end, we argue formally. Take $\varphi=(\varphi_{1},\varphi_{2})\in C^{\infty}(\mathcal{O};{\mathbb{R}}^{2})$ such that $\operatorname{div}\varphi=0$,
$\varphi=0,\quad\text{ on }\quad\Gamma_{b}\cup\Gamma_{u}.$
A formal integration by parts shows that (2.4) implies
(2.5) $\int_{\mathcal{O}}u\cdot\Delta\varphi\,\mathrm{d}x\mathrm{d}z=\int_{\mathbb{T}}g(x)\partial_{2}\varphi_{1}(x,a)\,\mathrm{d}x.$
In particular, the RHS of (2.5) makes sense even in case $g$ is a distribution if we interpret $\int_{\mathbb{T}}g(x)\partial_{2}\varphi_{1}(x,a)\,\mathrm{d}x=\langle\partial_{2}\varphi_{1}(\cdot,a),g\rangle$. The well-posedness of (2.4) is, as for the properties of the Stokes operator, a well-known fact. Indeed, Theorem 2.2 below holds. The reader is referred to [6, 7, 25, 30, 57] for its proof and more general results on the Dirichlet boundary values problem above even in case of weighted $L^{r}$ spaces of Muckenhoupt class and $u\cdot\widehat{n}|_{\Gamma_{u}\cup\Gamma_{b}}\neq 0$.
Theorem 2.2.
For all $g\in H^{-\frac{1}{2}}(\Gamma_{u})$ there exists a unique $(u,\pi)\in H\times H^{-1}(\mathcal{O})/\mathbb{R}$ very weak solution of (2.4). Moreover $(u,\pi)$ satisfy
(2.6) $\lVert u\rVert+\lVert\pi\rVert_{H^{-1}(\mathcal{O})/\mathbb{R}}\leq C\lVert g\rVert_{H^{-\frac{1}{2}}(\Gamma_{u})}.$
Finally, if $g\in H^{\frac{3}{2}}(\Gamma_{u})$, then $(u,\pi)\in\mathbb{H}^{2}\times H^{1}(\mathcal{O})/\mathbb{R}$ and
(2.7) $\lVert u\rVert_{\mathbb{H}^{2}(\mathcal{O};{\mathbb{R}}^{2})}+\lVert\pi\rVert_{H^{1}(\mathcal{O})/\mathbb{R}}\leq C\lVert g\rVert_{H^{3/2}(\Gamma_{u})}.$
Next, we denote by $\mathcal{D}$ the solution map defined by Theorem 2.2 which associate to a boundary datum $g$ the velocity $u$ solution of (2.4), i.e. $\mathcal{D}g:=u$. From the above result, we obtain
Corollary 2.3.
Let $\mathcal{D}$ and $U$ be the Dirichlet map and a separable Hilbert space, respectively. Then
$\mathcal{D}\in\mathscr{L}(H^{-\alpha}(\Gamma_{u};U),\mathscr{L}_{2}(U,\mathsf{D}(A^{-\frac{\alpha}{2}+\frac{1}{4}})))\quad\text{for }\alpha\in\left[-\frac{1}{2},0\right).$
Proof.
To begin, recall that $H^{s}(\Gamma_{u};U)=\mathscr{L}_{2}(U,H^{s}(\Gamma_{u}))$ for all $s\in{\mathbb{R}}$, see subsection 1.3. Hence, due to the ideal property of Hilbert-Schmidt operators, it is enough to consider the scalar case $U={\mathbb{R}}$.
By complex interpolation, the estimates in Theorem 2.2 yield
$\mathcal{D}:H^{2\theta-\frac{1}{2}}(\Gamma_{u})\to\mathbb{H}^{2\theta}(\mathcal{O})\ \ \text{ for all }\theta\in(0,1).$
Hence, the claim now follows from the description of the fractional power of $A$ in 2.1. ∎
2.3. Deterministic Navier-Stokes equations
Let us consider the deterministic Navier-Stokes equations with homogeneous boundary conditions
(2.8) $\partial_{t}\overline{u}+\overline{u}\cdot\nabla\overline{u}+\nabla\overline{\pi}&=\Delta\overline{u}+\overline{f},\quad&\text{ on }&(0,T)\times\mathcal{O},\\ \operatorname{div}\overline{u}&=0,&\text{ on }&(0,T)\times\mathcal{O},\\ \overline{u}&=0,&\text{ on }&(0,T)\times(\Gamma_{b}\cup\Gamma_{u}),\\ \overline{u}(0)&=\overline{u}_{0},&\text{ on }&\mathcal{O}.\end{aligned}\right.$
Define the trilinear form
(2.9) $b\left(u,v,w\right)=\sum_{i,j=1}^{2}\int_{\mathcal{O}}u_{i}\partial_{i}v_{j}w_{j}\,\mathrm{d}x\mathrm{d}z=\int_{\mathcal{O}}\left(u\cdot\nabla v\right)\cdot w\,\mathrm{d}x\mathrm{d}z$
which is well–defined and continuous on $\mathbb{L}^{p}\times\mathbb{H}^{1,q}\times\mathbb{L}^{r}$ by Hölder’s inequality, whenever
$\frac{1}{p}+\frac{1}{q}+\frac{1}{r}=1.$
Finally, we introduce the operator
$B:\mathbb{L}^{p}\times\mathbb{L}^{r}\rightarrow X_{-1/2,A_{q^{\prime}}}$
defined by the identity
$\left\langle B\left(u,v\right),\phi\right\rangle_{X_{-1/2,A_{q^{\prime}}},X_{1/2,A_{q}}}=-b\left(u,\phi,v\right)=-\int_{\mathcal{O}}\left(u\cdot\nabla\phi\right)\cdot v\,\mathrm{d}x\mathrm{d}z$
for all $\phi\in X_{1/2,A_{q}}$. Moreover, when $u\cdot\nabla v\in L^{r}\left(\mathcal{O};\mathbb{R}^{2}\right)$, it is explicitly given by
$B\left(u,v\right)=\mathbb{P}(u\cdot\nabla v).$
We have to define our notion of a weak solution for problem (2.8).
Definition 2.4.
Given $\overline{u}_{0}\in H$ and $\overline{f}\in L^{2}\left(0,T;V^{*}\right)$, we say that
$\overline{u}\in C\left(\left[0,T\right];H\right)\cap L^{2}\left(0,T;V\right)$
is a weak solution of equation (2.8) if for all $\phi\in\mathsf{D}\left(A\right)$ and $t\in[0,T]$,
$\left\langle\overline{u}\left(t\right),\phi\right\rangle-\int_{0}^{t}b\left(\overline{u}\left(s\right),\phi,\overline{u}\left(s\right)\right)\,\mathrm{d}s$
$=\left\langle\overline{u}_{0},\phi\right\rangle-\int_{0}^{t}\left\langle\overline{u}\left(s\right),A\phi\right\rangle\,\mathrm{d}s+\int_{0}^{t}\left\langle\overline{f}\left(s\right),\phi\right\rangle_{V^{*},V}\,\mathrm{d}s.$
The well-posedness of (2.8) in the sense of 2.4 is a well-known fact. Indeed the following theorem holds, see for instance [43, 59, 60].
Theorem 2.5.
For every $\overline{u}_{0}\in H$ and $\overline{f}\in L^{2}\left(0,T;V^{*}\right)$ there exists a unique weak solution of equation (2.8). It satisfies
$\lVert\overline{u}\left(t\right)\rVert^{2}+2\int_{0}^{t}\lVert\nabla\overline{u}\left(s\right)\rVert_{L^{2}}^{2}\,\mathrm{d}s=\lVert\overline{u}_{0}\rVert^{2}+2\int_{0}^{t}\left\langle\overline{u}\left(s\right),\overline{f}\left(s\right)\right\rangle_{V^{*},V}\,\mathrm{d}s.$
If $\left(\overline{u}_{0}^{n}\right)_{n\in\mathbb{N}}$ is a sequence in $H$ converging to $\overline{u}_{0}\in H$ and $\left(\overline{f}^{n}\right)_{n\in\mathbb{N}}$ is a sequence in $L^{2}\left(0,T;V^{*}\right)$ converging to $\overline{f}\in L^{2}\left(0,T;V^{*}\right)$, then the corresponding unique solutions $\left(\overline{u}^{n}\right)_{n\in\mathbb{N}}$ converge to the corresponding solution $\overline{u}$ in $C\left(\left[0,T\right];H\right)$ and in $L^{2}\left(0,T;V\right)$.
We end this section with the following lemma, which is the analogous of [29, Lemma 1.14] to the $L^{p}_{t}L^{q}_{x}$ framework.
Lemma 2.6.
If $p\geq\frac{2q}{q-2}$, $u\in C([0,T];H)\cap L^{2}(0,T;V),\ v\in L^{p}(0,T;\mathbb{L}^{q})$, then
(2.10) $B(u,v)\in L^{2}(0,T;V^{*}),$
(2.11) $B(v,u)\in L^{2}(0,T;V^{*}).$
In particular for each $t\in[0,T],\ \varepsilon,\ \varepsilon^{\prime}>0$ and $\phi\in L^{2}(0,T;V)$ it holds
$\int_{0}^{t}\lvert\langle u(s)\cdot$ $\nabla\phi(s),v(s)\rangle\rvert\,\mathrm{d}s$
(2.12) $\leq\varepsilon\lVert\phi\rVert_{L^{2}(0,t;V)}^{2}+\varepsilon^{\prime}\int_{0}^{t}\lVert u(s)\rVert_{V}^{2}\,\mathrm{d}s+\frac{C}{\varepsilon^{\frac{q}{q-2}}{\varepsilon^{\prime}}^{\frac{2}{q-2}}}\int_{0}^{t}\lVert u(s)\rVert^{2}\lVert v(s)\rVert_{\mathbb{L}^{q}}^{\frac{2q}{q-2}}\,\mathrm{d}s,$
$\int_{0}^{t}\lvert\langle v(s)\cdot$ $\nabla\phi(s),u(s)\rangle\rvert\,\mathrm{d}s$
(2.13) $\leq\varepsilon\lVert\phi\rVert_{L^{2}(0,t;V)}^{2}+\varepsilon^{\prime}\int_{0}^{t}\lVert u(s)\rVert_{V}^{2}\,\mathrm{d}s+\frac{C}{\varepsilon^{\frac{q}{q-2}}{\varepsilon^{\prime}}^{\frac{2}{q-2}}}\int_{0}^{t}\lVert u(s)\rVert^{2}\lVert v(s)\rVert_{\mathbb{L}^{q}}^{\frac{2q}{q-2}}\,\mathrm{d}s,$
where $C$ is a constant independent from $\varepsilon,\ \varepsilon^{\prime}.$
Proof.
By Hölder inequality, Sobolev embedding theorem and interpolation, for each $\phi\in V$ we have
$\lvert\langle B(u(s),v(s)),\phi\rangle\rvert$ $=\lvert\langle u(s)\cdot\nabla\phi,v(s)\rangle\rvert$
$\leq\lVert\phi\rVert_{V}\lVert v(s)\rVert_{\mathbb{L}^{q}}\lVert u(s)\rVert_{\mathbb{L}^{2q/(q-2)}}$
$\lesssim_{q}\lVert\phi\rVert_{V}\lVert v(s)\rVert_{\mathbb{L}^{q}}\lVert u(s)\rVert_{\mathsf{D}(A^{1/q})}$
$\leq\lVert\phi\rVert_{V}\lVert v(s)\rVert_{\mathbb{L}^{q}}\lVert u(s)\rVert^{1-\frac{2}{q}}\lVert u(s)\rVert_{V}^{\frac{2}{q}}.$
Therefore for each
$\int_{0}^{t}\lvert\langle u(s)\cdot$ $\nabla\phi(s),v(s)\rangle\rvert\,\mathrm{d}s\leq\varepsilon\lVert\phi\rVert_{L^{2}(0,t;V)}^{2}+\frac{C}{\varepsilon}\int_{0}^{t}\lVert v(s)\rVert^{2}_{\mathbb{L}^{q}}\lVert u(s)\rVert^{2(1-\frac{2}{q})}\lVert u(s)\rVert_{V}^{\frac{4}{q}}\,\mathrm{d}s$
$\leq\varepsilon\lVert\phi\rVert_{L^{2}(0,t;V)}^{2}+\frac{C}{\varepsilon}\left(\int_{0}^{t}\lVert u(s)\rVert_{V}^{2}\,\mathrm{d}s\right)^{2/q}\left(\int_{0}^{t}\lVert u(s)\rVert^{2}\lVert v(s)\rVert_{\mathbb{L}^{q}}^{\frac{2q}{q-2}}\,\mathrm{d}s\right)^{\frac{q-2}{q}}$
$\leq\varepsilon\lVert\phi\rVert_{L^{2}(0,t;V)}^{2}+\varepsilon^{\prime}\lVert u\rVert_{L^{2}(0,t,V)}^{2}+\frac{C}{\varepsilon^{\frac{q}{q-2}}{\varepsilon^{\prime}}^{\frac{2}{q-2}}}\int_{0}^{t}\lVert u(s)\rVert^{2}\lVert v(s)\rVert_{\mathbb{L}^{q}}^{\frac{2q}{q-2}}\,\mathrm{d}s.$
The relation above implies (2.10) and (2.6). The proof of (2.11) and (2.6) is analogous and we omit the details. ∎
2.4. Stochastic convolutions with fractional noise
Definition 2.7.
Let $U$ be a separable Hilbert space. A $U$-cylindrical fractional Brownian motion $(W^{\mathcal{H}}(t))_{t\geq 0}$ with Hurst index $H\in(0,1)$ is defined by the formal series
$W^{\mathcal{H}}(t)=\sum\limits_{n=1}^{\infty}b^{\mathcal{H}}_{n}(t)e_{n},$
where $\{e_{n}\}$ is an orthonormal basis in $U$ and $(b^{\mathcal{H}}(t))_{n\in\mathbb{N}}$ is a sequence of independent standard one-dimensional fractional Brownian motions, i.e. $\mathbf{E}[b^{\mathcal{H}}_{n}(t)]=0$ and
$\mathbf{E}[b^{\mathcal{H}}_{n}(t)b^{\mathcal{H}}_{n}(s)]=\frac{1}{2}(t^{2\mathcal{H}}+s^{2\mathcal{H}}-|t-s|^{2\mathcal{H}}),~{}~{}s,t\geq 0.$
For $\mathcal{H}=1/2$ one obtains a cylindrical Brownian motion. However for $\mathcal{H}\neq 1/2$ the fbm exhibits a totally different behaviour, in particular is neither Markov nor a semimartingale.
For our aims in 3.1, we need the following results on the regularity of stochastic convolutions established in [22, Corollary 3.1].
Lemma 2.8.
([22, Corollary 3.1]) Let $A$ be the generator of an analytic $C_{0}$-semigroup $(S(t))_{t\geq 0}$ on a separable Hilbert space $U_{1}$, $\Phi\in\mathcal{L}(U,U_{1})$. Assume that
(2.14) $\|S(t)\Phi\|_{\mathcal{L}_{2}(U_{1},U)}\leq t^{-\gamma}\text{ for}\gamma<\mathcal{H}.$
Then the stochastic convolution $\int_{0}^{t}S(t-s)\Phi~{}\mathrm{d}W^{\mathcal{H}}(s)$ has $\mathbf{P}$-a.s. $\gamma_{1}$-Hölder continuous trajectories in $\mathsf{D}(A^{\gamma_{2}})$, for $0\leq\gamma_{1}+\gamma_{2}<\mathcal{H}$. If $\Phi\in\mathcal{L}_{2}(U_{1},U)$, then the assumption (2.14) is satisfied for $\gamma=0$.
3. Global well-posedness
Here we prove Theorem 1.3(1). This section is organized as follows. Firstly, in subsection 4.1 we first prove that the solution $w_{g}$ of the 2D Stokes equations with boundary noise (1.3) satisfies (1.5). Secondly, in subsection 3.2, we prove the existence of a $q$-solution to (1.2) by studying the auxiliary Navier-Stokes problem (1.4) for a given forcing term $w=w_{g}$ satisfying the regularity assumption as in (1.5) for a given $q$. Finally, in subsection 3.3, we prove the uniqueness of solutions to (1.2) therefore concluding the proof of Theorem 1.3(1). Recall that ($q$-)solutions of (1.2) are defined in 1.2.
3.1. Stokes equations
As discussed in subsection 1.1, we start by considering the linear problem (1.3). According to [20] and [21, Chapter 15], the mild solution $w_{g}$ of the former problem is formally given by
(3.1) $w_{g}(t)=A\int_{0}^{t}S(t-s)\mathcal{D}[g]\,\mathrm{d}W^{\mathcal{H}}(s).$
Here $A$ is (minus) the Stokes operator with homogeneous boundary conditions (see Subsection 2.1).
Next, we prove that $w_{g}$ is well-defined in sufficiently regular function spaces therefore allowing us to treat the nonlinearity in the Navier-Stokes equations.
Proposition 3.1.
Let 1.1 be satisfied. Then the process $w_{g}$ is well-defined, progressively measurable, and for all $T>0$ and $\varepsilon>0$,
(3.2) $w_{g}\in L^{p}(\Omega;C([0,T];\mathsf{D}(A^{\mathcal{H}-\frac{3}{4}-\frac{s}{2}-\varepsilon})))\ \text{ for all }p\in(1,\infty).$
In particular, for all $r\in(2,2q_{\mathcal{H}})$
[…]
Proof.
Note that, thanks to 2.3
$\mathcal{D}g\in\mathcal{L}_{2}(U,\mathsf{D}(A^{\frac{1}{4}-\frac{s+\varepsilon}{2}})).$
Hence, by 2.8, a.s.,
$w_{g}=\underbrace{A^{\frac{3}{4}+\frac{s+\varepsilon}{2}}\underbrace{\int_{0}^{\cdot}S(\cdot-s)A^{\frac{1}{4}-\frac{s+\varepsilon}{2}}\mathcal{D}g\,\mathrm{d}W^{\mathcal{H}}(s)}_{\in\,C([0,T];\mathsf{D}(A^{\mathcal{H}-\frac{\varepsilon}{2}}))}}_{\in\,C([0,T];\mathsf{D}(A^{\mathcal{H}-\frac{3}{4}-\frac{s}{2}-\varepsilon}))}.$
The arbitrariness of $\varepsilon>0$ yields (3.2). To prove (3.3), note that, by 2.1,
$\mathsf{D}(A^{\mathcal{H}-\frac{3}{4}-\frac{s}{2}-\varepsilon})\subset H^{2\mathcal{H}-\frac{3}{2}-s-2\varepsilon}(\mathcal{O};{\mathbb{R}}^{2}).$
The above space embeds into $L^{r}(\mathcal{O};{\mathbb{R}}^{2})$ for some $r>2$ provided
$02\mathcal{H}-\frac{3}{2}-s-2\varepsilon>$ 0 .
The above is exactly our assumption due to the arbitrariness of $\varepsilon>0$. In particular, by the arbitrariness of $\varepsilon$ and Sobolev’s embedding we can choose whatever $r<2q_{\mathcal{H}}$.
Remark 3.2 (Necessity of the $L^{p}$-setting for $v$).
In the setting of 3.1, we have $2\mathcal{H}-\frac{3}{2}-s<\frac{1}{2}.$ Therefore, for all choices of $\mathcal{H}$ and $s$ in 1.1, it follows that
$H^{2\mathcal{H}-\frac{3}{2}-s-2\varepsilon}(\mathcal{O};{\mathbb{R}}^{2})\not\hookrightarrow L^{4}(\mathcal{O};{\mathbb{R}}^{2}).$
Thus, (3.3) holds with $r<4$ and therefore $B(w_{g},w_{g})\not\in L^{2}(0,T;V^{*}).$ In particular, in the next subsection, we cannot avoid the use of $L^{p}$-setting in space, cf., the comments below 3.4.
Remark 3.3.
Previous results with white noise boundary conditions [10], [1] exploited stochastical maximal $L^{p}$ regularity techniques to study the linear part of the problem. Here is worth mentioning that we employed the more standard Hilbert value framework because it produces the sharpest result on the regularity of the stochastic convolution in terms of the Hurst parameter $\mathcal{H}$. Indeed, assuming just for simplicity the case $s=0$ and $g\in L^{p}(\Gamma_{u};U)$ for some $p\in[2,+\infty)$, then by 2.3, [17, Proposition 4.5] and arguing as above we have
$w_{g}\in C([0,T];\mathsf{D}(A_{p}^{\mathcal{H}-1+\frac{1}{2p}-\varepsilon})).$
In particular $w_{g}\in C([0,T];\mathbb{L}^{r})$ for some $r>2$ if $\mathcal{H}>1-\frac{1}{2p}$. Therefore the right-hand side is minimized and we can use the rougher noise for $p=2$.
3.2. Auxiliary Navier-Stokes type equations
Motivated by the auxiliary problem (1.4) and by the results of the previous subsection, here we study the well-posedness of the following abstract PDE
[…]
with $A_{q}$ is the Stokes operator on $\mathbb{L}^{q}$ and $B$ is the bilinear nonlinearity as in Subsection 2.3. Finally, $q=r/2$ and $(w,r)$ satisfies the following
Assumption 3.4.
$w\in C([0,T];\mathbb{L}^{r})$ for some $r\in(2,4)$.
Note that the above assumption is satisfied $\mathbf{P}-a.s.$ with $w=w_{g}$, as it follows from 3.1. Moreover, the limitation $r<4$ is motivated by 3.2. In particular, the arguments used in [1] do not apply to (3.4). Indeed, if 3.4 holds, then
$B(w,w)\not\in L^{2}(0,T;H^{-1}(\mathcal{O};{\mathbb{R}}^{2})).$
Hence, the (potential) energy of solutions for (3.4), i.e.
$\int_{0}^{t}\int_{\mathcal{O}}|\nabla v|^{2}\,\mathrm{d}x\,\mathrm{d}s\ \ \text{ for }\ t>0,$
is ill-defined even in absence of the terms $B(v,v),B(w,v)$ and $B(v,w)$. In particular, one cannot expect energy (or Leray’s) type solutions for (3.4) to be defined and the analysis carried on in [1] does not work in our framework. However, $B(v,v)\in L^{\infty}(0,T;H^{-1,q}(\mathcal{O};{\mathbb{R}}^{2}))$ for some $q>1$ as $r>2$, and therefore $L^{q}$-theory for (3.4) can be built.
Next, let us describe the main idea behind our construction of a $r/2$-solution to (3.4). In what follows, the subcriticality of $L^{\infty}(0,T;\mathbb{L}^{r})$ with $r>1$ for the 2D Navier-Stokes equations (cf., the discussion below 1.2) plays a central role. Indeed, by subcriticality, given $q_{0}=\frac{r}{2}$, the solution $v_{0}$ to
[…]
satisfies $v_{0}\in L^{p}(0,T;L^{r_{0}}(\mathcal{O};{\mathbb{R}}^{2}))$ for some $r_{0}>r$ and each $p<+\infty$. Hence, we obtained a small gain of space regularity. In particular, $\overline{v}_{1}=v-v_{0}$ solves
[…]
In the above, we would like to take $q_{1}>q_{0}$ due to the increased regularity of the forcing terms. Indeed, as $v_{0}\in L^{p}(0,T;L^{r_{0}}(\mathcal{O};{\mathbb{R}}^{2}))$ for some $r_{0}>r$ and each $p<+\infty$ one obtains that the terms $B(w+v_{0},v_{0})$ and $B(w,v_{0})$ belong to $L^{p}(0,T;H^{-1,q_{1}}(\mathcal{O};{\mathbb{R}}^{2}))$ where $\frac{1}{q_{1}}=\frac{1}{r_{0}}+\frac{1}{r}$ satisfies $q_{1}>q_{0}$. In particular, the terms appearing in the problem above are more regular in space than $B(w,w)$. This opens the door to a further iteration. In particular, by considering the solution $v_{1}$ to
[…]
and studying the problem for $\overline{v}_{2}=\overline{v}_{1}-v_{1}$, one can check that the above procedure leads to a further improvement. The idea is to stop the iteration whenever the forcing terms appearing in the procedure are regular enough to build Leray-type solutions to the corresponding PDE. Before going further, let us stress that the above procedure is reminiscent of the so-called ‘DaPrato-Debussche trick’ introduced in [19] and now is widely used in the context of stochastic PDEs.
Let us now turn to the construction of a $q$-solution to (3.4). The above argument motivates the following splitting. Let $N$ be a positive integer such that
(3.5) $r\in\left[\frac{2(N+2)}{N+1},\frac{2(N+1)}{N}\right)$
then we search of a solution $v$ to (3.4) given by a sum of $N+1$ terms
(3.6) $v=\sum_{i=0}^{N-1}v_{i}+\overline{v}$
where $v_{i}$ and $\overline{v}$ solve the following system of PDEs on $[0,T]$:
[…]
where, $\sum_{j=0}^{-1}:=0$, $i\in\{1,\dots,N-1\}$ and
(3.8) $q_{i}=\frac{2r}{r+2+(i+1)(2-r)}.$
Note that $v_{i}$ for $i\in\{1,\dots,N-1\}$ solves a (linear) Stokes problem, while the problem for $\overline{v}$ is a modified version of the Navier-Stokes equations.
At least formally, it is clear that $v$ solves (3.4). The latter fact is a straightforward consequence of the following identity (letting $v_{N}:=\overline{v}$ for simplicity)
$\sum_{i,j=0}^{N}B(v_{i},v_{j})$ $=\sum_{i=0}^{N}B(v_{i},v_{i})+\sum_{0\leq j<i\leq N}B(v_{i},v_{j})+\sum_{0\leq i<j\leq N}B(v_{i},v_{j})$
$=\sum_{i=0}^{N}B(v_{i},v_{i})+\sum_{i=0}^{N-1}\sum_{j=0}^{i-2}B(v_{i-1},v_{j})+\sum_{j=0}^{N-1}\sum_{i=0}^{j-2}B(v_{i},v_{j-1})$
$=B(\overline{v},\overline{v})+\sum_{i=0}^{N-1}\sum_{j=0}^{i-1}B(v_{i-1},v_{j})+\sum_{j=0}^{N-1}\sum_{i=0}^{j-2}B(v_{i},v_{j-1}).$
To show rigorously that $v$ given in (3.6) with $(v_{0},\dots,v_{N-1},\overline{v})$ solving (3.7) is a $q$-solution to (3.4) we need to check that $v_{0},\dots,v_{N-1}$ and $\overline{v}$ are sufficiently regular. The appropriate regularity class for $v$ in (3.6) to obtain a $q$-solution is given in the following definition, see also 3.6 below.
Definition 3.5.
Given $r\in(2,4)$, $N$ given by (3.5), $p\geq 2^{N}\frac{r}{r-2}$, we say that
$(v_{0},\dots,v_{N-1},\overline{v})$
is a $(p,r)$-solution of (3.7) if
(3.9) $v_{i}$ $\in W^{1,{p/2^{i}}}(0,T;X_{-1/2,A_{q_{i}}})\cap L^{{p/2^{i}}}(0,T;X_{1/2,A_{q_{i}}}),$
$\overline{v}$ $\in C([0,T];H)\cap L^{2}(0,T;V),$
where $q_{i}$ is as in (3.8), and for each
$(\phi_{0},\dots,\phi_{N-1},\overline{\phi})\quad\text{s.t. }\phi_{i}\in\mathsf{D}(A_{q_{i}^{\prime}}),\ \overline{\phi}\in\mathsf{D}(A)$
we have, for all $t\in[0,T]$,
(3.10) $\langle v_{0}(t),\phi_{0}\rangle$ $=-\int_{0}^{t}\langle v_{0}(s),A_{q_{0}^{\prime}}\phi_{0}\rangle\,\mathrm{d}s+\int_{0}^{t}\langle w(s)\otimes w(s),\nabla\phi_{0}\rangle\,\mathrm{d}s,$
(3.11) $\langle v_{i}(t),\phi_{i}\rangle$ $=-\int_{0}^{t}\langle v_{i}(s),A_{q_{i}^{\prime}}\phi_{i}\rangle\,\mathrm{d}s+\int_{0}^{t}\langle v_{i-1}(s)\otimes v_{i-1}(s),\nabla\phi_{i}\rangle\,\mathrm{d}s$
$+\int_{0}^{t}\langle v_{i-1}(s)\otimes\left(w(s)+\sum_{j=0}^{i-2}v_{j}(s)\right),\nabla\phi_{i}\rangle\,\mathrm{d}s$
$+\int_{0}^{t}\langle\left(w(s)+\sum_{j=0}^{i-2}v_{j}(s)\right)\otimes v_{i-1}(s),\nabla\phi_{i}\rangle\,\mathrm{d}s,$
(3.12) $\langle\overline{v}(t),\overline{\phi}\rangle$ $=\langle u_{\mathrm{in}},\overline{\phi}\rangle-\int_{0}^{t}\langle\overline{v}(s),A\overline{\phi}\rangle\,\mathrm{d}s+\int_{0}^{t}\langle\overline{v}(s)\otimes\overline{v}(s),\nabla\overline{\phi}\rangle\,\mathrm{d}s$
$+\int_{0}^{t}\langle\overline{v}(s)\otimes\left(w(s)+\sum_{j=0}^{N-1}v_{j}(s)\right),\nabla\overline{\phi}\rangle\,\mathrm{d}s$
$+\int_{0}^{t}\langle\left(w(s)+\sum_{j=0}^{N-1}v_{j}(s)\right)\otimes\overline{v}(s),\nabla\overline{\phi}\rangle\,\mathrm{d}s$
$+\int_{0}^{t}\langle v_{N-1}(s)\otimes v_{N-1}(s),\nabla\overline{\phi}\rangle\,\mathrm{d}s$
$+\int_{0}^{t}\langle v_{N-1}(s)\otimes\left(w(s)+\sum_{j=0}^{N-2}v_{j}(s)\right),\nabla\overline{\phi}\rangle\,\mathrm{d}s$
$+\int_{0}^{t}\langle\left(w(s)+\sum_{j=0}^{N-2}v_{j}(s)\right)\otimes v_{N-1}(s),\nabla\overline{\phi}\rangle\,\mathrm{d}s.$
Remark 3.6.
We observe that $q_{i}>1$ for all $i\in\{0,\dots,N-1\}$ and is increasing in $i$. As an immediate consequence of 3.5, Sobolev embedding theorem and interpolation we have that
$v_{i}\in L^{p/2^{i}}(0,T;\mathbb{L}^{r_{i}})\cap C([0,T];H),\quad r_{i}=\frac{2r}{(i+1)(2-r)+2}.$
In particular, $r_{i}>2$ for all $i\in\{0,\dots,N-1\}$ and is increasing in $i$. Therefore one can easily check that all the duality pairings in 3.5 are well defined. Moreover, for all $i\in\{1,\dots,N-1\}$,
$v_{i},\overline{v}\in L^{\frac{2r}{r-2}}(0,T;\mathbb{L}^{r}).$
Indeed, the above assertion for $v_{i}$ follows from $p\geq 2^{N}\frac{r}{r-2}$. While for $\overline{v}$ we can use the standard interpolation inequality $L^{2}(0,T;H^{1})\cap L^{\infty}(0,T;L^{2})\subseteq L^{2/\theta}(0,T;H^{\theta})$ with $\theta=\frac{r-2}{r}\in(0,1)$ and the Sobolev embedding $H^{\theta}(\mathcal{O})\hookrightarrow L^{r}(\mathcal{O})$.
In particular, if $(v_{0},\dots,v_{N-1},\overline{v})$ is a $(p,r)$ solution of (3.7), then, given $v:=\sum_{i=0}^{N-1}v_{i}+\overline{v}$, $u=v+w$ is a $r/2$-solution of (1.2) in the sense of 1.2.
The following yields the well-posedness of (3.4) in the sense of 3.5.
Theorem 3.7.
Let 3.4 be satisfied. For each $u_{\mathrm{in}}\in H$, $p\geq 2^{N}\frac{r}{r-2}$ there exists a unique $(p,r)$-solution $(v_{0},\dots,v_{N-1},\overline{v})$ of (3.7) in the sense of 3.5. Moreover $\overline{v}$ satisfies the energy relation
$\lVert\overline{v}(t)\rVert^{2}+2\int_{0}^{T}\lVert\nabla\overline{v}(s)\rVert_{L^{2}}^{2}\,\mathrm{d}s$ $=\lVert u_{\mathrm{in}}\rVert^{2}+2\int_{0}^{t}\langle\overline{v}(s)\cdot\nabla\overline{v}(s),w(s)+\sum_{j=0}^{N-1}v_{j}(s)\rangle\,\mathrm{d}s$
$+2\int_{0}^{t}\langle v_{N-1}(s)\cdot\nabla\overline{v}(s),w(s)+\sum_{j=0}^{N-1}v_{j}(s)\rangle\,\mathrm{d}s$
(3.13) $+2\int_{0}^{t}\langle(w(s)+\sum_{j=0}^{N-2}v_{j}(s))\cdot\nabla\overline{v}(s),v_{N-1}(s)\rangle\,\mathrm{d}s.$
If $\left(u_{\mathrm{in}}^{n}\right)_{n\in{\mathbb{N}}}$ is a sequence in $H$ converging to $u_{\mathrm{in}}\in H$ and $(w^{n})_{n\in{\mathbb{N}}}$ is a sequence in $C([0,T];\mathbb{L}^{r})$ converging to $w\in C([0,T];\mathbb{L}^{r})$, then the corresponding unique solutions $((v_{0}^{n},\dots,v_{N-1}^{n},\overline{v}^{n}))_{n\in{\mathbb{N}}}$ converge to the corresponding solution $(v_{0},\dots,v_{N-1},\overline{v})$, each one in the topologies of 3.5.
Proof.
We exploit strongly the triangle structure of (3.7) and split the proof in several steps.
Step 1: Linear part of (3.7). We argue by induction and exploit maximal $L^{p}$ regularity techniques, see [53, Chapter 3]. The existence and uniqueness of $v_{0}$ satisfying the corresponding PDE in the sense of 3.5 and the continuous dependence from data, i.e. $w$ in the topology of $C([0,T];\mathbb{L}^{r})$, follows if
$B(w,w)\in L^{p}(0,T;X_{-1/2,A_{q_{0}}}).$
The claim is true, indeed $q_{0}=\frac{r}{2}$ and by Hölder’s inequality we have
$\int_{0}^{T}\lVert B(w(s),w(s))\rVert_{X_{-1/2,A_{r/2}}}^{p}\,\mathrm{d}s$ $\leq\int_{0}^{T}\lVert w(s)\rVert_{\mathbb{L}^{r}}^{2p}\,\mathrm{d}s\leq T\lVert w\rVert_{C([0,T];\mathbb{L}^{r})}^{2p}.$
Now assume we have already proved the existence and uniqueness of $(v_{i})_{i\in\{0,\dots,l-1\}},$ $l\leq N-1$ solving (3.7) in the sense of 3.5 and depending continuously from the data, i.e. $w$ in the topology of $C([0,T];\mathbb{L}^{r})$. Let us check that there exists a unique $v_{l}$ solving the corresponding PDE in (3.7) in the sense of 3.5 and depending continuously from $w$ in the topology of $C([0,T];\mathbb{L}^{r})$. Again, due to maximal $L^{p}$ regularity techniques, it is enough to show that
$B(v_{l-1},v_{l-1})+B(v_{l-1},w+\sum_{j=0}^{l-2}v_{j})-B(w+\sum_{j=0}^{l-2}v_{j},v_{l-1})\in L^{p/2^{l}}(0,T;X_{-1/2,A_{q_{l}}}).$
The claim is true, indeed due to 3.6
$w,\ v_{i}\in L^{p/2^{l-1}}(0,T;\mathbb{L}^{r})\quad\text{if }i\in\{0,\dots,l-2\}$
and by induction hypothesis
$v_{l-1}\in L^{p/2^{l-1}}(0,T;\mathbb{L}^{r_{l-1}}).$
Moreover all $v_{i},\ i\in{1,\dots,l-1}$ depends continuously from $w\in C([0,T];\mathbb{L}^{r})$ in the corresponding topologies. Therefore by Hölder’s inequality we have
$\int_{0}^{T}\lVert B(v_{l-1}$ $(s),w(s)+\sum_{j=0}^{l-2}v_{j}(s))\rVert_{X_{-1/2,A_{q_{l}}}}^{p/2^{l}}\,\mathrm{d}s$
$\leq\int_{0}^{T}\left\lVert v_{l-1}(s)\otimes\left(w(s)+\sum_{j=0}^{l-2}v_{j}(s)\right)\right\rVert_{\mathbb{L}^{q_{l}}}^{p/2^{l}}\,\mathrm{d}s$
$\lesssim_{p,l}\int_{0}^{T}\lVert v_{l-1}(s)\rVert_{\mathbb{L}^{r_{l-1}}}^{p/2^{l}}\left(\lVert w(s)\rVert_{\mathbb{L}^{r_{l-1}}}^{p/2^{l}}+\sum_{j=0}^{l-2}\lVert v_{j}(s)\rVert_{\mathbb{L}^{r_{l-1}}}^{p/2^{l}}\right)\,\mathrm{d}s$
$\lesssim\lVert v_{l-1}\rVert^{p/2^{l-1}}_{L^{p/2^{l-1}}(0,T;\mathbb{L}^{r_{l-1}})}+\lVert w\rVert^{p/2^{l-1}}_{L^{p/2^{l-1}}(0,T;\mathbb{L}^{q})}+\sum_{j=0}^{l-2}\lVert v_{j}\rVert^{p/2^{l-1}}_{L^{p/2^{l-1}}(0,T;\mathbb{L}^{r})}.$
Step 2: Introduction to the nonlinear part of (3.7). First, we observe that due to Step 1 we have that
(3.14) $\overline{f}$ $=-B(v_{N-1},w+\sum_{j=0}^{N-1}v_{j})-B(w+\sum_{j=0}^{N-2}v_{j},v_{N-1})\in L^{2}(0,T;V^{*}),$
(3.15) $\tilde{v}$ $=w+\sum_{j=0}^{N-1}v_{j}\in L^{p/2^{N-1}}(0,T;\mathbb{L}^{r}).$
Therefore we are left to study the well-posedness in the weak setting of the following PDE
[…]
This can be treated similarly to [1, Section 3.2] and is the object of the remaining steps.
Step 3: Uniqueness. Let $\overline{v}^{\left(i\right)}$ be two solutions. The function $z=\overline{v}^{\left(1\right)}-\overline{v}^{\left(2\right)}$ satisfies hence
$\left\langle z\left(t\right),\overline{\phi}\right\rangle+\int_{0}^{t}\left\langle z\left(s\right),A\overline{\phi}\right\rangle\,\mathrm{d}s-\int_{0}^{t}\langle z(s)\cdot\nabla\overline{\phi},z(s)\rangle\,\mathrm{d}s=\int_{0}^{t}\left\langle\widetilde{f}\left(s\right),\overline{\phi}\right\rangle\,\mathrm{d}s$
where
$\widetilde{f}=-B\left(\overline{v}^{\left(2\right)}+\tilde{v},z\right)-B\left(z,\overline{v}^{\left(2\right)}+\tilde{v}\right).$
By 2.6 and [29, Lemma 1.14], $\widetilde{f}\in L^{2}\left(0,T;V^{*}\right)$. Then, by Theorem 2.5,
$\lVert z\left(t\right)\rVert^{2}+2\int_{0}^{t}\lVert\nabla z\left(s\right)\rVert_{L^{2}}^{2}\,\mathrm{d}s=2\int_{0}^{t}\langle z(s)\cdot\nabla z(s),\overline{v}^{\left(2\right)}(s)+\tilde{v}(s)\rangle\,\mathrm{d}s.$
Again by 2.6, we have
$\int_{0}^{t}\langle z(s)\cdot$ $\nabla z(s),\overline{v}^{\left(2\right)}(s)+\tilde{v}(s)\rangle\,\mathrm{d}s$
$\leq\left\lvert\int_{0}^{t}\langle z(s)\cdot\nabla z(s),\overline{v}^{\left(2\right)}(s)\rangle\,\mathrm{d}s\right\rvert+\left\lvert\int_{0}^{t}\langle z(s)\cdot\nabla z(s),\tilde{v}(s)\rangle\,\mathrm{d}s\right\rvert$
$\leq 2\varepsilon\lVert z\rVert_{L^{2}(0,t;V)}^{2}+\frac{C}{\varepsilon^{3}}\int_{0}^{t}\lVert z(s)\rVert^{2}\lVert\overline{v}^{\left(2\right)}(s)\rVert_{\mathbb{L}^{4}}^{4}\,\mathrm{d}s$
$+2\varepsilon\lVert z\rVert_{L^{2}(0,t;V)}^{2}+\frac{C}{\varepsilon^{\frac{r+2}{r-2}}}\int_{0}^{t}\lVert z(s)\rVert^{2}\lVert\tilde{v}(s)\rVert_{\mathbb{L}^{r}}^{\frac{2r}{r-2}}\,\mathrm{d}s$
$=4\varepsilon\int_{0}^{t}\lVert\nabla z(s)\rVert_{L^{2}}^{2}\,\mathrm{d}s+\frac{C}{\varepsilon^{\frac{r+2}{r-2}}}\int_{0}^{t}\lVert z(s)\rVert^{2}\left(\lVert\overline{v}^{\left(2\right)}(s)\rVert_{\mathbb{L}^{4}}^{4}+\lVert\tilde{v}(s)\rVert_{\mathbb{L}^{r}}^{\frac{2r}{r-2}}\right)\,\mathrm{d}s.$
Applying the above with $4\varepsilon=\frac{1}{2}$ and renaming the constant $C$, it follows that
$\lVert z\left(t\right)\rVert^{2}+\int_{0}^{t}\lVert\nabla z\left(s\right)\rVert_{L^{2}}^{2}\,\mathrm{d}s\leq C\int_{0}^{t}\lVert z\left(s\right)\rVert^{2}\left(\lVert\overline{v}^{\left(2\right)}(s)\rVert_{\mathbb{L}^{4}}^{4}+\lVert\tilde{v}(s)\rVert_{\mathbb{L}^{r}}^{\frac{2r}{r-2}}\right)\,\mathrm{d}s.$
We conclude $z=0$ by the Gronwall lemma, using (3.15) and the integrability properties of $\overline{v}^{\left(2\right)}$.
Step 4: Existence. Define the sequence $\left(\overline{v}^{n}\right)$ by setting $\overline{v}^{0}=0$ and for every $n\geq 0$, given $\overline{v}^{n}\in C\left(\left[0,T\right];H\right)\cap L^{2}\left(0,T;V\right)$, let $\overline{v}^{n+1}$ be the solution of equation (2.8) with initial condition $u_{\mathrm{in}}$ and with
$f=-B\left(\overline{v}^{n},\tilde{v}\right)-B\left(\tilde{v},\overline{v}^{n}\right)+\overline{f}.$
In particular
$\left\langle\overline{v}^{n+1}\left(t\right),\overline{\phi}\right\rangle+\int_{0}^{t}\left\langle\overline{v}^{n+1}\left(s\right),A\overline{\phi}\right\rangle\,\mathrm{d}s-\int_{0}^{t}\left\langle\overline{v}^{n+1}\left(s\right)\cdot\nabla\overline{\phi},\overline{v}^{n+1}\left(s\right)\right\rangle\,\mathrm{d}s$
$=\left\langle u_{\mathrm{in}},\overline{\phi}\right\rangle+\int_{0}^{t}\langle f(s),\overline{\phi}\rangle\,\mathrm{d}s$
for every $\overline{\phi}\in\mathsf{D}\left(A\right)$. The above is well-defined as
$B\left(\overline{v}^{n},\tilde{v}\right),B\left(\tilde{v},\overline{v}^{n}\right),\overline{f}\in L^{2}\left(0,T;V^{*}\right)$
by 2.6 and (3.14).
Then let us investigate the convergence of $\left(\overline{v}^{n}\right)$. First, let us prove a bound. From the previous identity and Theorem 2.5 we get
$\lVert\overline{v}^{n+1}\left(t\right)\rVert^{2}+2\int_{0}^{t}\lVert\nabla\overline{v}^{n+1}\left(s\right)\rVert_{L^{2}}^{2}\,\mathrm{d}s$
$=\lVert u_{\mathrm{in}}\rVert^{2}+2\int_{0}^{t}\left(b\left(\overline{v}^{n},\overline{v}^{n+1},\tilde{v}\right)+b\left(\tilde{v},\overline{v}^{n+1},\overline{v}^{n}\right)+\langle\overline{f},\overline{v}^{n+1}\rangle\right)\left(s\right)\,\mathrm{d}s.$
It gives us (using 2.6 and (3.14))
$\lVert\overline{v}^{n+1}\left(t\right)\rVert^{2}+\int_{0}^{t}\lVert\nabla\overline{v}^{n+1}\left(s\right)\rVert_{L^{2}}^{2}\,\mathrm{d}s$
$=\lVert u_{\mathrm{in}}\rVert^{2}+\varepsilon\int_{0}^{t}\lVert\overline{v}^{n}\left(s\right)\rVert_{V}^{2}\,\mathrm{d}s$
$+C_{\varepsilon}\int_{0}^{t}\lVert\overline{v}^{n}\left(s\right)\rVert^{2}\lVert\tilde{v}(s)\rVert_{\mathbb{L}^{r}}^{\frac{2r}{r-2}}\,\mathrm{d}s+C_{\varepsilon}\int_{0}^{t}\lVert\overline{f}\left(s\right)\rVert_{V^{*}}^{2}\,\mathrm{d}s.$
Choosing a small constant $\varepsilon$, one can find $R>\lVert u_{\mathrm{in}}\rVert^{2}$ and $\overline{T}$ small enough, depending only from $\lVert u_{\mathrm{in}}\rVert$ and $\lVert\tilde{v}\rVert_{L^{\frac{2r}{r-2}}(0,T;\mathbb{L}^{r})}$, such that if
(3.17) $\sup_{t\in\left[0,\overline{T}\right]}\lVert\overline{v}^{n}\left(t\right)\rVert^{2}\leq R,\qquad\int_{0}^{\overline{T}}\lVert\overline{v}^{n}\left(s\right)\rVert_{V}^{2}\,\mathrm{d}s\leq R$
then the same inequalities hold for $\overline{v}^{n+1}$.
Set $z_{n}=\overline{v}^{n}-\overline{v}^{n-1}$, for $n\geq 1$. From the identity above,
$\left\langle z_{n+1}\left(t\right),\overline{\phi}\right\rangle-\int_{0}^{t}\left(b\left(\overline{v}^{n+1},\overline{\phi},\overline{v}^{n+1}\right)-b\left(\overline{v}^{n},\phi,\overline{v}^{n}\right)\right)\left(s\right)\,\mathrm{d}s$
$=-\int_{0}^{t}\left\langle z_{n+1}\left(s\right),A\overline{\phi}\right\rangle\,\mathrm{d}s-\int_{0}^{t}\left\langle\left(B\left(\overline{v}^{n},\tilde{v}\right)-B\left(\overline{v}^{n-1},\tilde{v}\right)\right)\left(s\right),\overline{\phi}\right\rangle\,\mathrm{d}s$
$-\int_{0}^{t}\left\langle\left(B\left(\tilde{v},\overline{v}^{n}\right)-B\left(\tilde{v},\overline{v}^{n-1}\right)\right)\left(s\right),\overline{\phi}\right\rangle\,\mathrm{d}s.$
Since
$b\left(\overline{v}^{n+1},\overline{\phi},\overline{v}^{n+1}\right)-b\left(\overline{v}^{n},\overline{\phi},\overline{v}^{n}\right)-b\left(z_{n+1},\overline{\phi},z_{n+1}\right)$
$=b\left(\overline{v}^{n},\overline{\phi},z_{n+1}\right)+b\left(z_{n+1},\overline{\phi},\overline{v}^{n}\right)$
we may rewrite it as
$\left\langle z_{n+1}\left(t\right),\overline{\phi}\right\rangle-\int_{0}^{t}b\left(z_{n+1}\left(s\right),\overline{\phi},z_{n+1}\left(s\right)\right)\,\mathrm{d}s$
$=-\int_{0}^{t}\left\langle z_{n+1}\left(s\right),A\overline{\phi}\right\rangle\,\mathrm{d}s-\int_{0}^{t}\left\langle\left(B\left(z_{n},\tilde{v}\right)+B\left(\tilde{v},z_{n}\right)\right)\left(s\right),\phi\right\rangle\,\mathrm{d}s$
$+\int_{0}^{t}\left(b\left(\overline{v}^{n},\overline{\phi},z_{n+1}\right)+b\left(z_{n+1},\overline{\phi},\overline{v}^{n}\right)\right)\left(s\right)\,\mathrm{d}s.$
One can check as above the applicability of Theorem 2.5 and get
$\lVert z_{n+1}\left(t\right)\rVert^{2}+2\int_{0}^{t}\lVert\nabla z_{n+1}\left(s\right)\rVert_{L^{2}}^{2}\,\mathrm{d}s$
$=2\int_{0}^{t}\left(b\left(z_{n},z_{n+1},\tilde{v}\right)+b\left(\tilde{v},z_{n+1},z_{n}\right)\right)\left(s\right)\,\mathrm{d}s$
$+2\int_{0}^{t}b\left(z_{n+1},z_{n+1},\overline{v}^{n}\right)\left(s\right)\,\mathrm{d}s.$
As above, thanks to [29, Lemma 1.14] we deduce that
$\int_{0}^{t}\lvert b\left(z_{n+1},z_{n+1},\overline{v}^{n}\right)(s)\rvert\,\mathrm{d}s\leq\frac{1}{4}\int_{0}^{t}\lVert z_{n+1}(s)\rVert_{V}^{2}\,\mathrm{d}s+C\int_{0}^{t}\lVert z_{n+1}(s)\rVert^{2}\lVert\overline{v}^{n}(s)\rVert_{\mathbb{L}^{4}}^{4}\,\mathrm{d}s.$
But
$\int_{0}^{t}\lvert b\left(z_{n},z_{n+1},\tilde{v}\right)(s)+b\left(\tilde{v},z_{n+1},z_{n}\right)(s)\rvert\,\mathrm{d}s$
$\leq\frac{1}{4}\int_{0}^{t}\lVert z_{n+1}(s)\rVert_{V}^{2}\,\mathrm{d}s+\frac{1}{8}\int_{0}^{t}\lVert z_{n}(s)\rVert_{V}^{2}\,\mathrm{d}s+C\int_{0}^{t}\lVert z_{n}(s)\rVert^{2}\lVert\tilde{v}(s)\rVert_{\mathbb{L}^{r}}^{\frac{2r}{r-2}}\,\mathrm{d}s.$
Hence
$\lVert z_{n+1}\left(t\right)\rVert^{2}+\int_{0}^{t}\lVert\nabla z_{n+1}\left(s\right)\rVert_{L^{2}}^{2}\,\mathrm{d}s$
$\leq C\int_{0}^{t}\lVert z_{n+1}\left(s\right)\rVert^{2}\lVert\overline{v}^{n}\left(s\right)\rVert_{\mathbb{L}^{4}}^{4}\,\mathrm{d}s$
$+\frac{1}{4}\int_{0}^{t}\lVert z_{n}\left(s\right)\rVert_{V}^{2}\,\mathrm{d}s+C\int_{0}^{t}\lVert z_{n}\left(s\right)\rVert^{2}\lVert\tilde{v}\left(s\right)\rVert_{\mathbb{L}^{r}}^{\frac{2r}{r-2}}\,\mathrm{d}s.$
Now we work under the bounds (3.17) and deduce, using the Gronwall lemma, for $\overline{T}$, depending only from $\lVert u_{\mathrm{in}}\rVert$ and $\lVert\tilde{v}\rVert_{L^{\frac{2r}{r-2}}(0,T;\mathbb{L}^{r})}$, possibly smaller than the previous one,
$\sup_{t\in\left[0,\overline{T}\right]}\lVert z_{n+1}\left(t\right)\rVert^{2}+\int_{0}^{\overline{T}}\lVert z_{n+1}\left(s\right)\rVert_{V}^{2}\,\mathrm{d}s$
$\leq\frac{1}{2}\left(\sup_{t\in\left[0,\overline{T}\right]}\lVert z_{n}\left(t\right)\rVert^{2}+\int_{0}^{\overline{T}}\lVert z_{n}\left(s\right)\rVert_{V}^{2}\,\mathrm{d}s\right).$
It implies that the sequence $\left(\overline{v}^{n}\right)$ is Cauchy in $C\left(\left[0,\overline{T}\right];H\right)\cap L^{2}\left(0,\overline{T};V\right)$. The limit $\overline{v}$ has the right regularity to be a weak solution and satisfies the weak formulation; in the identity above for $\overline{v}^{n+1}$ and $\overline{v}^{n}$ we may prove that
$\int_{0}^{t}b\left(\overline{v}^{n+1}\left(s\right),\overline{\phi},\overline{v}^{n+1}\left(s\right)\right)\,\mathrm{d}s$ $\rightarrow\int_{0}^{t}b\left(\overline{v}\left(s\right),\overline{\phi},\overline{v}\left(s\right)\right)\,\mathrm{d}s$
$\int_{0}^{t}b\left(\overline{v}^{n}\left(s\right),\overline{\phi},\tilde{v}\left(s\right)\right)\,\mathrm{d}s$ $\rightarrow\int_{0}^{t}b\left(\overline{v}\left(s\right),\overline{\phi},\tilde{v}\left(s\right)\right)\,\mathrm{d}s$
$\int_{0}^{t}b\left(\tilde{v}\left(s\right),\overline{\phi},\overline{v}^{n}\left(s\right)\right)\,\mathrm{d}s$ $\rightarrow\int_{0}^{t}b\left(\tilde{v}\left(s\right),\overline{\phi},\overline{v}\left(s\right)\right)\,\mathrm{d}s.$
All these convergences can be proved easily by recalling the definition of $b$. Similarly, we can pass to the limit in the energy identity. After proving existence and uniqueness in $[0,\overline{T}]$ we can reiterate the existence procedure and in a finite number of steps cover the interval $[0,T]$.
Step 5: Continuity dependence on the data Let $\overline{v}^{n}$ (resp. $\overline{v}$) the unique solution of (3.16) with data $u_{\mathrm{in}}^{n},\ \bar{f}^{n},\ \tilde{v}^{n}$ (resp. $u_{\mathrm{in}},\ \bar{f},\ \tilde{v}$). Since $u_{0}^{n}\rightarrow u_{0}$ in $H$ (resp. $\overline{f}^{n}\rightarrow\overline{f}$ in $L^{2}(0,T;V^{*})$, $\tilde{v}^{n}\rightarrow\tilde{v}$ in $L^{p/2^{N-1}}(0,T;\mathbb{L}^{r})$) the family $(u_{\mathrm{in}}^{n})_{n\in\mathbb{N}}$ is bounded in $H$ (resp. the family $(\overline{f}^{n})_{n\in\mathbb{N}}$ is bounded in $L^{2}(0,T;V^{*})$, the family $(\widetilde{v}^{n})_{n\in\mathbb{N}}$ is bounded in $L^{p/2^{N-1}}(0,T;\mathbb{L}^{r})$), by (3.7) one can show easily that the family $(\overline{v}^{n})_{n\in\mathbb{N}}$ is bounded in $C([0,T];H)\cap L^{2}(0,T;V).$ Moreover for each $t\in[0,T]$, $z^{n}=\overline{v}^{n}-\overline{v}$ satisfies the energy relation
$\frac{1}{2}\lVert z^{n}(t)\rVert^{2}+\int_{0}^{t}\lVert\nabla z^{n}(s)\rVert_{L^{2}}^{2}\,\mathrm{d}s$ $=\frac{1}{2}\lVert u_{\mathrm{in}}^{n}-u_{\mathrm{in}}\rVert^{2}$
$+\int_{0}^{t}b(z^{n}(s),z^{n}(s),\overline{v}(s))\,\mathrm{d}s$
$+\int_{0}^{t}b(\overline{v}^{n}(s),z^{n}(s),\widetilde{v}^{n}(s)-\widetilde{v}(s))\,\mathrm{d}s$
$+\int_{0}^{t}b(z^{n}(s),z^{n}(s),\widetilde{v}(s))\,\mathrm{d}s$
$+\int_{0}^{t}b(\widetilde{v}^{n}(s)-\widetilde{v}(s),z^{n}(s),\overline{v}(s))\,\mathrm{d}s$
(3.18) $+\int_{0}^{t}\langle\overline{f}^{n}(s)-\overline{f}(s),z^{n}(s)\rangle\,\mathrm{d}s.$
We can easily bound the right-hand side of relation (3.2) by Young’s inequality and Hölder’s inequality obtaining
$\frac{1}{2}\lVert z^{n}(t)\rVert^{2}+\frac{1}{2}\int_{0}^{t}\lVert\nabla z^{n}(s)\rVert_{L^{2}}^{2}\,\mathrm{d}s$ $\leq\frac{1}{2}\lVert u_{\mathrm{in}}^{n}-u_{\mathrm{in}}\rVert^{2}+\int_{0}^{t}\lVert\overline{f}^{n}(s)-\overline{f}(s)\rVert_{V^{*}}^{2}\,\mathrm{d}s$
$+C\int_{0}^{t}\lVert z^{n}(s)\rVert^{2}\left(\lVert\overline{v}(s)\rVert_{\mathbb{L}^{4}}^{4}+\lVert\tilde{v}(s)\rVert_{\mathbb{L}^{r}}^{\frac{2r}{r-2}}\right)\,\mathrm{d}s$
$+C\lVert\widetilde{v}^{n}-\widetilde{v}\rVert_{L^{\frac{2r}{r-2}}(0,T;\mathbb{L}^{r})}^{2}\lVert\overline{v}^{n}\rVert_{C([0,T];H)}^{\frac{2(r-2)}{r}}\lVert\overline{v}^{n}\rVert_{L^{2}(0,T;V)}^{\frac{4}{r}}$
(3.19) $+C\lVert\widetilde{v}^{n}-\widetilde{v}\rVert_{L^{\frac{2r}{r-2}}(0,T;\mathbb{L}^{r})}^{2}\lVert\overline{v}\rVert_{C([0,T];H)}^{\frac{2(r-2)}{r}}\lVert\overline{v}\rVert_{L^{2}(0,T;V)}^{\frac{4}{r}}.$
Applying Gronwall’s inequality to relation (3.2) the claimed continuity follows. ∎
Remark 3.8.
Freezing the variable $\omega\in\Omega$ and solving (3.4) for each $\omega$ does not allow us to obtain information about the measurability properties of $v$. However, the measurability of $v$ with respect to the progressive $\sigma$-algebra follows from the continuity of the solution map with respect to $u_{\mathrm{in}}$ and $w$. Therefore we have the required measurability properties for $v$ with $w$ being the mild solution of (1.3). In particular $v$ has $\mathbf{P}$-a.s. paths in $C(0,T;H)\cap L^{\frac{2q}{q-1}}(0,T;\mathbb{L}^{2q})$ for each $r\in(1,q_{\mathcal{H}})$, it is progressively measurable with respect to these topologies.
Combining 3.1, Theorem 3.7 and 3.8 we get immediately the existence of a $q-$solution of equation (1.2) in the sense of 1.2 for each $q\in(1,q_{\mathcal{H}})$.
3.3. Proof of Theorem 1.3(1)
As discussed above, the results of subsection 3.1, subsection 3.2 provide the existence of a $q$-solution of equation (1.2) in the sense of 1.2 for each $q\in(1,q_{\mathcal{H}})$, moreover such a solution is adapted with paths in $C([0,T];H)$ due to 3.6. In order to conclude the proof of Theorem 1.31 it is enough to show the following:
Proposition 3.9 (Uniqueness).
Let $u_{1}$ be a $q_{1}$-solution of (1.2) and $u_{2}$ be a $q_{2}$-solution of (1.2). Then $u_{1}\equiv u_{2}$.
The uniqueness result in the Ladyzhenskaya–Prodi–Serrin class of 3.9 might be known to experts. Here, for completeness, we provide a relatively short proof relying on maximal $L^{p}$-regularity techniques which seem not standard even in the absence of noise.
Proof of 3.9.
Recall that $u_{i}=w_{g}+v_{i}$ where $v_{i}$ solves the modified Navier-Stokes equations for $i\in\{1,2\}$ by 1.2. We split the proof into two cases.
[…]
$\langle\delta(t),\varphi\rangle$ $-\int_{0}^{t}\langle\delta(s),\Delta\varphi\rangle\,\mathrm{d}s$
$=\int_{0}^{t}b(\delta(s),\varphi,v_{1}(s)+w_{g}(s))\,\mathrm{d}s+\int_{0}^{t}b(v_{2}(s)+w_{g}(s),\varphi,\delta(s))\,\mathrm{d}s.$
As $v_{i}\in L^{2q^{\prime}}(0,T;L^{2q}(\mathcal{O};{\mathbb{R}}^{2}))$ for $i\in\{1,2\}$ by 1.2, we obtain
$B(\delta,v_{1}+w_{g}),B(v_{2}+w_{g},\delta)\in L^{q^{\prime}}(0,T;X_{-1/2,A_{q}})\ \ \mathbf{P}-a.s.$
Hence, by the density of divergence-free $\varphi\in C^{\infty}(\mathcal{O};{\mathbb{R}}^{2})$ such that $\varphi=0$ on $\Gamma_{b}\cup\Gamma_{u}$ in the domain of the Stokes operator $A_{q}$ and from the maximal $L^{q}$-regularity of $A_{q}$, it follows that
$\delta$ $\in W^{1,q^{\prime}}(0,T;X_{-1/2,A_{q}})\cap L^{q^{\prime}}(0,T;X_{1/2,A_{q}})$
$\subset C([0,T];B^{1-2/q^{\prime}}_{q,q^{\prime}}(\mathcal{O};{\mathbb{R}}^{2}))\ \ \mathbf{P}-a.s.$
where in the last step we used the trace embedding [53, Theorem 3.4.8] applied with $A=A_{q}$.
By real interpolation (see e.g. [8, Chapter 6]), we obtain
$(B^{1-2/q^{\prime}}_{q,q^{\prime}}(\mathcal{O}),H^{1,q}(\mathcal{O}))_{1/2,1}\hookrightarrow B^{1-1/q^{\prime}}_{q,1}(\mathcal{O})\hookrightarrow L^{2q}(\mathcal{O})$
where in the last step we applied the Sobolev embedding and $1-\frac{1}{q^{\prime}}-\frac{2}{q}=-\frac{1}{q}$. In particular,
(3.20) $\|f\|_{L^{2q}(\mathcal{O})}\lesssim\|f\|_{B^{1-2/q^{\prime}}_{q,q^{\prime}}(\mathcal{O})}^{1/2}\|f\|_{H^{1,q}(\mathcal{O})}^{1/2}$
for all $f$ for which the right-hand side is finite.
Hence, by maximal $L^{q}$-regularity of $A_{q}$, again the trace embedding [53, Theorem 3.4.8] as well as the Hölder inequality, there exists a constant $C_{0}>0$ independent of $v_{1},v_{2}$ and $\delta$ such that, for all $t\in[0,T]$ and $\mathbf{P}-a.s.$,
$\sup_{r\in[0,t]}\|\delta(r)\|_{B^{1-2/q^{\prime}}_{q,q^{\prime}}}^{q^{\prime}}+\int_{0}^{t}\|\delta(r)\|_{H^{1,q}}^{q^{\prime}}\,\mathrm{d}r$
$\leq C_{0}\int_{0}^{t}\big{(}\max_{i}\|v_{i}(r)\|_{L^{2q}}^{q^{\prime}}+\|w_{g}(r)\|_{L^{2q}}^{q^{\prime}}\big{)}\|\delta(r)\|_{L^{2q}}^{q^{\prime}}\,\mathrm{d}r$
$\leq C_{1}\int_{0}^{t}\big{(}\max_{i}\|v_{i}(r)\|_{L^{2q}}^{2q^{\prime}}+\|w_{g}(r)\|_{L^{2q}}^{2q^{\prime}}\big{)}\|\delta(r)\|_{B^{1-2/q^{\prime}}_{q,q^{\prime}}}^{q^{\prime}}\,\mathrm{d}r+\frac{1}{2}\int_{0}^{t}\|\delta(r)\|_{H^{1,q}}^{q^{\prime}}\,\mathrm{d}r$
where in the last step we used the Young inequality and (3.20).
Now the conclusion follows from the Grownall lemma and the integrability conditions on $v_{1},v_{2}$ due to 1.2 and $w_{g}\in C([0,\infty);L^{2q})$ for all $q<q_{\mathcal{H}}$ by 3.1.
[…]
Lemma 3.10 (Compatibility).
Let $w\in C([0,T];\mathbb{L}^{r})$ for some $r\in(2,4)$ and $2<\widetilde{r}\leq r$. If $(v_{0},\dots,v_{N-1},\overline{v})$ is a solution of (3.7) in the sense of 3.5 with
$p\geq 2^{N}\frac{r}{r-2},\qquad q_{i}=\frac{2r}{r+2+(i+1)(2-r)}$
and $(\widetilde{v}_{0},\dots,\widetilde{v}_{N-1},\widehat{v})$ is a solution of (3.4) in the sense of 3.5 with
$\widetilde{p}\geq 2^{\widetilde{N}}\frac{\widetilde{r}}{\widetilde{r}-2},\qquad\widetilde{q}_{i}=\frac{2\widetilde{r}}{\widetilde{r}+2+(i+1)(2-\widetilde{r})}$
then $v=\widetilde{v}$.
Proof.
The case of $r=\widetilde{r}$ is obvious since in such a case $N=\widetilde{N},\ q_{i}=\widetilde{q}_{i}$ and our construction does not rely on the choice of $p$ so far that $p\geq 2^{N}\frac{r}{r-2}$. In the general case we have two sequences $(v_{0},\dots,v_{N-1},\overline{v})$ and $(\widetilde{v}_{0},\dots,\widetilde{v}_{\widetilde{N}-1},\widetilde{\overline{v}})$. If $N=\widetilde{N}$ the claim is still trivial since our construction does not rely on the choice of $p$ so far that $p\geq 2^{N}\frac{r}{r-2}$ and of the precise choice of the $q_{i}$ since $S_{\widetilde{q}_{i}}(t)|_{\mathbb{L}^{q_{i}}}=S_{q_{i}}(t)$. If $\widetilde{N}>N$ arguing as above we have
$v_{i}=\widetilde{v}_{i}\quad\forall i\in{0,\dots,N-1}$
and we are left to show that $\overline{v}=\sum_{i=N}^{\widetilde{N}-1}\widetilde{v}_{i}+\widetilde{\overline{v}}=:\widehat{v}$. Due to previous steps we can assume that $v$ is $(\widetilde{p},r)$ solution since $\widetilde{p}\geq 2^{\widetilde{N}}\frac{\widetilde{r}}{\widetilde{r}-2}>2^{{N}}\frac{{r}}{{r}-2}$ We observe that due to 3.5 and 3.6,
(3.21) $\overline{v},\widehat{v}\in C([0,T];H)\cap L^{\frac{2\widetilde{r}}{\widetilde{r}-2}}(0,T;\mathbb{L}^{\widetilde{r}}),\ f:=w+\sum_{i=0}^{N-1}v_{i}\in L^{\frac{2\widetilde{r}}{\widetilde{r}-2}}(0,T;\mathbb{L}^{\widetilde{r}}).$
Therefore either $\overline{v}$ and $\widehat{v}$ satisfy for all divergence free vector field $\varphi\in C^{\infty}(\mathcal{O};{\mathbb{R}}^{2})$ such that $\varphi=0$ on $\Gamma_{b}\cup\Gamma_{u}$ equation (3.12). Therefore, denoting by $\delta(t)=\overline{v}-\widehat{v}$ we have that $\delta$ satisfies
$\left\langle\delta\left(t\right),\varphi\right\rangle$ $-\int_{0}^{t}\left\langle z\left(s\right),\Delta\varphi\right\rangle\,\mathrm{d}s$
$=\int_{0}^{t}b(\delta(s),\varphi,\overline{v}(s)+f(s))\,\mathrm{d}s+\int_{0}^{t}b(\widehat{v}(s)+f(s),\varphi,\delta(s))\,\mathrm{d}s.$
Denoting by $\widetilde{q}=\frac{\widetilde{r}}{2}\in(1,2)$, due to relation (3.21)
$B(\delta,\overline{v}+f),\ B(\widehat{v}+f,\delta)\in L^{\widetilde{q}^{\prime}}(0,T;X_{-1/2,A_{\widetilde{q}}}).$
Now the proof proceeds as in the first case of 3.9 and we omit the details. ∎
4. Interior regularity
As announced below subsection 1.1, we prove Theorem 1.3(2). To this end, we first prove the interior regularity of $w_{g}$ and afterwards the one of $v$ by exploiting the decomposition introduced in subsection 3.2.
4.1. Stokes equations
We start showing a lemma concerning the relation between the mild and the weak formulation of (1.3) defined below.
Definition 4.1.
Let 1.1 be satisfied. A stochastic process $w$ is a weak solution of (1.3) if it is $\mathcal{F}$-progressively measurable with $\mathbf{P}-a.s.$ paths in
$w_{g}\in C(0,T;\mathbb{L}^{r})$
for some $r\geq 2$, and $\mathbf{P}-a.s.$ for all $\phi\in\mathsf{D}(A)$ and $t\in[0,T]$,
(4.1) $\langle w_{g}(t),\phi\rangle=-\int_{0}^{t}\langle w_{g}(s),A\phi\rangle\,\mathrm{d}s-\langle g,\widehat{n}\cdot\nabla\phi\rangle_{H^{-s}(\Gamma_{u}),H^{s}(\Gamma_{u})}W^{\mathcal{H}}_{t}.$
Since $g$ is time-independent, the last term in (4.1) can be rewritten as a stochastic integral as
$\langle g,\widehat{n}\cdot\nabla\phi\rangle_{H^{-s}(\Gamma_{u}),H^{s}(\Gamma_{u})}W^{\mathcal{H}}_{t}=\int_{0}^{t}\langle g,\widehat{n}\cdot\nabla\phi\rangle_{H^{-s}(\Gamma_{u}),H^{s}(\Gamma_{u})}\,\mathrm{d}W^{\mathcal{H}}_{s}.$
Lemma 4.2.
Let 1.1 be satisfied. There exists a unique weak solution of (1.3) in the sense of 4.1 and it is given by the formula (3.1).
Proof.
We split the proof into two steps.
Step 1: There exists a unique weak solution of (1.3) and it is necessarily given by the mild formula (3.1). Let $\psi\in C^{1}([0,T];\mathsf{D}(A))$. Arguing as in the first step of the proof of [29, Theorem 1.7], see also [44, Lemma 3], one can readily check that $w_{g}$ satisfies
$\langle w_{g}(t),\psi(t)\rangle=$ $\int_{0}^{t}\langle w_{g}(s),\partial_{s}\psi(s)\rangle\,\mathrm{d}s-\int_{0}^{t}\langle w_{g}(s),A\psi(s)\rangle\,\mathrm{d}s$
(4.2) $-\int_{0}^{t}\langle g,\widehat{n}\cdot\nabla\psi(s)\rangle_{H^{-s}(\Gamma_{u}),H^{s}(\Gamma_{u})}\,\mathrm{d}W^{\mathcal{H}}_{s}$
for each $t\in[0,T],\ \mathbf{P}-a.s.$ The stochastic integral in the relation above is well-defined as a real-valued stochastic integral. Recalling that $(W^{\mathcal{H}}_{t})_{t\geq 0}$ is a $U$-cylindrical fractional Brownian motion we observe that $\langle g,\widehat{n}\cdot\nabla\psi(s)\rangle_{H^{-s}(\Gamma_{u}),H^{s}(\Gamma_{u})}$ is given by the linear operator on $U$
$h^{\prime}\mapsto\langle gh^{\prime},\widehat{n}\cdot\nabla\psi(s)\rangle_{H^{-s}(\Gamma_{u}),H^{s}(\Gamma_{u})}=L_{\psi}(gh^{\prime}),$
where $L_{\psi}:=\langle\cdot,\widehat{n}\cdot\nabla\psi(\cdot)\rangle_{H^{-s}(\Gamma_{u}),H^{s}(\Gamma_{u})}.$ By the ideal property of the Hilbert-Schmidt operators we have $\mathscr{L}_{2}(U,H^{-s}(\Gamma_{u}))=H^{-s}(\Gamma_{u};U)$ and obtain that
$\|\langle g,\widehat{n}\cdot\nabla\psi(s)\rangle_{H^{-s}(\Gamma_{u}),H^{s}(\Gamma_{u})}\|_{U^{*}}\lesssim\|g\|_{H^{-s}(\Gamma_{u})}\|\nabla\psi(s)\|_{H^{s}(\Gamma_{u})}~{}~{}\text{a.e. on }\Omega\times(0,T).$
In conclusion, the stochastic integral in (4.1) is well-defined as a real-valued one, since
$\mathbf{E}\Big{|}\int_{0}^{t}\langle g,\widehat{n}\cdot\nabla\psi(s)\rangle_{H^{-s}(\Gamma_{u}),H^{s}(\Gamma_{u})}~{}\mathrm{d}W^{\mathcal{H}}_{s}\Big{|}^{2}$
$\leq\mathcal{H}(2\mathcal{H}-1)\int_{0}^{t}\int_{0}^{t}\|\langle g,\widehat{n}\cdot\nabla\psi(s)\rangle\|_{U^{*}}\|\langle g,\widehat{n}\cdot\nabla\psi(v)\rangle\|_{U^{*}}|s-v|^{2\mathcal{H}-2}~{}\mathrm{d}s~{}\mathrm{d}v$
$\leq\mathcal{H}(2\mathcal{H}-1)\|g\|^{2}_{H^{-s}(\Gamma_{u};U)}\int_{0}^{t}\int_{0}^{t}\|\nabla\psi(s)\|_{H^{s}(\Gamma_{u})}\|\nabla\psi(v)\|_{H^{s}(\Gamma_{u})}|s-v|^{2\mathcal{H}-2}~{}\mathrm{d}s~{}\mathrm{d}v,$
which is finite since $\psi\in C^{1}([0,T];\mathsf{D}(A))$ and $\mathcal{H}>1/2$. Now consider $\phi\in\mathsf{D}(A^{2})$ and use $\psi_{t}(s)=S(t-s)\phi,\ s\in[0,t]$ as test function in (4.1) obtaining
(4.3) $\langle w_{g}(t),\phi\rangle=$ $-\int_{0}^{t}\langle g,\widehat{n}\cdot\nabla S(t-s)\phi\rangle_{H^{-s}(\Gamma_{u}),H^{s}(\Gamma_{u})}\,\mathrm{d}W^{\mathcal{H}}_{s}.$
Recalling the definition of the Dirichlet map $\mathcal{D}$, (4.3) can be rewritten as
(4.4) $\langle w(t),\phi\rangle=$ $\int_{0}^{t}\langle\mathcal{D}[g],AS(t-s)\phi\rangle\,\mathrm{d}W^{\mathcal{H}}_{s}.$
Then, exploiting the self-adjointness property of $S$ and $A$ we have that weak solutions of (1.3) satisfy the mild formulation. Therefore they are unique.
Step 2: The mild formula (3.1) is a weak solution of (1.3) in the sense of 4.1. We begin by noticing that $w_{g}$ has the required regularity due to 3.1. Let us test our mild formulation (3.1) against functions $\phi\in\mathsf{D}(A^{2})$. It holds, exploiting self-adjointness property of $S$ and $A$
$\langle w(t),\phi\rangle$ $=\int_{0}^{t}\langle\mathcal{D}[g],AS(t-s)\phi\rangle\,\mathrm{d}W^{\mathcal{H}}_{s}$
$=\int_{0}^{t}\langle g,\widehat{n}\cdot\nabla S(t-s)\phi\rangle_{H^{-s}(\Gamma_{u}),H^{s}(\Gamma_{u})}\,\mathrm{d}W^{\mathcal{H}}_{s}\quad\mathbf{P}-a.s.,$
where in the last step we used the definition of Dirichlet map. To complete the proof of this step it is enough to show that
$\int_{0}^{t}\langle g,\widehat{n}\cdot\nabla S(t-s)\phi\rangle_{H^{-s}(\Gamma_{u}),H^{s}(\Gamma_{u})}\,\mathrm{d}W^{\mathcal{H}}_{s}=-\int_{0}^{t}\langle w_{g}(s),A\phi\rangle\,\mathrm{d}s$
(4.5) $+\langle g,\widehat{n}\cdot\nabla\phi\rangle_{H^{-s}(\Gamma_{u}),H^{s}(\Gamma_{u})}\,W^{\mathcal{H}}_{t}\quad\mathbf{P}-a.s.$
The relation (4.1) is true. Indeed,
(4.6) $\int_{0}^{t}\langle w_{g}(s),A\phi\rangle\,\mathrm{d}s$ $=\int_{0}^{t}\,\mathrm{d}s\int_{0}^{s}\langle\mathcal{D}[g],S(s-\tau)A^{2}\phi\rangle\,\mathrm{d}W^{\mathcal{H}}(\tau)\quad\mathbf{P}-a.s.$
The double integrals in (4.6) can be exchanged via stochastic Fubini’s Theorem, see [48, 3]. Therefore the double integral in the right-hand side of (4.6) can be rewritten as
$\int_{0}^{t}\,\mathrm{d}s\int_{0}^{s}\langle\mathcal{D}[g],S(s-\tau)A^{2}\phi\rangle\,\mathrm{d}W^{\mathcal{H}}_{\tau}$
$\qquad=\int_{0}^{t}\,\mathrm{d}W^{\mathcal{H}}(\tau)\int_{\tau}^{t}\langle\mathcal{D}[g],S(s-\tau)A^{2}\phi\rangle\,\mathrm{d}s$
$\qquad=\langle\mathcal{D}[g],A\phi\rangle W^{\mathcal{H}}_{t}-\int_{0}^{t}\langle\mathcal{D}[g],AS(t-\tau)\phi\rangle\,\mathrm{d}W^{\mathcal{H}}_{\tau}$
$\qquad=\langle g,\widehat{n}\cdot\nabla\phi\rangle_{H^{-s}(\Gamma_{u}),H^{s}(\Gamma_{u})}\,W^{\mathcal{H}}_{t}$
$\qquad-\int_{0}^{t}\langle g,\widehat{n}\cdot\nabla S(t-\tau)\phi\rangle_{H^{-s}(\Gamma_{u}),H^{s}(\Gamma_{u})}\,\mathrm{d}W^{\mathcal{H}}_{\tau}\quad\mathbf{P}-a.s.$
Inserting this expression in (4.6), (4.1) holds and the proof is complete. ∎
Let $(v_{1},\dots,v_{N-1},\overline{v})$ be the $(p,r)$-solution to (3.7) as defined in 3.5 given by Theorem 3.7. Let $N_{0}$ be the $\mathbf{P}$ null measure set where at least one between
$w_{g}$ $\notin C([0,T];\mathbb{L}^{r}),\quad v_{i}\notin W^{1,p/2^{i}}(0,T;X_{-1/2,A_{q_{i}}})\cap L^{p/2^{i}}(0,T;X_{1/2,A_{q_{i}}}),$
$\overline{v}$ $\notin C([0,T];H)\cap L^{2}(0,T;V),$
(3.10), (3.11), (3.12), (4.1) is not satisfied. In the following, we will work pathwise in $\Omega\setminus N_{0}$ even if not specified. Thanks to the weak formulation guaranteed by 4.2 we can easily obtain the interior regularity of the linear stochastic problem (1.3). Indeed, we are exactly in the same position of [1, Corollary 4.4] and the following holds. We omit the proof as it follows verbatim the one of [1, Corollary 4.4].
Corollary 4.3.
Let 1.1 be satisfied. Let $w_{g}$ be the unique weak solution of (1.3) in the sense of 4.1. Then, for all $0<t_{1}\leq t_{2}<T,$ $x_{0}\in\mathcal{O}$, $\rho>0$ such that $\mathrm{dist}({B(x_{0},\rho)},\partial\mathcal{O})>0$,
$w_{g}\in C([t_{1},t_{2}],C^{\infty}({B(x_{0},\rho)};\mathbb{R}^{2}))\quad\mathbf{P}-a.s.$
4.2. Auxiliary Navier–Stokes equations and proof of Theorem 1.3(2)
To deal with the interior regularity of (3.4) we perform a Serrin type argument, see [41, 58]. In contrast to [1], as $w_{g}\notin C([0,T];\mathbb{L}^{4})$, we cannot work directly on $v$. However, recalling that the solution $v$ to (1.2) proven in subsection 3.3 satisfies
(4.7) $v=\sum_{i=0}^{N-1}v_{i}+\overline{v}$
where, again, $(v_{1},\dots,v_{N-1},\overline{v})$ is the $(p,r)$-solution to (3.7), cf., subsection 3.2. The advantage of having the splitting (4.7) at our disposal is that $v_{i}$ satisfies a linear problem where the forcing terms only depend on $v_{1},\dots,v_{i-1}$. Thus, by 4.3 and an induction argument, we can prove that $v_{i}$ is smooth inside $(0,T)\times\mathcal{O}$. While to prove the corresponding statement for $\overline{v}$, we can exploit that $\overline{v}$ is a Leray solution (i.e., it has finite energy) and therefore the Serrin regularization can be adjusted to our situation.
We begin with analyzing the interior regularity of $v_{i}$ for $i\in\{1,\dots,N-1\}$.
Lemma 4.4.
Let 1.1, $r\in(2,4)$ and $p\geq 2^{N}\frac{r}{r-2}$. Let $(v_{1},\dots,v_{N-1},\overline{v})$ be the $(p,r)$-solution of (3.7) in the sense of 3.5. Then for all $i\in\{0,\dots,N-1\},$ $0<t_{1}\leq t_{2}<T,$ $x_{0}\in\mathcal{O}$, $\rho>0$ such that $\mathrm{dist}({B(x_{0},\rho)},\partial\mathcal{O})>0$,
$v_{i}\in C([t_{1},t_{2}],C^{\infty}({B(x_{0},\rho)};\mathbb{R}^{2}))\quad\mathbf{P}-a.s.$
Proof.
As in the first step of the proof of Theorem 3.7 we argue by induction exploiting strongly the linear and triangle structure of (3.7). Before starting we observe that, by [53, Theorem 3.4.8], it follows that
(4.8) $v_{i}\in C([0,T];\mathbb{L}^{q_{i}})\cap L^{p/{2^{i}}}(0,T;X_{1/2,A_{q_{i}}}).$
[…]
$\langle v_{0}(t),\phi(t)\rangle$ $=\int_{0}^{t}\langle v_{0}(s),\partial_{s}\phi(s)\rangle\,\mathrm{d}s-\int_{0}^{t}\left\langle v_{0}\left(s\right),A_{q_{0}^{\prime}}\phi(s)\right\rangle\,\mathrm{d}s$
$+\int_{0}^{t}b\left(w_{g}\left(s\right),\phi(s),w_{g}\left(s\right)\right)\,\mathrm{d}s\quad\mathbf{P}-a.s.$
Choosing $\phi=-\nabla^{\perp}\chi,\ \chi\in C^{\infty}_{c}((0,T)\times\mathcal{O})$ in the weak formulation above and denoting by
$\omega_{0}=\operatorname{curl}v_{0}$ $\in C([0,T];H^{-1,q_{0}}(\mathcal{O}))\cap L^{p}(0,T;L^{q_{0}}(\mathcal{O})),$
$\omega_{w}=\operatorname{curl}w_{g}$ $\in C([t_{1}-2\varepsilon,t_{2}+2\varepsilon],C^{\infty}({B(x_{0},\rho+2\varepsilon)}))\quad\mathbf{P}-a.s.$
it follows that
$-\int_{0}^{t}\langle\omega_{0}(s),\partial_{s}\chi(s)\rangle+\langle\omega(s),\Delta\chi(s)\rangle\,\mathrm{d}s$ $=\int_{0}^{t}\langle\operatorname{curl}(w_{g}(s)\otimes w_{g}(s)),\nabla\chi(s)\rangle\,\mathrm{d}s.$
This means that $\omega_{0}$ is a distributional solution in $(0,T)\times\mathcal{O}$ of the partial differential equation
$\partial_{t}\omega_{0}$ $=\Delta\omega_{0}-\operatorname{div}\operatorname{curl}(w_{g}(s)\otimes w_{g}(s)).$
Let us consider $\psi_{0}\in C^{\infty}_{c}((0,T)\times\mathcal{O})$ supported in $[t_{1}-\varepsilon,t_{2}+\varepsilon]\times B(x_{0},\rho+\varepsilon)$ such that it is equal to one in $[t_{1}-\varepsilon/2,t_{2}+\varepsilon/2]\times B(x_{0},\rho+\varepsilon/2)$. Let us denote by ${\omega}^{*}_{0}=\omega_{0}\psi_{0}\in L^{p}(0,T;L^{q_{0}}(\mathbb{R}^{2}))$ supported in $[t_{1}-\varepsilon,t_{2}+\varepsilon]\times B(x_{0},\rho+\varepsilon)$, then ${\omega}^{*}_{0}$ is a distributional solution in $(0,T)\times\mathbb{R}^{2}$ of
(4.9) $\partial_{t}{\omega}^{*}_{0}$ $=\Delta{\omega}^{*}_{0}+h_{0}$
with
$h_{0}$ $=\partial_{t}\psi_{0}\omega_{0}-2\nabla\psi_{0}\cdot\nabla\omega_{0}-\Delta\psi_{0}\omega_{0}-\psi_{0}w_{g}\cdot\nabla\omega_{w}.$
Due to 4.3
$h_{0}\in L^{p}(0,T;H^{-1,q_{0}}(\mathbb{R}^{2}))\quad\mathbf{P}-a.s.$
Then, again by maximal $L^{p}$-regularity techniques for the heat equation (see e.g. [53, Theorem 4.4.4] [37, Theorems 10.2.25 and 10.3.4]) and the trace embedding of [53, Theorem 3.4.8],
${\omega}^{*}_{0}\in C([0,T];L^{q_{0}}(\mathbb{R}^{2}))\cap L^{p}(0,T;H^{1,q_{0}}(\mathbb{R}^{2})).$
Therefore,
$\omega_{0}\in$ $C([t_{1}-\varepsilon/4,t_{2}+\varepsilon/4],L^{q_{0}}(B(x_{0},\rho+\varepsilon/4)))$
$\cap L^{p}(t_{1}-\varepsilon/4,t_{2}+\varepsilon/4,H^{1,q_{0}}(B(x_{0},\rho+\varepsilon/4)))\quad\mathbf{P}-a.s.$
Introducing $\phi_{0}\in C^{\infty}_{c}(B(x_{0},\rho+\varepsilon/4))$ equal to one in $B(x_{0},\rho+\varepsilon/8)$, since $\omega_{0}=\operatorname{curl}v_{0}$, then $\phi_{0}v_{0}$ satisfies
(4.10) $\Delta(\phi_{0}v_{0})=\nabla^{\perp}\omega_{0}\phi_{0}+\Delta\phi_{0}v_{0}+2\nabla\phi_{0}\cdot\nabla v_{0},\quad(\phi_{0}v)|_{\partial B(x_{0},\rho+\varepsilon/4)}=0.$
From the regularity of $\omega_{0}$, by standard elliptic regularity theory (see for example [61, Chapter 4]), it follows that $\phi_{0}v_{0}\in C([t_{1}-\varepsilon/4,t_{2}+\varepsilon/4];H^{1,q_{0}}(B(x_{0},\rho+\varepsilon/4);\mathbb{R}^{2}))\cap L^{p}(t_{1}-\varepsilon/4,t_{2}+\varepsilon/4;H^{2,q_{0}}(B(x_{0},\rho+\varepsilon/4);\mathbb{R}^{2}))\ \mathbf{P}-a.s$. Therefore, since $\phi_{0}\equiv 1$ on $B(x_{0},\rho+\varepsilon/8)$
$v_{0}\in$ $C([t_{1}-{\varepsilon}/{16},t_{2}+{\varepsilon}/{16}];H^{1,q_{0}}(B(x_{0},\rho+{\varepsilon}/{16});\mathbb{R}^{2}))$
(4.11) $\cap L^{p}(t_{1}-{\varepsilon}/{16},t_{2}+{\varepsilon}/{16};H^{2,q_{0}}(B(x_{0},\rho+{\varepsilon}/{16});\mathbb{R}^{2}))\quad\mathbf{P}-a.s.$
Reiterating the argument, i.e. considering for each $j\in{\mathbb{N}},\ j\geq 0$, first $\psi_{j}\in C^{\infty}_{c}((0,T)\times\mathcal{O})$ supported in $[t_{1}-\varepsilon/2^{4j},t_{2}+\varepsilon/2^{4j}]\times B(x_{0},\rho+\varepsilon/2^{4j})$ identically equal to one in $[t_{1}-\varepsilon/2^{4j+1},t_{2}+\varepsilon/2^{4j+1}]\times B(x_{0},\rho+\varepsilon/2^{4j+1})$ and $\phi_{j}\in C^{\infty}_{c}(B(x_{0},\rho+\varepsilon/2^{4j+2}))$ identically equal to one in $B(x_{0},\rho+\varepsilon/2^{4j+3})$ we get iteratively that $\mathbf{P}-$a.s.
$\omega_{0}\in$ $C([t_{1}-\varepsilon/2^{4j+2},t_{2}+\varepsilon/2^{4j+2}],H^{j,q_{0}}(B(x_{0},\rho+\varepsilon/2^{4j+2})))$
$\cap L^{p}(t_{1}-\varepsilon/2^{4j+2},t_{2}+\varepsilon/2^{4j+2},H^{j+1,q_{0}}(B(x_{0},\rho+\varepsilon/2^{4j+2})))$
$v_{0}\in$ $C([t_{1}-\varepsilon/2^{4(j+1)},t_{2}+\varepsilon/2^{4(j+1)}],H^{j+1,q_{0}}(B(x_{0},\rho+\varepsilon/2^{4(j+1)});\mathbb{R}^{2}))$
$\cap L^{p}(t_{1}-\varepsilon/2^{4(j+1)},t_{2}+\varepsilon/2^{4(j+1)},H^{j+2,q_{0}}(B(x_{0},\rho+\varepsilon/2^{4(j+1)});\mathbb{R}^{2})).$
and the claimed interior regularity for $v_{0}$ follows.
[…]
$\langle v_{l}(t),\phi(t)\rangle$ $=\int_{0}^{t}\langle v_{l}(s),\partial_{s}\phi(s)\rangle\,\mathrm{d}s-\int_{0}^{t}\left\langle v_{0}\left(s\right),A_{q_{l}^{\prime}}\phi(s)\right\rangle\,\mathrm{d}s$
$+\int_{0}^{t}b(v_{l-1}(s),\phi(s),w(s)+\sum_{j=0}^{l-1}v_{j}(s))\,\mathrm{d}s$
$+\int_{0}^{t}b(w(s)+\sum_{j=0}^{i-2}v_{j}(s),\phi(s),v_{i-1}(s))\,\mathrm{d}s\quad\mathbf{P}-a.s.$
Choosing $\phi=-\nabla^{\perp}\chi,\ \chi\in C^{\infty}_{c}((0,T)\times\mathcal{O})$ in the weak formulation above and, for $i\in\{0,\dots,l-1\}$, denoting by
$\omega_{l}=\operatorname{curl}v_{l}$ $\in C([0,T];H^{-1,q_{l}}(\mathcal{O}))\cap L^{p/2^{l}}(0,T;L^{q_{l}}(\mathcal{O})),$
$\omega_{i}=\operatorname{curl}v_{i}$ $\in C([t_{1}-2\varepsilon,t_{2}+2\varepsilon],C^{\infty}({B(x_{0},\rho+2\varepsilon)})),$
$\omega_{w}=\operatorname{curl}w_{g}$ $\in C([t_{1}-2\varepsilon,t_{2}+2\varepsilon],C^{\infty}({B(x_{0},\rho+2\varepsilon)}))\quad\mathbf{P}-a.s.$
arguing as in Step 1 it follows that $\omega_{l}$ is a distributional solution in $(0,T)\times\mathcal{O}$ of the partial differential equation
$\partial_{t}\omega_{l}$ $=\Delta\omega_{l}-\operatorname{div}\operatorname{curl}(v_{l-1}(s)\otimes v_{l-1}(s))$
$-\operatorname{div}\operatorname{curl}\left(v_{l-1}(s)\otimes\left(w_{g}(s)+\sum_{j=0}^{l-2}v_{j}(s)\right)\right)$
$-\operatorname{div}\operatorname{curl}\left(\left(w_{g}(s)+\sum_{j=0}^{l-2}v_{j}(s)\right)\otimes v_{l-1}(s)\right).$
Let us consider $\psi_{0}\in C^{\infty}_{c}((0,T)\times\mathcal{O})$ supported in $[t_{1}-\varepsilon,t_{2}+\varepsilon]\times B(x_{0},\rho+\varepsilon)$ such that it is equal to one in $[t_{1}-\varepsilon/2,t_{2}+\varepsilon/2]\times B(x_{0},\rho+\varepsilon/2)$. Let us denote by ${\omega}^{*}_{l}=\omega_{l}\psi_{0}\in L^{p}(0,T;L^{q_{l}}(\mathbb{R}^{2}))$ supported in $[t_{1}-\varepsilon,t_{2}+\varepsilon]\times B(x_{0},\rho+\varepsilon)$, then ${\omega}^{*}_{l}$ is a distributional solution in $(0,T)\times\mathbb{R}^{2}$ of
(4.12) $\partial_{t}{\omega}^{*}_{l}$ $=\Delta{\omega}^{*}_{l}+h_{l}$
with
$h_{l}$ $=\partial_{t}\psi_{0}\omega_{l}-2\nabla\psi_{0}\cdot\nabla\omega_{l}-\Delta\psi_{0}\omega_{l}-\psi_{0}w_{l-1}\cdot\nabla\omega_{l-1}$
$-\psi_{0}w_{l-1}\cdot\nabla\left(\omega_{w}+\sum_{j=0}^{l-2}\omega_{j}\right)-\psi_{0}\left(w_{j}+\sum_{j=0}^{l-2}v_{j}\right)\cdot\nabla\omega_{l-1}.$
Due to 4.3 and the inductive hypothesis
$h_{l}\in L^{p}(0,T;H^{-1,q_{l}}(\mathbb{R}^{2}))\quad\mathbf{P}-a.s.$
Now we can argue as in Step 1 obtaining the claim. We omit the easy details. ∎
Now we are in the position to apply similar ideas of [1, Section 4.2] for the equation satisfied by $\overline{v}$. For the sake of completeness, we provide some details.
Lemma 4.5.
Let 1.1, $r\in(2,4)$ and $p\geq 2^{N}\frac{r}{r-2}$. Let $(v_{1},\dots,v_{N-1},\overline{v})$ be the $(p,r)$-solution of (3.7) in the sense of 3.5. Then, for all $0<t_{1}\leq t_{2}<T,$ $x_{0}\in\mathcal{O}$, $\rho>0$ such that $\mathrm{dist}({B(x_{0},\rho)},\partial\mathcal{O})>0$,
$\overline{v}\in C([t_{1},t_{2}],H^{3/2}({B(x_{0},\rho)};\mathbb{R}^{2}))\quad\mathbf{P}-a.s.$
Proof.
[…]
$\tilde{v}$ $=w+\sum_{j=0}^{N-1}v_{j},\ \widetilde{\omega}=\operatorname{curl}\widetilde{v}.$
As described in 4.2, arguing as in the proof of [29, Theorem 7], we can extend the weak formulation satisfied by $\overline{v}$ to time-dependent test functions $\phi\in C^{1}([0,T];H)\cap C([0,T];\mathsf{D}(A))$ obtaining that for each $t\in[0,T]$
$\langle\overline{v}(t),\phi(t)\rangle-\langle u_{\mathrm{in}},\phi(0)\rangle$ $=\int_{0}^{t}\langle\overline{v}(s),\partial_{s}\phi(s)\rangle\,\mathrm{d}s-\int_{0}^{t}\left\langle\overline{v}\left(s\right),A\phi(s)\right\rangle\,\mathrm{d}s$
$+\int_{0}^{t}b\left(\overline{v}\left(s\right)+\widetilde{v}\left(s\right),\phi(s),\overline{v}\left(s\right)\right)\,\mathrm{d}s$
$+\int_{0}^{t}b\left(\overline{v}\left(s\right),\phi(s),\widetilde{v}\left(s\right)\right)\,\mathrm{d}s$
$+\int_{0}^{t}b\left(v_{N-1}\left(s\right),\phi(s),\widetilde{v}(s)\right)\,\mathrm{d}s$
$+\int_{0}^{t}b\left(\widetilde{v}\left(s\right)-v_{N-1}\left(s\right),\phi(s),v_{N-1}\left(s\right)\right)\,\mathrm{d}s\quad\mathbf{P}-a.s.$
Choosing $\phi=-\nabla^{\perp}\chi,\ \chi\in C^{\infty}_{c}((0,T)\times\mathcal{O})$ in the weak formulation above and, for $i\in\{0,\dots,N-1\},$ denoting by
$\omega=\operatorname{curl}v$ $\in C([0,T];H^{-1})\cap L^{2}((0,T)\times\mathcal{O}),$
$\omega_{i}=\operatorname{curl}v_{i}$ $\in C([t_{1}-2\varepsilon,t_{2}+2\varepsilon],C^{\infty}({B(x_{0},\rho+2\varepsilon)})),$
$\omega_{w}=\operatorname{curl}w$ $\in C([t_{1}-2\varepsilon,t_{2}+2\varepsilon],C^{\infty}({B(x_{0},\rho+2\varepsilon)}))\quad\mathbf{P}-a.s.$
it follows that
$-\int_{0}^{t}\langle\omega(s),\partial_{s}\chi(s)\rangle+\langle\omega(s),\Delta\chi(s)\rangle\,\mathrm{d}s$ $=\int_{0}^{t}\langle\operatorname{curl}(v_{N-1}(s)\otimes\widetilde{v}(s)),\nabla\chi(s)\rangle\,\mathrm{d}s$
$+\int_{0}^{t}\langle\operatorname{curl}(\left(\widetilde{v}(s)-v_{N-1}(s)\right)\otimes v_{N-1}(s)),\nabla\chi(s)\rangle\,\mathrm{d}s$
$+\int_{0}^{t}\langle\operatorname{curl}(\overline{v}(s)\otimes\widetilde{v}(s)),\nabla\chi(s)\rangle\,\mathrm{d}s$
$+\int_{0}^{t}\langle\operatorname{curl}(\widetilde{v}(s)\otimes\overline{v}(s)),\nabla\chi(s)\rangle\,\mathrm{d}s$
$+\int_{0}^{t}\langle\omega(s),\overline{v}(s)\cdot\nabla\chi(s)\rangle\,\mathrm{d}s.$
This means that $\omega$ is a distributional solution in $(0,T)\times\mathcal{O}$ of the partial differential equation
$\partial_{t}\omega+\overline{v}\cdot\nabla\omega$ $=\Delta\omega-\operatorname{div}\bigg{(}\operatorname{curl}(v_{N-1}(s)\otimes\widetilde{v}(s))$
$+\operatorname{curl}(\widetilde{v}(s)-v_{N-1}(s)\otimes v_{N-1}(s))$
$+\operatorname{curl}(\widetilde{v}(s)\otimes\overline{v}(s))+\operatorname{curl}(\overline{v}(s)\otimes\widetilde{v}(s))\bigg{)}.$
Let us consider $\psi\in C^{\infty}_{c}((0,T)\times\mathcal{O})$ supported in $[t_{1}-\varepsilon,t_{2}+\varepsilon]\times B(x_{0},\rho+\varepsilon)$ such that it is equal to one in $[t_{1}-\varepsilon/2,t_{2}+\varepsilon/2]\times B(x_{0},\rho+\varepsilon/2)$. Let us denote by ${\omega}^{*}=\omega\psi\in L^{2}((0,T)\times\mathbb{R}^{2})$ supported in $[t_{1}-\varepsilon,t_{2}+\varepsilon]\times B(x_{0},\rho+\varepsilon)$, then ${\omega}^{*}$ is a distributional solution in $(0,T)\times\mathbb{R}^{2}$ of
(4.13) $\partial_{t}{\omega}^{*}$ $=\Delta{\omega}^{*}-\overline{v}\cdot\nabla{\omega}^{*}-\widetilde{v}\cdot\nabla{\omega}^{*}+h$
with
$h$ $=\partial_{t}\psi\omega-2\nabla\psi\cdot\nabla\omega-\Delta\psi\omega+\overline{v}\cdot\nabla\psi\omega+\widetilde{v}\cdot\nabla\psi\omega-\psi\left(\widetilde{v}-v_{N-1}\right)\cdot\nabla\omega_{N-1}$
$-\psi\overline{v}\cdot\nabla\widetilde{\omega}-\psi v_{N-1}\cdot\nabla\widetilde{\omega}.$
Due to 4.3 and 4.4 the terms
$\widetilde{v}\cdot\nabla\psi\omega-\psi\left(\widetilde{v}-v_{N-1}\right)\cdot\nabla\omega_{N-1}-\psi\overline{v}\cdot\nabla\widetilde{\omega}$
$-\psi v_{N-1}\cdot\nabla\widetilde{\omega}\in L^{2}((0,T)\times\mathbb{R}^{2})\quad\mathbf{P}-a.s.$
Therefore $h\in L^{2}(0,T;H^{-1}(\mathbb{R}^{2}))+L^{1}(0,T;L^{2}(\mathbb{R}^{2}))\ \mathbf{P}-a.s.$ Then, arguing as in the first step of the proof of [41, Theorem 13.2], the fact that ${\omega}^{*}$ is a distributional solution of (4.13) implies that ${\omega}^{*}\in C([0,T];L^{2}(\mathbb{R}^{2}))\cap L^{2}(0,T;H^{1}(\mathbb{R}^{2})).$ Therefore
$\omega\in$ $C([t_{1}-\varepsilon/4,t_{2}+\varepsilon/4],L^{2}(B(x_{0},\rho+\varepsilon/4)))$
$\cap L^{2}(t_{1}-\varepsilon/4,t_{2}+\varepsilon/4,H^{1}(B(x_{0},\rho+\varepsilon/4)))\quad\mathbf{P}-a.s.$
Introducing $\phi\in C^{\infty}_{c}(B(x_{0},\rho+\varepsilon/4))$ equal to one in $B(x_{0},\rho+\varepsilon/8)$, since $\omega=\operatorname{curl}\overline{v}$, then $\phi\overline{v}$ satisfies
(4.14) $\Delta(\phi\overline{v})=\nabla^{\perp}\omega\phi+\Delta\phi\overline{v}+2\nabla\phi\cdot\nabla\overline{v},\quad(\phi\overline{v})|_{\partial B(x_{0},\rho+\varepsilon/4)}=0.$
From the regularity of $\omega$, by standard elliptic regularity theory (see for example [5]), it follows that $\phi\overline{v}\in C([t_{1}-\varepsilon/4,t_{2}+\varepsilon/4];H^{1}(B(x_{0},\rho+\varepsilon/4);\mathbb{R}^{2}))\cap L^{2}(t_{1}-\varepsilon/4,t_{2}+\varepsilon/4;H^{2}(B(x_{0},\rho+\varepsilon/4);\mathbb{R}^{2}))\ \mathbf{P}-a.s$. Therefore, since $\phi\equiv 1$ on $B(x_{0},\rho+\varepsilon/8)$
$\overline{v}\in$ $C([t_{1}-{\varepsilon}/{16},t_{2}+{\varepsilon}/{16}];H^{1}(B(x_{0},\rho+{\varepsilon}/{16});\mathbb{R}^{2}))$
(4.15) $\cap L^{2}(t_{1}-{\varepsilon}/{16},t_{2}+{\varepsilon}/{16};H^{2}(B(x_{0},\rho+{\varepsilon}/{16});\mathbb{R}^{2}))\quad\mathbf{P}-a.s.$
Let us now consider $\widehat{\psi}\in C^{\infty}_{c}((t_{1}-{\varepsilon}/{16},t_{2}+{\varepsilon}/{16})\times B(x_{0},\rho+{\varepsilon}/{16}))$ such that it is equal to one in $[t_{1}-{\varepsilon}/{32},t_{2}+{\varepsilon}/{32}]\times B(x_{0},\rho+{\varepsilon}/{32})$. Let us denote by $\widehat{\omega}=\omega\widehat{\psi}\in C([0,T];L^{2}(\mathbb{R}^{2}))\cap L^{2}(0,T;H^{1}(\mathbb{R}^{2}))$ supported in $(t_{1}-{\varepsilon}/{16},t_{2}+{\varepsilon}/{16})\times B(x_{0},\rho+{\varepsilon}/{16})$, then $\widehat{\omega}$ is a distributional solution in $(0,T)\times\mathbb{R}^{2}$ of
(4.16) $\partial_{t}\widehat{\omega}$ $=\Delta\widehat{\omega}+\widehat{h}$
with
$\widehat{h}$ $=-\overline{v}\cdot\nabla\widehat{\omega}-\widetilde{v}\cdot\nabla\widehat{\omega}+\partial_{t}\widehat{\psi}\omega-2\nabla\widehat{\psi}\cdot\nabla\omega-\Delta\widehat{\psi}\omega+\overline{v}\cdot\nabla\widehat{\psi}\omega+\widetilde{v}\cdot\nabla\widehat{\psi}\omega$
$-\widehat{\psi}\left(\widetilde{v}-v_{N-1}\right)\cdot\nabla\omega_{N-1}-\widehat{\psi}\overline{v}\cdot\nabla\widetilde{\omega}-\widehat{\psi}v_{N-1}\cdot\nabla\widetilde{\omega}.$
By 4.3, 4.4 and relation (4.2) it follows that
$\widehat{h}\in L^{2}(0,T;H^{-1/2}(\mathbb{R}^{2}))\ \mathbf{P}-a.s.$
Therefore $\widehat{\omega}\in C([0,T];H^{1/2}(\mathbb{R}^{2}))\cap L^{2}(0,T;H^{3/2}(\mathbb{R}^{2}))\ \mathbf{P}-a.s.$ and arguing as above
$\overline{v}\in$ $C([t_{1}-{\varepsilon}/{64},{t}_{2}+{\varepsilon}/{64}],H^{3/2}({B(x_{0},{r}+{\varepsilon}/{64})};\mathbb{R}^{2}))$
$\cap L^{2}({t}_{1}-{\varepsilon}/{64},{t}_{2}+{\varepsilon}/{64},H^{5/2}({B(x_{0},{\rho}+{\varepsilon}/{64})};\mathbb{R}^{2}))\quad\mathbf{P}-a.s.$
This concludes the proof of 4.5. ∎
Corollary 4.6.
Let 1.1, $r\in(2,4)$ and $p\geq 2^{N}\frac{r}{r-2}$. Let $(v_{1},\dots,v_{N-1},\overline{v})$ be the $(p,r)$-solution of (3.7) in the sense of 3.5. Then, for all $0<t_{1}\leq t_{2}<T,$ $x_{0}\in\mathcal{O}$, $\rho>0$ such that $\mathrm{dist}({B(x_{0},\rho)},\partial\mathcal{O})>0$,
$\overline{v}\in C([t_{1},t_{2}];C^{\infty}({B(x_{0},\rho)};\mathbb{R}^{2}))\quad\mathbf{P}-a.s.$
Proof.
[…]
$\tilde{v}$ $=w+\sum_{j=0}^{N-1}v_{j},\ \widetilde{\omega}=\operatorname{curl}\widetilde{v},$
$\omega=\operatorname{curl}v$ $\in C([0,T];H^{-1})\cap L^{2}((0,T)\times\mathcal{O}),$
$\omega_{i}=\operatorname{curl}v_{i}$ $\in C([t_{1}-2\varepsilon,t_{2}+2\varepsilon],C^{\infty}({B(x_{0},\rho+2\varepsilon)})),$
$\omega_{w}=\operatorname{curl}w$ $\in C([t_{1}-2\varepsilon,t_{2}+2\varepsilon],C^{\infty}({B(x_{0},\rho+2\varepsilon)}))\quad\mathbf{P}-a.s.$
and $\omega^{*}=\omega\psi\in L^{2}((0,T)\times\mathbb{R}^{2})$ supported in $[t_{1}-\varepsilon,t_{2}+\varepsilon]\times B(x_{0},\rho+\varepsilon)$, then, arguing as in the proof of 4.5, it follows that $\omega^{*}$ is a distributional solution in $(0,T)\times B(x_{0},\rho+\varepsilon)$ of
(4.17) $\partial_{t}\omega^{*}$ $=\Delta\omega^{*}+\widetilde{h}$
with
$\widetilde{h}$ $=-\overline{v}\cdot\nabla{\omega}^{*}-\widetilde{v}\cdot\nabla{\omega}^{*}+\partial_{t}{\psi}\omega-2\nabla{\psi}\cdot\nabla\omega-\Delta{\psi}\omega+\overline{v}\cdot\nabla{\psi}\omega+\widetilde{v}\cdot\nabla{\psi}\omega$
$-{\psi}\left(\widetilde{v}-v_{N-1}\right)\cdot\nabla\omega_{N-1}-{\psi}\overline{v}\cdot\nabla\widetilde{\omega}-{\psi}v_{N-1}\cdot\nabla\widetilde{\omega}.$
From the regularity of $\omega,\ \overline{v},\ \widetilde{\omega},\ \widetilde{v},\ \omega_{N-1},v_{N-1}$, then $\widetilde{h}\in L^{2}(t_{1}-\varepsilon,t_{2}+\varepsilon;H^{-1}(B(x_{0},\rho+\varepsilon)))\ \mathbf{P}-a.s$. By standard regularity theory for the heat equation, see for example Step 2 in [41, Theorem 13.1], a solution of (4.17) with $\widetilde{h}\in L^{2}(t_{1}-\varepsilon,t_{2}+\varepsilon;H^{k-1}(B(x_{0},\rho+\varepsilon))),\ k\in\mathbb{N}$, belongs to $C([{t}_{1}-\varepsilon/2,{t}_{2}+\varepsilon/2];H^{k}(B(x_{0},{\rho}+\varepsilon/2)))\cap L^{2}({t}_{1}-\varepsilon/2,{t}_{2}+\varepsilon/2;H^{k+1}(B(x_{0},{\rho}+\varepsilon/2)))$. Therefore
${\omega}^{*}$ $\in C([{t}_{1}-\varepsilon/2,{t}_{2}+\varepsilon/2];L^{2}(B(x_{0},{\rho}+\varepsilon/2)))$
$\cap L^{2}({t}_{1}-\varepsilon/2,{t}_{2}+\varepsilon/2;H^{1}(B(x_{0},{\rho}+\varepsilon/2)))\quad\mathbf{P}-a.s.$
which implies
$\omega$ $\in C([{t}_{1}+\varepsilon/4,{t}_{2}-\varepsilon/4;L^{2}(B(x_{0},{\rho}+\varepsilon/4)))$
$\cap L^{2}({t}_{1}-\varepsilon/4,{t}_{2}+\varepsilon/4;H^{1}(B(x_{0},{\rho}+\varepsilon/4)))\quad\mathbf{P}-a.s.$
since $\psi\equiv 1$ on $({t}_{1}-\varepsilon/2,{t}_{2}+\varepsilon/2)\times B(x_{0},{\rho}+\varepsilon/2).$ Considering now $\phi\in C^{\infty}_{c}(\mathcal{O})$ supported on $B(x_{0},\rho+\varepsilon/4)$ such that $\phi\equiv 1$ on $B(x_{0},{\rho}+\varepsilon/8)$, since $\operatorname{curl}\overline{v}=\omega$ then $\phi\overline{v}$ satisfies
(4.18) $\Delta(\phi\overline{v})=\nabla^{\perp}\omega\phi+\Delta\phi\overline{v}+2\nabla\phi\cdot\nabla\overline{v},\quad(\phi\overline{v})|_{\partial B(x_{0},\rho+\varepsilon/4)}=0.$
Since
$\nabla^{\perp}\omega\phi+\Delta\phi\overline{v}+2\nabla\phi\cdot\nabla\overline{v}$ $\in C([{t}_{1}+\varepsilon/4,{t}_{2}-\varepsilon/4;H^{-1}(B(x_{0},{\rho}+\varepsilon/4);\mathbb{R}^{2}))$
$\cap L^{2}({t}_{1}-\varepsilon/4,{t}_{2}+\varepsilon/4;L^{2}(B(x_{0},{\rho}+\varepsilon/4);\mathbb{R}^{2}))\quad\mathbf{P}-a.s.,$
by standard elliptic regularity theory (see for example [5]),
$\phi\overline{v}$ $\in C([{t}_{1}+\varepsilon/4,{t}_{2}-\varepsilon/4;H^{1}(B(x_{0},{\rho}+\varepsilon/4);\mathbb{R}^{2}))$
$\cap L^{2}({t}_{1}-\varepsilon/4,{t}_{2}+\varepsilon/4;H^{2}(B(x_{0},{\rho}+\varepsilon/4);\mathbb{R}^{2}))\quad\mathbf{P}-a.s..$
Since $\phi\equiv 1$ on $B(x_{0},{\rho}+\varepsilon/8)$ then
$\overline{v}$ $\in C([{t}_{1}+{\varepsilon}/{16},{t}_{2}-{\varepsilon}/{16};H^{1}(B(x_{0},\rho+{\varepsilon}/{16})))$
$\cap L^{2}({t}_{1}-{\varepsilon}/{16},{t}_{2}+{\varepsilon}/{16};H^{2}(B(x_{0},{\rho}+{\varepsilon}/{16})))\quad\mathbf{P}-a.s.$
Reiterating the argument as in Step 3 in [41, Theorem 13.1] the claim follows. ∎
Proof of Theorem 1.3(2).
The claim follows by 4.3, 4.4 and 4.6 and a localization argument. To begin, recall from the proof of Theorem 1.3(1) in subsection 3.3 that there exists a solution (1.2) on the time interval $[0,T+1]$ and it is given by $\widetilde{u}=w_{g}+\sum_{i=0}^{N-1}v_{i}+\overline{v}$ where $(v_{0},\dots,v_{N-1},\overline{v})$ is the $(p,r)$-solution to (3.7) on $[0,T+1]$ for $r<2q_{\mathcal{H}}$, $N$ as in (3.5) and $p\geq 2^{N}\frac{r}{r-2}$. Then, by 4.3, 4.4, 4.6 and a standard covering argument, for all $t_{0}\in(0,T)$, $\mathcal{O}_{0}\subset\mathcal{O}$ such that $\mathrm{dist}({\mathcal{O}_{0}},\partial\mathcal{O})>0$,
(4.19) $\widetilde{u}\in C([t_{0},T];C^{\infty}(\mathcal{O}_{0};\mathbb{R}^{2}))\quad\mathbf{P}-a.s.$
Now, let $u$ be the unique solution (1.2) provided by Theorem 1.3(1) on $[0,T]$. By uniqueness, we have $u=\widetilde{u}|_{[0,T]}$ and the conclusion follows from (4.19). ∎

### v2

@@
 Acknowledgements.
-The authors thank Robert Denk and Tim Seitz for several useful discussions.
+The authors thank Robert Denk and Tim Seitz for several useful discussions. Finally, the authors thank the anonymous referees for helpful comments which improved the paper from its initial version.
 Key words and phrases:
@@
 […]
+A.B. acknowledges support from DFG CRC/TRR 388, project A06.
 1. Introduction
@@
 In particular, in the case of Dirichlet boundary conditions the solution is only a distribution.
-This allowed us to treat only a restricted class of nonlinearities, exploiting specific properties of the heat semigroup and studying carefully the blow-up of the solution close to the boundary.
+This allowed to treat only a restricted class of nonlinearities, exploiting specific properties of the heat semigroup and studying carefully the blow-up of the solution close to the boundary.
 For some results in this direction, the reader is referred to [2, 12, 25, 36].
-On the contrary, in the last few years, maximal $L^{p}$ regularity techniques provided new ideas to treat partial differential equations with white noise Neumann boundary conditions with more severe nonlinearities such as those coming from fluid dynamical models.
-Indeed, some results on the global and local well-posedness of the 2D Navier-Stokes equations and the 3D primitive equations with boundary noise perturbations of Neumann time have been proven in [1] and [10], respectively.
+On the contrary, partial differential equations with white noise Neumann boundary conditions with more severe nonlinearities have been considered in [14, 61], and in the last few years, maximal $L^{p}$ regularity techniques provided new ideas to treat some of those arising in fluid dynamics.
+Indeed, some results on the global and local well-posedness of the 2D Navier-Stokes equations and the 3D primitive equations with boundary noise perturbations of Neumann type have been proven in [1] and [10], respectively.
 Besides the physical interests in studying the Navier-Stokes equations with boundary noise in virtue of its connection with the Couette flow (see also below for further motivations), the present manuscript also aims at (partially) filling the gap in the literature between Dirichlet and Neumann type boundary conditions for fluid dynamical models.
-Throughout the manuscript, we fix a finite time horizon $T>0$ and we let $\mathcal{O}=\mathbb{T}\times(0,a)$ where $\mathbb{T}$ is the one-dimensional torus and $a>0$. Moreover, we denote by
-(1.1) $\Gamma_{b}=\mathbb{T}\times\{0\}\quad\text{ and }\quad\Gamma_{u}=\mathbb{T}\times\{a\},$
-the bottom and the upper part of the boundary of $\mathcal{O}$, respectively.
-In this paper we are interested in the global well-posedness and the interior regularity of the 2D Navier-Stokes equations with boundary noise for the unknown velocity field $u(t,\omega,x,z)=(u_{1},u_{2}):(0,T)\times\Omega\times\mathcal{O}\to{\mathbb{R}}^{2}$ and pressure $P:(0,T)\times\Omega\times\mathcal{O}\to{\mathbb{R}}$, formally written as
+Throughout the manuscript, we fix a finite time horizon $T>0$ and consider a spatial domain $\mathcal{O}=\mathbb{T}\times(0,a)$ where $\mathbb{T}$ is the one-dimensional torus and $a>0$.
+We further define the lower and upper parts of the boundary of $\mathcal{O}$ by
+(1.1) $\Gamma_{b}=\mathbb{T}\times\{0\}\quad\text{ and }\quad\Gamma_{u}=\mathbb{T}\times\{a\}.$
+In this work, we focus on the global well-posedness and interior regularity of the two-dimensional Navier–Stokes equations with fractional boundary noise.
+The unknowns are the velocity field $u(t,\omega,x,z)=(u_{1},u_{2}):(0,T)\times\Omega\times\mathcal{O}\to{\mathbb{R}}^{2}$ and the pressure $P:(0,T)\times\Omega\times\mathcal{O}\to{\mathbb{R}}$, which formally satisfy the system
 (1.2) $\partial_{t}u&=\Delta u+\nabla P-(u\cdot\nabla)u\quad&\text{ on }&(0,T)\times\mathcal{O},\\ \operatorname{div}u&=0\quad&\text{ on }&(0,T)\times\mathcal{O},\\ u_{1}&=g\,\dot{W}^{\mathcal{H}}\quad&\text{ on }&(0,T)\times\Gamma_{u},\\ u_{2}&=0\quad&\text{ on }&(0,T)\times\Gamma_{u},\\ u&=0\quad&\text{ on }&(0,T)\times\Gamma_{b},\\ u(0)&=u_{\mathrm{in}}\quad&\text{ on }&\mathcal{O},\end{aligned}\right.$
@@
 According to [35, 52, 53], see also the discussion in the introduction of [1], the geometry considered in (1.2) can be seen as an idealization of the ocean dynamics (more precisely, a vertical slice of the ocean).
-The model (1.2), describes a Couette flow, namely a viscous fluid in the space between two surfaces, one of which is moving tangentially relative to the other.
+The model (1.2), describes a Couette flow, namely a viscous fluid in the space between two surfaces, one of which is moving tangentially relative to the other (see also Remark 1.4).
 The relative motion of the surfaces imposes a shear stress on the fluid and induces the flow.
 Let us recall that the onset of turbulence is often related to the randomness of background movement [47].
-Moreover, according to [52, Chapter 3] in any turbulent flow there are unavoidably perturbations in boundary conditions and material properties.
+Moreover, according to [54, Chapter 3] in any turbulent flow there are unavoidably perturbations in boundary conditions and material properties.
 We model these features by the noise term $g\,\dot{W}^{\mathcal{H}}$.
@@
 1.1. Main result
-We begin by introducing some notation.
-Throughout this manuscript, we consider a complete filtered probability space $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\geq 0},\mathbf{P})$, a separable Hilbert space $U$.
-We say that a process $\Phi$ is $\mathcal{F}$-progressive measurable if $\Phi|_{(0,t)\times\Omega}$ is $\mathcal{F}_{t}\times\mathcal{B}((0,t))$-measurable for all $t>0$, where $\mathcal{B}$ denotes the Borel $\sigma$-algebra.
-We postpone to subsection 1.3 the relevant notation on function spaces.
+We begin by introducing some basic notation.
+Throughout this manuscript, we work on a complete filtered probability space $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\geq 0},\mathbf{P})$ and consider a separable Hilbert space $U$.
+A process $\Phi$ is said to be $\mathcal{F}$-progressive measurable if, for every $t>0$, the restriction $\Phi|_{(0,t)\times\Omega}$ is measurable with respect to $\mathcal{F}_{t}\otimes\mathcal{B}((0,t))$, where $\mathcal{B}$ denotes the Borel $\sigma$-algebra.
+Further notation concerning function spaces is deferred to Section 1.3.
 On the noise $W^{\mathcal{H}}$ we enforce the following
@@
 Note that 1.1 is consistent with the results obtained in [23] for the stochastic heat equation with Dirichlet fractional noise.
-The reader is referred to 1.4 for the case of a time-dependent $g$.
-Following [21, Chapter 15] and [19], we look for solutions to (1.2) in the form
-$u=w_{g}+v,$
+The reader is referred to Remark 1.4 for the case of a time-dependent $g$.
+Following [22, Chapter 15] and [20], we construct solutions to (1.2) by the splitting
+(1.3) $u=w_{g}+v,$
 where $w_{g}$ is a mild solution of the linear problem with non-homogeneous boundary conditions
@@
 In (1.5), due to the divergence-free of $v$ and $w_{g}$, we rewrote the Navier-Stokes nonlinearity in the conservative form to accommodate the weak (PDE) setting.
-As discussed in [21, Chapter 13], if $g$, $u_{\mathrm{in}}$, $W^{\mathcal{H}}(t)$ were regular enough, then $u=v+w_{g}$ would be a classical solution of the Navier-Stokes equations with non-homogeneous boundary conditions (1.2).
+As discussed in [22, Chapter 13], if $g$, $u_{\mathrm{in}}$, $W^{\mathcal{H}}(t)$ were sufficiently regular, then $u=v+w_{g}$ would be a classical solution of the Navier-Stokes equations with non-homogeneous boundary conditions (1.2).
 Next, we introduce the class of solutions we are going to consider.
@@
 This has drastic consequences in our analysis.
-As we will show in 3.1, we have, $\mathbf{P}-a.s.$,
+As we will show in Proposition 3.1, we have, $\mathbf{P}-a.s.$,
 (1.6) $w_{g}\in C([0,T];L^{2q}(\mathcal{O};{\mathbb{R}}^{2}))\ \text{ for all }q\in(1,q_{\mathcal{H}})$
@@
 (1.7) $q_{\mathcal{H}}:=\frac{2}{2s+5-4\mathcal{H}}\in(1,2).$
-Let us stress that $q_{\mathcal{H}}<2$ and $\lim_{\mathcal{H}\downarrow 3/4}q_{\mathcal{H}}\leq 1$ even if $s=0$.
+Let us stress that $q_{\mathcal{H}}<2$ and $\lim_{\mathcal{H}\downarrow 3/4}q_{\mathcal{H}}{=}1$ even if $s=0$.
 As we will see below, this fact creates major difficulties in our analysis of the auxiliary Navier-Stokes equations (1.5).
@@
 For this reason, our proofs rely on different methods.
-For details, the reader is referred to the text before 1.4.
+For details, the reader is referred to the text before Remark 1.4.
 In light of the previous discussion, we are now ready to define solutions to (1.2).
-Below, we set $A:B=\sum_{i,j=1}^{2}A^{i,j}B^{i,j}$ for two matrices $A$ and $B$ and $\mathbb{L}^{q}$ the image of $L^{q}(\mathcal{O};{\mathbb{R}}^{2})$ via the Leray projection $\mathbb{P}$ defined rigorously in subsection 2.1.
+Below, we set $A:B=\sum_{i,j=1}^{2}A^{i,j}B^{i,j}$ for two matrices $A$ and $B$ and $\mathbb{L}^{q}$ the image of $L^{q}(\mathcal{O};{\mathbb{R}}^{2})$ via the Leray projection $\mathbb{P}$ defined rigorously in Section 2.1.
 Definition 1.2.
 Let $T<\infty$, $u_{\mathrm{in}}\in L^{0}_{\mathscr{F}_{0}}(\Omega;\mathbb{L}^{2})$ and $q\in(1,q_{\mathcal{H}})$.
-  • ($q$-solution) A progressively measurable process $u$ with $\mathbf{P}-a.s.$ paths in $L^{2q^{\prime}}(0,T;\mathbb{L}^{2q})$, is a pathwise weak $q$-solution of (1.2) if $u=v+w_{g}$ and $v$ satisfies for all divergence-free $\varphi\in C^{\infty}(\mathcal{O};{\mathbb{R}}^{2})$ such that $\varphi=0$ on $\Gamma_{b}\cup\Gamma_{u}$ and a.e.
+  • ($q$-solution) A progressively measurable process $u$ with $\mathbf{P}-a.s.$ paths in $L^{2q^{\prime}}(0,T;\mathbb{L}^{2q})$, is a pathwise weak $q$-solution of (1.2) if for all divergence-free $\varphi=(\varphi_{1},\varphi_{2})\in C^{\infty}(\mathcal{O};{\mathbb{R}}^{2})$ such that $\varphi=0$ on $\Gamma_{b}\cup\Gamma_{u}$ and a.e.
 $t\in(0,T)$,
-    $\int_{\mathcal{O}}v(x,t)\varphi(x)\,\mathrm{d}x-\int_{\mathcal{O}}u_{\mathrm{in}}(x)\varphi(x)\,\mathrm{d}x$
-    $\qquad\qquad=\int_{0}^{t}\int_{\mathcal{O}}\big{(}v\cdot\Delta\varphi+[(v+w_{g})\otimes(v+w_{g})]:\nabla\varphi\big{)}\,\mathrm{d}x\,\mathrm{d}s.$
-  • (unique $q$-solution) A $q$-solution $u$ to (1.2) is said to be a unique if for any other $q$-solution $\widetilde{u}$ we have $u=\widetilde{u}$ a.e. on $[0,T]\times\Omega$.
+    $\int_{\mathcal{O}}u(x,t)\varphi(x)\,\mathrm{d}x-\int_{\mathcal{O}}u_{\mathrm{in}}(x)\varphi(x)\,\mathrm{d}x$
+    $\qquad\qquad=\int_{0}^{t}\int_{\mathcal{O}}\big{(}u\cdot\Delta\varphi+[u\otimes u]:\nabla\varphi\big{)}\,\mathrm{d}x\,\mathrm{d}r-{\langle g,\partial_{2}\varphi_{1}\rangle_{H^{-s}(\Gamma_{u}),H^{s}(\Gamma_{u})}W^{\mathcal{H}}_{t}}.$
+  • (unique $q$-solution) A $q$-solution $u$ to (1.2) is said to be a unique solution if for any other $q$-solution $\widetilde{u}$ we have $u=\widetilde{u}$ a.e. on $[0,T]\times\Omega$.
   • (unique solution) A $q$-solution $u$ is said to be a unique solution to (1.2) if it is also a $\widetilde{q}$-solution for all $\widetilde{q}\in(1,q_{\mathcal{H}})$.
@@
 Due to the argument below (1.5), one cannot expect solutions to (1.2) with integrability in space larger or equal to $2q_{\mathcal{H}}$.
-Furthermore, unique solution of (1.2) are independent of the choice of $q\in(1,q_{\mathcal{H}})$.
+Furthermore, even if solutions are constructed by (1.3), the uniqueness class is independent of this splitting.
+Moreover, the unique solution of (1.2) is independent of the choice of $q\in(1,q_{\mathcal{H}})$.
 Such independence is expected from solutions to (1.2) in light of (1.6).
@@
   1. (1)
-    There exists a unique solution of (1.2) in the sense of 1.2 with paths in
+    There exists a unique solution of (1.2) in the sense of Definition 1.2 with paths in
     $u\in C([0,T];\mathbb{L}^{2})\ \ \mathbf{P}-a.s.$
@@
     $u\in C([t_{0},T];C^{\infty}({\mathcal{O}_{0}};\mathbb{R}^{2}))\ \ \mathbf{P}-a.s.$
-The proof of Theorem 1.3(1) and (2) are given subsection 3.3 and subsection 4.2, respectively. Routine extensions of the above are commented in 1.4 below.
+The proof of Theorem 1.3(1) and (2) are given in Section 3.3 and Section 4.2, respectively. Routine extensions of the above are commented in Remark 1.4 below.
 Next, let us discuss the main ideas behind the proof of Theorem 1.3.
@@
 In particular, $N$ is chosen large enough such that the equation for $\overline{v}$ has a forcing in $L^{2}(0,T;H^{-1})$ and therefore is regular enough to prove the existence and uniqueness of Leray solutions, see Theorem 2.6 below.
-The reader is referred to subsection 3.2 for further discussions.
+The reader is referred to Section 3.2 for further discussions.
 The interior regularity of $u$ in Theorem 1.3(2) is treated considering again the splitting $u=v+w_{g}$ introduced above.
@@
 On the contrary, the low regularity of $v$ does not allow us to study directly its interior regularity by Serrin’s argument as in [1].
-For this reason, we rely on the splitting (1.7) analysed in subsection 3.2 to study the well-posedness of (1.4).
+For this reason, we rely on the splitting (1.8) analysed in Section 3.2 to study the well-posedness of (1.5).
 Combining maximal $L^{p}$ regularity techniques for studying the interior regularity of the $v_{i}$’s, an induction argument and a Serrin argument for treating the interior regularity of $\overline{v}$, we obtain the required regularity of $v$.
-According to [58] (see also [41, Section 13.1]), it seems not possible to gain higher-order interior time-regularity for the Navier-Stokes problem.
-This fact is in contrast to the case of the heat equation with white noise boundary conditions, see [12].
-The reason behind this is the presence of the unknown pressure $P$ which, due to its non-local nature, provides a connection between the interior and the boundary regularity where the noise acts.
-To conclude, let us note that, in contrast to [11, 31], we employ a different splitting scheme to prove existence due to the presence of the transport-type terms originated by $w_{g}$.
+As shown in [60] (see also [43, Section 13.1]) and similarly to [1], higher-order interior time regularity does not appear to be attainable for our Navier–Stokes problem with stochastic boundary conditions.
+This stands in contrast to the case of the heat equation with white noise boundary conditions, as studied in [12].
+The underlying reason for this phenomenon is the presence of the unknown pressure $P$, whose non-local nature creates a link between interior and boundary regularity, where the noise acts.
+To conclude, let us point out that, in contrast to [11, 33], we employ a different splitting scheme to prove existence due to the presence of the transport-type terms originated by $w_{g}$.
 Moreover, the number of splitting $N$ depends on how much the Sobolev index of the space $C([0,T];L^{2q}(\mathcal{O};{\mathbb{R}}^{2}))$, i.e.
@@
 One can readily check that Theorem 1.3 extends in the following cases:
-  • (The case of bounded domains) If $\mathcal{O}$ is replaced by a smooth $C^{2}$-bounded domain in ${\mathbb{R}}^{2}$.
+  • (Bounded domains) If $\mathcal{O}$ is replaced by a smooth $C^{2}$-bounded domain in ${\mathbb{R}}^{2}$.
 However, we prefer to keep the same geometry of [1] for two reasons.
@@
 Stochastic convolutions with respect to such processes were analyzed in [13, 18, 19].
-  • (Time-dependent $g$) The term $g$ in the boundary noise $g\dot{W}^{\mathcal{H}}$ depends on time as long as it is progressively measurable and the corresponding process $w_{g}$ satisfies (1.5) holds.
+  • (Time-dependent $g$) The term $g$ in the boundary noise $g\dot{W}^{\mathcal{H}}$ depends on time as long as it is progressively measurable and the corresponding process $w_{g}$ satisfies (1.6).
+  • (Full stochastic Couette flow) More general boundary conditions like
+    $\begin{cases}u_{1}=U_{up}+g\dot{W}^{\mathcal{H}}\quad&\text{on }(0,T)\times\Gamma_{u},\\ u_{1}=U_{b}\quad&\text{on }(0,T)\times\Gamma_{b}.\end{cases}$
+    can be considered for sufficiently smooth velocity fields $U_{up},U_{b}$ such that the corresponding process $w_{g}$ satisfies (1.6).
+For example the case $U_{up},U_{b}\in L^{2}((0,T)\times\mathbb{T})$ can be treated.
+The above can be seen as a Couette flow with uncertainty on the velocity of one of the two surfaces.
 1.2. Overview
-The paper is organized as follows.
-In section 2 we introduce the functional setting to deal with problem (1.2).
-The proof of Theorem 1.3 is the object of section 3 and section 4.
-In particular, the global well-posedness, i.e.
-item 1, is addressed in section 3 considering the linear problem (1.3) in subsection 3.1 and the nonlinear one (1.4) in subsection 3.2.
-The interior regularity, i.e.
-item 2, is the object of section 4.
-In particular, in subsection 4.1 we study the interior regularity of the solution of the linear problem (1.3), while in subsection 4.2 we consider the nonlinear problem (1.4).
+In Section 2, we introduce the functional framework required to study problem (1.2).
+The proof of Theorem 1.3 is developed in Section 3 and Section 4.
+Specifically, global well-posedness, i.e.
+Item 1, is addressed in Section 3, where we first analyze the linear problem (1.4) in Section 3.1, followed by the nonlinear problem (1.5) in Section 3.2.
+Interior regularity, i.e.
+Item 2, is the focus of Section 4.
+In particular, Section 4.1 is devoted to the interior regularity of the solution to the linear problem (1.4), while Section 4.2 deals with the nonlinear problem (1.5).
 1.3. Notation
 Here we collect some notation which will be used throughout the paper.
-Further notation will be introduced where needed.
-By $C$ we will denote several constants, perhaps changing value line by line.
-If we want to keep track of the dependence of $C$ from some parameter $\xi$ we will use the symbol $C(\xi)$.
-Sometimes we will use the notation $a\lesssim b$ (resp.
-$a\lesssim_{\xi}b$) if it exists a constant such that $a\leq Cb$ (resp.
+Additional notation will be introduced where needed.
+We use $C$ to denote a generic constant, which may vary from line to line.
+When it is important to emphasize the dependence of $C$ on a parameter $\xi$, we write $C(\xi)$.
+Moreover, we sometimes write $a\lesssim b$ (resp.
+$a\lesssim_{\xi}b$) to mean that there exists a constant $C$ (resp.
+$C(\xi)$) such that $a\leq Cb$ (resp.
 $a\leq C(\xi)b$).
-Fix $q\in(1,\infty)$.
-For an integer $k\geq 1$, $W^{k,q}$ denotes the usual Sobolev spaces.
-In the non-integer case $s\in(0,\infty)\setminus{\mathbb{N}}$, we let $W^{s,q}=B^{s}_{q,q}$ where $B^{s}_{q,q}$ is the Besov space with smoothness $s$, and integrability $q$ and microscopic integrability $q$.
-Moreover, $H^{s,q}$ denotes the Bessel potential spaces.
-Both Besov and Bessel potential spaces can be defined using Littlewood-Paley decompositions and restrictions (see e.g.
-[56], [55, Section 6]) or using the interpolation methods starting with the standard Sobolev spaces $W^{k,q}$ (see e.g.
-[8, Chapter 6]).
-Finally, we set $\mathcal{A}^{s,q}(D;{\mathbb{R}}^{d})=(\mathcal{A}^{s,q}(D))^{d}$ for an integer $d\geq 1$, a domain $D$ and $\mathcal{A}\in\{W,H\}$.
-Let $\mathcal{K}_{1}$ and $\mathcal{K}_{2}$ be two Hilbert spaces.
+Let $q\in(1,\infty)$ be fixed.
+For any integer $k\geq 1$, we denote by $W^{k,q}$ the standard Sobolev spaces.
+In the case of non-integer smoothness $s\in(0,\infty)\setminus\mathbb{N}$, we define $W^{s,q}=B^{s}_{q,q}$ where $B^{s}_{q,q}$ is the Besov space with smoothness $s$, and integrability $q$ and microscopic integrability $q$.
+We also denote by $H^{s,q}$ the Bessel potential spaces.
+Both Besov and Bessel potential spaces can be defined via Littlewood–Paley theory (see, e.g., [58], [57, Section 6]), or through interpolation methods based on the classical Sobolev spaces $W^{k,q}$ (see, e.g., [8, Chapter 6]).
+For a domain $D\subset{\mathbb{R}}^{n}$, integer $d\geq 1$, and $\mathcal{A}\in\{W,H\}$, we define the vector-valued spaces by $\mathcal{A}^{s,q}(D;{\mathbb{R}}^{d})=(\mathcal{A}^{s,q}(D))^{d}$.
+Let $\mathcal{K}_{1}$ and $\mathcal{K}_{2}$ be two separable Hilbert spaces.
 We denote by $\mathscr{L}_{2}(\mathcal{K}_{1},\mathcal{K}_{2})$ the set of Hilbert-Schmidt operators from $\mathcal{K}_{1}$ to $\mathcal{K}_{2}$.
-Below, we need the following Fubini-type result:
-$H^{s}(D;\mathcal{K}_{1})=\mathscr{L}_{2}(\mathcal{K}_{1},H^{s}(D))\ \ \text{ for all }s\in{\mathbb{R}}.$
-The above follows from [37, Theorem 9.3.6] and interpolation.
+We will use the following Fubini-type identity:
+$H^{s}(D;\mathcal{K}_{1})=\mathscr{L}_{2}(\mathcal{K}_{1},H^{s}(D))\ \ \text{ for all }s\in{\mathbb{R}},$
+which follows from [39, Theorem 9.3.6] and interpolation theory.
 2. Preliminaries
@@
 We begin by introducing the Helmholtz projection on $L^{r}(\mathcal{O};{\mathbb{R}}^{2})$, see e.g.
-[53, Subsection 7.4].
-Let $f\in L^{r}(\mathcal{O};{\mathbb{R}}^{2})$ and let $\psi_{f}\in W^{1,r}(\mathcal{O})$ be the unique solution of the following elliptic problem
-(2.1) $\Delta\psi_{f}&=\operatorname{div}f\quad&\text{ on }&\mathcal{O},\\ \partial_{n}\psi_{f}&=f\cdot\widehat{n}&\text{ on }&\Gamma_{u}\cup\Gamma_{b}.\end{aligned}\right.$
-Here $n$ denotes the exterior normal vector field on $\partial\mathcal{O}$. Of course, the above elliptic problem is interpreted in its naturally weak formulation:
+[55, Subsection 7.4].
+As described also in [1, Section 2.1], the projection can be defined via an elliptic problem.
+We recall its construction here for the sake of completeness.
+Let $f\in L^{r}(\mathcal{O};{\mathbb{R}}^{2})$ and let $\psi_{f}$ be the unique solution of the following elliptic problem
+(2.1) $\Delta\psi_{f}&=\operatorname{div}f\quad&\text{ on }&\mathcal{O},\\ \partial_{\widehat{n}}\psi_{f}&=f\cdot\widehat{n}&\text{ on }&\Gamma_{u}\cup\Gamma_{b}.\end{aligned}\right.$
+Here $\widehat{n}$ denotes the exterior normal vector field on $\partial\mathcal{O}$. This problem is understood in its standard weak formulation:
 (2.2) $\int_{\mathcal{O}}\nabla\psi_{f}\cdot\nabla\varphi\,\mathrm{d}x\mathrm{d}z=\int_{\mathcal{O}}f\cdot\nabla\varphi\,\mathrm{d}x\mathrm{d}z\ \ \text{ for all }\varphi\in C^{\infty}(\mathcal{O}).$
-By [53, Corollary 7.4.4] , we have $\psi_{f}\in W^{1,r}(\mathcal{O})$ and $\|\nabla\psi_{f}\|_{L^{r}(\mathcal{O};{\mathbb{R}}^{2})}\lesssim\|f\|_{L^{r}(\mathcal{O};{\mathbb{R}}^{2})}$.
-Then the Helmholtz projection is given by $\mathbb{P}_{r}$ is defined as
+By [55, Corollary 7.4.4] , we have $\psi_{f}\in W^{1,r}(\mathcal{O})$ and $\|\nabla\psi_{f}\|_{L^{r}(\mathcal{O};{\mathbb{R}}^{2})}\lesssim\|f\|_{L^{r}(\mathcal{O};{\mathbb{R}}^{2})}$.
+The Helmholtz projection $\mathbb{P}_{r}:L^{r}(\mathcal{O};{\mathbb{R}}^{2})\rightarrow L^{r}(\mathcal{O};{\mathbb{R}}^{2})$ is then defined by
 $\mathbb{P}_{r}f=f-\nabla\psi_{f},\quad f\in L^{r}(\mathcal{O};{\mathbb{R}}^{2}).$
-Next, we define the Stokes operator on $L^{r}(\mathcal{O};{\mathbb{R}}^{2})$.
+Next, we define the Stokes operator on $L^{r}(\mathcal{O};{\mathbb{R}}^{2})$ corresponding to the boundary conditions considered in (1.2).
 For notational convenience, we define $A_{r}$ as minus the Stokes operator so that $A_{r}$ is a positive operator for $r=2$ (i.e.
@@
 By [48, 42], it is enough to provide the boundedness of the $H^{\infty}$-calculus for $A_{r}$.
-The reader is referred to [53, Chapters 3 and 4] and [37, Chapter 10] for the main notation and basic results on the $H^{\infty}$-calculus.
+The reader is referred to [55, Chapters 3 and 4] and [39, Chapter 10] for the main notation and basic results on the $H^{\infty}$-calculus.
 Contrary to [1], the boundary conditions we are interested in here are much more classical.
@@
 The reader is referred e.g.
-to [36, Section 2.8], [47], [32] and [38, Section 9] for the proof of this nowadays classical statement.
+to [38, Section 2.8], [49], [34] and [40, Section 9] for the proof of this nowadays classical statement.
 Lemma 2.1.
@@
 $\mathsf{D}(A_{r}^{\alpha})=\left\{\begin{aligned} &\mathbb{H}^{2\alpha,r}\ \ &\text{ if }&\ \alpha<\frac{1}{2r},\\ &\big{\{}u\in\mathbb{H}^{2\alpha,r}\,:\,u|_{\partial\mathcal{O}}=0\big{\}}\ \ &\text{ if }&\ \frac{1}{2r}<\alpha\leq 1.\end{aligned}\right.$
-The above implies that $-A_{r}$ generates an analytic semigroup on $\mathbb{L}^{r}$ which admits stochastic and deterministic maximal $L^{p}$-regularity for all $p\in(1,+\infty)$, see [53, Chapter 3-4] and [46].
+The above implies that $-A_{r}$ generates an analytic semigroup on $\mathbb{L}^{r}$ which admits stochastic and deterministic maximal $L^{p}$-regularity for all $p\in(1,+\infty)$, see [55, Chapter 3-4] and [48].
 We denote such semigroup by $S_{r}(t)$.
@@
 where $\sim$ denotes the completion.
-Indeed, since $0\in\rho(A_{r})$ by 2.1, we have that $f\mapsto\|A_{r}^{{\alpha}}f\|_{\mathbb{L}^{r}}$ is a norm for all $\alpha<0$.
+Indeed, since $0\in\rho(A_{r})$ by Lemma 2.1, we have that $f\mapsto\|A_{r}^{{\alpha}}f\|_{\mathbb{L}^{r}}$ is a norm for all $\alpha<0$.
 Since $(A_{r})^{*}=A_{r^{\prime}}$, it follows that (see e.g.
@@
 (2.3) $(X_{\alpha,A_{r}})^{*}=X_{-\alpha,A_{r^{\prime}}}.$
-For convenience of notation, we simply write $A,\ S(t)$ in place of $A_{2}$ and $S(t)$. Moreover we define
+For notational convenience, we will write $A,\ S(t)$ instead of $A_{2}$ and $S(t)$. Moreover we define
 $H:=\mathbb{L}^{2},\quad V:=\mathsf{D}(A^{1/2}).$
 […]
+For the convenience of the reader, we recall here the definition of the deterministic maximal $L^{p}$-regularity, since our techniques heavily rely on it.
+Definition 2.2.
+([55, Definition 3.5.1]) We let $\mathcal{X}$ stand for a Banach space and $\mathcal{A}$ a linear closed operator on $\mathcal{X}$ with domain $\mathsf{D}(\mathcal{A})$.
+We say that the inhomogeneous initial value problem on $L^{p}(0,T;\mathcal{X})$ given by
+(2.4) $\dot{u}(t)+\mathcal{A}u(t)=f(t),~{}~{}u(0)=u_{0}$
+admits maximal $L^{p}$-regularity, if for each $f\in L^{p}(0,T;\mathcal{X})$ and $u_{0}\in(\mathcal{X},\mathsf{D}(\mathcal{A}))_{1-1/p,p}$ there exists a unique $u\in W^{1,p}(0,T;\mathcal{X})\cap L^{p}(0,T;\mathsf{D}(\mathcal{A}))$ satisfying (2.4) a.e.
+in $(0,T)$.
 2.2. The Dirichlet map
@@
 Theorem 2.3.
-For all $g\in H^{-\frac{1}{2}}(\Gamma_{u})$ there exists a unique $(u,\pi)\in H\times H^{-1}(\mathcal{O})/\mathbb{R}$ very weak solution of (2.4). Moreover $(u,\pi)$ satisfy
+For all $g\in H^{-\frac{1}{2}}(\Gamma_{u})$ there exists a unique very weak solution $(u,\pi)\in H\times H^{-1}(\mathcal{O})/\mathbb{R}$ of (2.5). Moreover $(u,\pi)$ satisfy
 (2.7) $\lVert u\rVert+\lVert\pi\rVert_{H^{-1}(\mathcal{O})/\mathbb{R}}\leq C\lVert g\rVert_{H^{-\frac{1}{2}}(\Gamma_{u})}.$
@@
 Proof.
-To begin, recall that $H^{s}(\Gamma_{u};U)=\mathscr{L}_{2}(U,H^{s}(\Gamma_{u}))$ for all $s\in{\mathbb{R}}$, see subsection 1.3.
+To begin, recall that $H^{s}(\Gamma_{u};U)=\mathscr{L}_{2}(U,H^{s}(\Gamma_{u}))$ for all $s\in{\mathbb{R}}$, see Section 1.3.
 Hence, due to the ideal property of Hilbert-Schmidt operators, it is enough to consider the scalar case $U={\mathbb{R}}$.
@@
 $\mathcal{D}:H^{2\theta-\frac{1}{2}}(\Gamma_{u})\to\mathbb{H}^{2\theta}(\mathcal{O})\ \ \text{ for all }\theta\in(0,1).$
-Hence, the claim now follows from the description of the fractional power of $A$ in 2.1. ∎
+Hence, the claim now follows from the description of the fractional power of $A$ in Lemma 2.1. ∎
 2.3. Deterministic Navier-Stokes equations
@@
 $\left\langle B\left(u,v\right),\phi\right\rangle_{X_{-1/2,A_{q^{\prime}}},X_{1/2,A_{q}}}=-b\left(u,\phi,v\right)=-\int_{\mathcal{O}}\left(u\cdot\nabla\phi\right)\cdot v\,\mathrm{d}x\mathrm{d}z$
-for all $\phi\in X_{1/2,A_{q}}$. Moreover, when $u\cdot\nabla v\in L^{r}\left(\mathcal{O};\mathbb{R}^{2}\right)$, it is explicitly given by
+for all $\phi\in X_{1/2,A_{q}}$. Moreover, if $u\cdot\nabla v\in L^{r}\left(\mathcal{O};\mathbb{R}^{2}\right)$ for some $r\in(1,\infty)$, it is explicitly given by
 $B\left(u,v\right)=\mathbb{P}(u\cdot\nabla v).$
@@
 $=\left\langle\overline{u}_{0},\phi\right\rangle-\int_{0}^{t}\left\langle\overline{u}\left(s\right),A\phi\right\rangle\,\mathrm{d}s+\int_{0}^{t}\left\langle\overline{f}\left(s\right),\phi\right\rangle_{V^{*},V}\,\mathrm{d}s.$
-The well-posedness of (2.8) in the sense of 2.4 is a well-known fact. Indeed the following theorem holds, see for instance [43, 59, 60].
+The well-posedness of (2.9) in the sense of Definition 2.5 is a well-known fact. Indeed the following theorem holds, see for instance [45, 62, 63].
 Theorem 2.6.
@@
 If $\left(\overline{u}_{0}^{n}\right)_{n\in\mathbb{N}}$ is a sequence in $H$ converging to $\overline{u}_{0}\in H$ and $\left(\overline{f}^{n}\right)_{n\in\mathbb{N}}$ is a sequence in $L^{2}\left(0,T;V^{*}\right)$ converging to $\overline{f}\in L^{2}\left(0,T;V^{*}\right)$, then the corresponding unique solutions $\left(\overline{u}^{n}\right)_{n\in\mathbb{N}}$ converge to the corresponding solution $\overline{u}$ in $C\left(\left[0,T\right];H\right)$ and in $L^{2}\left(0,T;V\right)$.
-We end this section with the following lemma, which is the analogous of [29, Lemma 1.14] to the $L^{p}_{t}L^{q}_{x}$ framework.
+We end this section with the following lemma, which generalizes [31, Lemma 1.14] to the $L^{p}_{t}L^{q}_{x}$ setting. In particular, [31, Lemma 1.14] corresponds to the specific case $p=q=4.$
 Lemma 2.7.
-If $p\geq\frac{2q}{q-2}$, $u\in C([0,T];H)\cap L^{2}(0,T;V),\ v\in L^{p}(0,T;\mathbb{L}^{q})$, then
+If $q>2$ and $p\geq\frac{2q}{q-2}$, $u\in C([0,T];H)\cap L^{2}(0,T;V),\ v\in L^{p}(0,T;\mathbb{L}^{q})$, then
 (2.11) $B(u,v)\in L^{2}(0,T;V^{*}),$
@@
 Definition 2.8.
-Let $U$ be a separable Hilbert space. A $U$-cylindrical fractional Brownian motion $(W^{\mathcal{H}}(t))_{t\geq 0}$ with Hurst index $H\in(0,1)$ is defined by the formal series
+Let $U$ be a separable Hilbert space. A $U$-cylindrical fractional Brownian motion $(W^{\mathcal{H}}(t))_{t\geq 0}$ with Hurst index $\mathcal{H}\in(0,1)$ is defined by the formal series
 $W^{\mathcal{H}}(t)=\sum\limits_{n=1}^{\infty}b^{\mathcal{H}}_{n}(t)e_{n},$
@@
 For $\mathcal{H}=1/2$ one obtains a cylindrical Brownian motion. However for $\mathcal{H}\neq 1/2$ the fbm exhibits a totally different behaviour, in particular is neither Markov nor a semimartingale.
-For our aims in 3.1, we need the following results on the regularity of stochastic convolutions established in [22, Corollary 3.1].
+For our aims in Proposition 3.1, we need the following results on the regularity of stochastic convolutions established in [23, Corollary 3.1] and [24, Proposition 11.6].
 Lemma 2.9.
-([22, Corollary 3.1]) Let $A$ be the generator of an analytic $C_{0}$-semigroup $(S(t))_{t\geq 0}$ on a separable Hilbert space $U_{1}$, $\Phi\in\mathcal{L}(U,U_{1})$. Assume that
-(2.14) $\|S(t)\Phi\|_{\mathcal{L}_{2}(U_{1},U)}\leq t^{-\gamma}\text{ for}\gamma<\mathcal{H}.$
+([23, Corollary 3.1] and [24, Proposition 11.6]) Let $A$ be the generator of an analytic $C_{0}$-semigroup $(S(t))_{t\geq 0}$ on a separable Hilbert space $U_{1}$, $\Phi\in\mathscr{L}(U,U_{1})$.
+Assume that
+(2.15) $\|S(t)\Phi\|_{\mathscr{L}_{2}(U,U_{1})}\leq t^{-\gamma}\text{ for}\gamma<\mathcal{H}.$
 Then the stochastic convolution $\int_{0}^{t}S(t-s)\Phi~{}\mathrm{d}W^{\mathcal{H}}(s)$ has $\mathbf{P}$-a.s.
-$\gamma_{1}$-Hölder continuous trajectories in $\mathsf{D}(A^{\gamma_{2}})$, for $0\leq\gamma_{1}+\gamma_{2}<\mathcal{H}$.
-If $\Phi\in\mathcal{L}_{2}(U_{1},U)$, then the assumption (2.14) is satisfied for $\gamma=0$.
+$\gamma_{1}$-Hölder continuous trajectories in $\mathsf{D}(A^{\gamma_{2}})$, for $0\leq\gamma_{1}+\gamma_{2}<\mathcal{H}-{\gamma}$.
+If $\Phi\in\mathscr{L}_{2}(U,U_{1})$, then the assumption (2.15) is satisfied for $\gamma=0$.
 3. Global well-posedness
@@
 This section is organized as follows.
-Firstly, in subsection 4.1 we first prove that the solution $w_{g}$ of the 2D Stokes equations with boundary noise (1.3) satisfies (1.5).
-Secondly, in subsection 3.2, we prove the existence of a $q$-solution to (1.2) by studying the auxiliary Navier-Stokes problem (1.4) for a given forcing term $w=w_{g}$ satisfying the regularity assumption as in (1.5) for a given $q$.
-Finally, in subsection 3.3, we prove the uniqueness of solutions to (1.2) therefore concluding the proof of Theorem 1.3(1).
-Recall that ($q$-)solutions of (1.2) are defined in 1.2.
+Firstly, in Section 3.1 we prove that the solution $w_{g}$ of the 2D Stokes equations with boundary noise (1.4) satisfies (1.6).
+Secondly, in Section 3.2, we prove the existence of a $q$-solution to (1.2) by studying the auxiliary Navier-Stokes problem (1.5) for a given forcing term $w=w_{g}$ satisfying the regularity assumption as in (1.6) for a given $q$.
+Finally, in Section 3.3, we prove the uniqueness of solutions to (1.2) therefore concluding the proof of Theorem 1.3(1).
+Recall that ($q$-)solutions of (1.2) are defined in Definition 1.2.
 3.1. Stokes equations
-As discussed in subsection 1.1, we start by considering the linear problem (1.3). According to [20] and [21, Chapter 15], the mild solution $w_{g}$ of the former problem is formally given by
+As discussed in Section 1.1, we start by considering the linear problem (1.4). According to [21] and [22, Chapter 15], the mild solution $w_{g}$ of the former problem is formally given by
 (3.1) $w_{g}(t)=A\int_{0}^{t}S(t-s)\mathcal{D}[g]\,\mathrm{d}W^{\mathcal{H}}(s).$
-Here $A$ is (minus) the Stokes operator with homogeneous boundary conditions (see Subsection 2.1).
+Here $A$ is (minus) the Stokes operator with homogeneous boundary conditions as defined in Section 2.1.
 Next, we prove that $w_{g}$ is well-defined in sufficiently regular function spaces therefore allowing us to treat the nonlinearity in the Navier-Stokes equations.
@@
 (3.2) $w_{g}\in L^{p}(\Omega;C([0,T];\mathsf{D}(A^{\mathcal{H}-\frac{3}{4}-\frac{s}{2}-\varepsilon})))\ \text{ for all }p\in(1,\infty).$
-In particular, for all $r\in(2,2q_{\mathcal{H}})$
+In particular, for all $r\in(2,2q_{\mathcal{H}})$,
 […]
 Proof.
-Note that, thanks to 2.3
-$\mathcal{D}g\in\mathcal{L}_{2}(U,\mathsf{D}(A^{\frac{1}{4}-\frac{s+\varepsilon}{2}})).$
-Hence, by 2.8, a.s.,
+Note that, thanks to Corollary 2.4
+$\mathcal{D}g\in\mathscr{L}_{2}(U,\mathsf{D}(A^{\frac{1}{4}-\frac{s+\varepsilon}{2}})).$
+Hence, by Lemma 2.9, a.s.,
 $w_{g}=\underbrace{A^{\frac{3}{4}+\frac{s+\varepsilon}{2}}\underbrace{\int_{0}^{\cdot}S(\cdot-s)A^{\frac{1}{4}-\frac{s+\varepsilon}{2}}\mathcal{D}g\,\mathrm{d}W^{\mathcal{H}}(s)}_{\in\,C([0,T];\mathsf{D}(A^{\mathcal{H}-\frac{\varepsilon}{2}}))}}_{\in\,C([0,T];\mathsf{D}(A^{\mathcal{H}-\frac{3}{4}-\frac{s}{2}-\varepsilon}))}.$
-The arbitrariness of $\varepsilon>0$ yields (3.2). To prove (3.3), note that, by 2.1,
+The arbitrariness of $\varepsilon>0$ yields (3.2). To prove (3.3), note that, by Lemma 2.1,
 $\mathsf{D}(A^{\mathcal{H}-\frac{3}{4}-\frac{s}{2}-\varepsilon})\subset H^{2\mathcal{H}-\frac{3}{2}-s-2\varepsilon}(\mathcal{O};{\mathbb{R}}^{2}).$
@@
 Remark 3.2 (Necessity of the $L^{p}$-setting for $v$).
-In the setting of 3.1, we have $2\mathcal{H}-\frac{3}{2}-s<\frac{1}{2}.$ Therefore, for all choices of $\mathcal{H}$ and $s$ in 1.1, it follows that
+In the setting of Proposition 3.1, we have $2\mathcal{H}-\frac{3}{2}-s<\frac{1}{2}.$ Therefore, for all choices of $\mathcal{H}$ and $s$ in 1.1, it follows that
 $H^{2\mathcal{H}-\frac{3}{2}-s-2\varepsilon}(\mathcal{O};{\mathbb{R}}^{2})\not\hookrightarrow L^{4}(\mathcal{O};{\mathbb{R}}^{2}).$
-Thus, (3.3) holds with $r<4$ and therefore $B(w_{g},w_{g})\not\in L^{2}(0,T;V^{*}).$ In particular, in the next subsection, we cannot avoid the use of $L^{p}$-setting in space, cf., the comments below 3.4.
+Thus, (3.3) holds with $r<4$ and therefore $B(w_{g},w_{g})\not\in L^{2}(0,T;V^{*}).$ In particular, in the next subsection, we cannot avoid the use of $L^{p}$-setting in space, cf.
+the comments below 3.6.
 Remark 3.3.
-Previous results with white noise boundary conditions [10], [1] exploited stochastical maximal $L^{p}$ regularity techniques to study the linear part of the problem.
+Previous results with white noise boundary conditions [1, 10] exploited stochastical maximal $L^{p}$ regularity techniques to study the linear part of the problem.
 Here is worth mentioning that we employed the more standard Hilbert value framework because it produces the sharpest result on the regularity of the stochastic convolution in terms of the Hurst parameter $\mathcal{H}$.
-Indeed, assuming just for simplicity the case $s=0$ and $g\in L^{p}(\Gamma_{u};U)$ for some $p\in[2,+\infty)$, then by 2.3, [17, Proposition 4.5] and arguing as above we have
+Indeed, assuming just for simplicity the case $s=0$ and $g\in L^{p}(\Gamma_{u};U)$ for some $p\in[2,+\infty)$, then by Corollary 2.4, [18, Proposition 4.5] and arguing as above we have
 $w_{g}\in C([0,T];\mathsf{D}(A_{p}^{\mathcal{H}-1+\frac{1}{2p}-\varepsilon})).$
 In particular $w_{g}\in C([0,T];\mathbb{L}^{r})$ for some $r>2$ if $\mathcal{H}>1-\frac{1}{2p}$. Therefore the right-hand side is minimized and we can use the rougher noise for $p=2$.
+We end this subsection showing a lemma concerning the relation between the mild and the weak formulation of (1.4) as defined below.
+Definition 3.4.
+Let 1.1 be satisfied. A stochastic process $w$ is a weak solution of (1.4) if it is $\mathcal{F}$-progressively measurable with $\mathbf{P}-a.s.$ paths in
+$w_{g}\in C(0,T;\mathbb{L}^{r})$
+for some $r\geq 2$, and $\mathbf{P}-a.s.$ for all $\phi\in\mathsf{D}(A)$ and $t\in[0,T]$,
+(3.4) $\langle w_{g}(t),\phi\rangle=-\int_{0}^{t}\langle w_{g}(s),A\phi\rangle\,\mathrm{d}s-\langle g,\widehat{n}\cdot\nabla\phi\rangle_{H^{-s}(\Gamma_{u}),H^{s}(\Gamma_{u})}W^{\mathcal{H}}_{t}.$
+As above, $\widehat{n}$ denotes the exterior normal vector field on $\partial\mathcal{O}$. Since $g$ is time-independent, the last term in (3.4) can be rewritten as a stochastic integral as
+$\langle g,\widehat{n}\cdot\nabla\phi\rangle_{H^{-s}(\Gamma_{u}),H^{s}(\Gamma_{u})}W^{\mathcal{H}}_{t}=\int_{0}^{t}\langle g,\widehat{n}\cdot\nabla\phi\rangle_{H^{-s}(\Gamma_{u}),H^{s}(\Gamma_{u})}\,\mathrm{d}W^{\mathcal{H}}_{s}.$
+Lemma 3.5.
+Let 1.1 be satisfied. There exists a unique weak solution of (1.4) in the sense of Definition 3.4 and it is given by the formula (3.1).
+Proof.
+We split the proof into two steps.
+Step 1: There exists a unique weak solution of (1.4) and it is necessarily given by the mild formula (3.1).
+Let $\psi\in C^{1}([0,T];\mathsf{D}(A))$.
+Arguing as in the first step of the proof of [31, Theorem 1.7], see also [46, Lemma 3], one can readily check that $w_{g}$ satisfies
+$\langle w_{g}(t),\psi(t)\rangle=$ $\int_{0}^{t}\langle w_{g}(s),\partial_{s}\psi(s)\rangle\,\mathrm{d}s-\int_{0}^{t}\langle w_{g}(s),A\psi(s)\rangle\,\mathrm{d}s$
+(3.5) $-\int_{0}^{t}\langle g,\widehat{n}\cdot\nabla\psi(s)\rangle_{H^{-s}(\Gamma_{u}),H^{s}(\Gamma_{u})}\,\mathrm{d}W^{\mathcal{H}}_{s}$
+for each $t\in[0,T],\ \mathbf{P}-a.s.$ The stochastic integral in the relation above is well-defined as a real-valued stochastic integral.
+Indeed, recalling that $(W^{\mathcal{H}}_{t})_{t\geq 0}$ is a $U$-cylindrical fractional Brownian motion we observe that $\langle g,\widehat{n}\cdot\nabla\psi(s)\rangle_{H^{-s}(\Gamma_{u}),H^{s}(\Gamma_{u})}$ is given by the linear operator on $U$
+$h^{\prime}\mapsto\langle gh^{\prime},\widehat{n}\cdot\nabla\psi(s)\rangle_{H^{-s}(\Gamma_{u}),H^{s}(\Gamma_{u})}=L_{\psi}(gh^{\prime}),$
+where $L_{\psi}:=\langle\cdot,\widehat{n}\cdot\nabla\psi(\cdot)\rangle_{H^{-s}(\Gamma_{u}),H^{s}(\Gamma_{u})}.$ By the ideal property of the Hilbert-Schmidt operators we have $\mathscr{L}_{2}(U,H^{-s}(\Gamma_{u}))=H^{-s}(\Gamma_{u};U)$ and obtain that
+$\|\langle g,\widehat{n}\cdot\nabla\psi(s)\rangle_{H^{-s}(\Gamma_{u}),H^{s}(\Gamma_{u})}\|_{U^{*}}\lesssim\|g\|_{H^{-s}(\Gamma_{u})}\|\nabla\psi(s)\|_{H^{s}(\Gamma_{u})}~{}~{}\text{a.e.
+on }\Omega\times(0,T).$
+In conclusion, the stochastic integral in (3.1) is well-defined as a real-valued one (see [23, (2.16)]), since
+$\mathbf{E}\Big{|}\int_{0}^{t}\langle g,\widehat{n}\cdot\nabla\psi(s)\rangle_{H^{-s}(\Gamma_{u}),H^{s}(\Gamma_{u})}~{}\mathrm{d}W^{\mathcal{H}}_{s}\Big{|}^{2}$
+$\leq\mathcal{H}(2\mathcal{H}-1)\int_{0}^{t}\int_{0}^{t}\|\langle g,\widehat{n}\cdot\nabla\psi(s)\rangle\|_{U^{*}}\|\langle g,\widehat{n}\cdot\nabla\psi(v)\rangle\|_{U^{*}}|s-v|^{2\mathcal{H}-2}~{}\mathrm{d}s~{}\mathrm{d}v$
+$\leq\mathcal{H}(2\mathcal{H}-1)\|g\|^{2}_{H^{-s}(\Gamma_{u};U)}\int_{0}^{t}\int_{0}^{t}\|\nabla\psi(s)\|_{H^{s}(\Gamma_{u})}\|\nabla\psi(v)\|_{H^{s}(\Gamma_{u})}|s-v|^{2\mathcal{H}-2}~{}\mathrm{d}s~{}\mathrm{d}v,$
+which is finite since $\psi\in C^{1}([0,T];\mathsf{D}(A))$ and $\mathcal{H}>1/2$.
+Now consider $\phi\in\mathsf{D}(A^{2})$ and use $\psi_{t}(s)=S(t-s)\phi,\ s\in[0,t]$ as test function in (3.1) obtaining
+(3.6) $\langle w_{g}(t),\phi\rangle=$ $-\int_{0}^{t}\langle g,\widehat{n}\cdot\nabla S(t-s)\phi\rangle_{H^{-s}(\Gamma_{u}),H^{s}(\Gamma_{u})}\,\mathrm{d}W^{\mathcal{H}}_{s}.$
+Recalling the definition of the Dirichlet map $\mathcal{D}$, (3.6) can be rewritten as
+(3.7) $\langle w(t),\phi\rangle=$ $\int_{0}^{t}\langle\mathcal{D}[g],AS(t-s)\phi\rangle\,\mathrm{d}W^{\mathcal{H}}_{s}.$
+Then, exploiting the self-adjointness property of $S$ and $A$ we have that weak solutions of (1.4) satisfy the mild formulation. Therefore they are unique.
+Step 2: The mild formula (3.1) is a weak solution of (1.4) in the sense of Definition 3.4.
+We begin by noticing that $w_{g}$ has the required regularity due to Proposition 3.1.
+Let us test our mild formulation (3.1) against functions $\phi\in\mathsf{D}(A^{2})$.
+It holds, exploiting self-adjointness property of $S$ and $A$
+$\langle w(t),\phi\rangle$ $=\int_{0}^{t}\langle\mathcal{D}[g],AS(t-s)\phi\rangle\,\mathrm{d}W^{\mathcal{H}}_{s}$
+$=\int_{0}^{t}\langle g,\widehat{n}\cdot\nabla S(t-s)\phi\rangle_{H^{-s}(\Gamma_{u}),H^{s}(\Gamma_{u})}\,\mathrm{d}W^{\mathcal{H}}_{s}\quad\mathbf{P}-a.s.,$
+where in the last step we used the definition of Dirichlet map. To complete the proof of this step it is enough to show that
+$\int_{0}^{t}\langle g,\widehat{n}\cdot\nabla S(t-s)\phi\rangle_{H^{-s}(\Gamma_{u}),H^{s}(\Gamma_{u})}\,\mathrm{d}W^{\mathcal{H}}_{s}=-\int_{0}^{t}\langle w_{g}(s),A\phi\rangle\,\mathrm{d}s$
+(3.8) $+\langle g,\widehat{n}\cdot\nabla\phi\rangle_{H^{-s}(\Gamma_{u}),H^{s}(\Gamma_{u})}\,W^{\mathcal{H}}_{t}\quad\mathbf{P}-a.s.$
+The relation (3.1) is true. Indeed,
+(3.9) $\int_{0}^{t}\langle w_{g}(s),A\phi\rangle\,\mathrm{d}s$ $=\int_{0}^{t}\,\mathrm{d}s\int_{0}^{s}\langle\mathcal{D}[g],S(s-\tau)A^{2}\phi\rangle\,\mathrm{d}W^{\mathcal{H}}(\tau)\quad\mathbf{P}-a.s.$
+The double integrals in (3.9) can be exchanged via stochastic Fubini’s Theorem, see [50, 3]. Therefore the double integral in the right-hand side of (3.9) can be rewritten as
+$\int_{0}^{t}\,\mathrm{d}s\int_{0}^{s}\langle\mathcal{D}[g],S(s-\tau)A^{2}\phi\rangle\,\mathrm{d}W^{\mathcal{H}}_{\tau}$
+$\qquad=\int_{0}^{t}\,\mathrm{d}W^{\mathcal{H}}(\tau)\int_{\tau}^{t}\langle\mathcal{D}[g],S(s-\tau)A^{2}\phi\rangle\,\mathrm{d}s$
+$\qquad=\langle\mathcal{D}[g],A\phi\rangle W^{\mathcal{H}}_{t}-\int_{0}^{t}\langle\mathcal{D}[g],AS(t-\tau)\phi\rangle\,\mathrm{d}W^{\mathcal{H}}_{\tau}$
+$\qquad=\langle g,\widehat{n}\cdot\nabla\phi\rangle_{H^{-s}(\Gamma_{u}),H^{s}(\Gamma_{u})}\,W^{\mathcal{H}}_{t}$
+$\qquad-\int_{0}^{t}\langle g,\widehat{n}\cdot\nabla S(t-\tau)\phi\rangle_{H^{-s}(\Gamma_{u}),H^{s}(\Gamma_{u})}\,\mathrm{d}W^{\mathcal{H}}_{\tau}\quad\mathbf{P}-a.s.$
+Inserting this expression in (3.9), (3.1) holds and the proof is complete. ∎
 3.2. Auxiliary Navier-Stokes type equations
@@
 […]
-with $A_{q}$ is the Stokes operator on $\mathbb{L}^{q}$ and $B$ is the bilinear nonlinearity as in Subsection 2.3. Finally, $q=r/2$ and $(w,r)$ satisfies the following
+with $A_{q}$ is the Stokes operator on $\mathbb{L}^{q}$ and $B$ is the bilinear nonlinearity as in Section 2.3. Finally, $q=r/2$ and $(w,r)$ satisfies the following
 Assumption 3.6.
 $w\in C([0,T];\mathbb{L}^{r})$ for some $r\in(2,4)$.
-Note that the above assumption is satisfied $\mathbf{P}-a.s.$ with $w=w_{g}$, as it follows from 3.1.
-Moreover, the limitation $r<4$ is motivated by 3.2.
+Note that the above assumption is satisfied $\mathbf{P}-a.s.$ with $w=w_{g}$, as it follows from Proposition 3.1.
+Moreover, the limitation $r<4$ is motivated by Remark 3.2.
 In particular, the arguments used in [1] do not apply to (3.10).
-Indeed, if 3.4 holds, then
+Indeed, if 3.6 holds, then
 $B(w,w)\not\in L^{2}(0,T;H^{-1}(\mathcal{O};{\mathbb{R}}^{2})).$
@@
 However, $B(v,v)\in L^{\infty}(0,T;H^{-1,q}(\mathcal{O};{\mathbb{R}}^{2}))$ for some $q>1$ as $r>2$, and therefore $L^{q}$-theory for (3.10) can be built.
-Next, let us describe the main idea behind our construction of a $r/2$-solution to (3.4).
-In what follows, the subcriticality of $L^{\infty}(0,T;\mathbb{L}^{r})$ with $r>1$ for the 2D Navier-Stokes equations (cf., the discussion below 1.2) plays a central role.
+Next, let us describe the main idea behind our construction of a solution to (3.10).
+In what follows, the subcriticality of $L^{\infty}(0,T;\mathbb{L}^{r})$ with $r>1$ for the 2D Navier-Stokes equations (cf.
+the discussion below Definition 1.2) plays a central role.
 Indeed, by subcriticality, given $q_{0}=\frac{r}{2}$, the solution $v_{0}$ to
@@
 Before going further, let us stress that the above procedure is reminiscent of the so-called ‘DaPrato-Debussche trick’ introduced in [20] and now is widely used in the context of stochastic PDEs.
-Let us now turn to the construction of a $q$-solution to (3.4). The above argument motivates the following splitting. Let $N$ be a positive integer such that
+Let us now turn to the construction of a solution to (3.10). The above argument motivates the following splitting. Let $N$ be a positive integer such that
 (3.11) $r\in\left[\frac{2(N+2)}{N+1},\frac{2(N+1)}{N}\right)$
@@
 Note that $v_{i}$ for $i\in\{1,\dots,N-1\}$ solves a (linear) Stokes problem, while the problem for $\overline{v}$ is a modified version of the Navier-Stokes equations.
-At least formally, it is clear that $v$ solves (3.4). The latter fact is a straightforward consequence of the following identity (letting $v_{N}:=\overline{v}$ for simplicity)
-$\sum_{i,j=0}^{N}B(v_{i},v_{j})$ $=\sum_{i=0}^{N}B(v_{i},v_{i})+\sum_{0\leq j<i\leq N}B(v_{i},v_{j})+\sum_{0\leq i<j\leq N}B(v_{i},v_{j})$
-$=\sum_{i=0}^{N}B(v_{i},v_{i})+\sum_{i=0}^{N-1}\sum_{j=0}^{i-2}B(v_{i-1},v_{j})+\sum_{j=0}^{N-1}\sum_{i=0}^{j-2}B(v_{i},v_{j-1})$
-$=B(\overline{v},\overline{v})+\sum_{i=0}^{N-1}\sum_{j=0}^{i-1}B(v_{i-1},v_{j})+\sum_{j=0}^{N-1}\sum_{i=0}^{j-2}B(v_{i},v_{j-1}).$
-To show rigorously that $v$ given in (3.6) with $(v_{0},\dots,v_{N-1},\overline{v})$ solving (3.7) is a $q$-solution to (3.4) we need to check that $v_{0},\dots,v_{N-1}$ and $\overline{v}$ are sufficiently regular.
-The appropriate regularity class for $v$ in (3.6) to obtain a $q$-solution is given in the following definition, see also 3.6 below.
+At least formally, it is clear that $v$ solves (3.10).
+The latter fact is a straightforward consequence of the following identity involving $B(v+w,v+w)$ (letting $v_{-1}=w,\ v_{N}:=\overline{v}$ for simplicity)
+$B(v+w,v+w)$ $=\sum_{i,j=-1}^{N}B(v_{i},v_{j})$
+$=\sum_{i=-1}^{N}B(v_{i},v_{i})+\sum_{i=-1}^{N}\sum_{j=-1}^{i-1}B(v_{i},v_{j})+\sum_{j=-1}^{N}\sum_{i=-1}^{j-1}B(v_{i},v_{j})$
+${=B(\overline{v},\overline{v})+\sum_{i=-1}^{N-1}B(v_{i},v_{i})+\sum_{i=-1}^{N-1}B(\overline{v},v_{i})+\sum_{i=-1}^{N-1}\sum_{j=-1}^{i-1}B(v_{i},v_{j})}$
+${+\sum_{i=-1}^{N-1}B(v_{i},\overline{v})+\sum_{j=-1}^{N-1}\sum_{i=-1}^{j-1}B(v_{i},v_{j})}$
+$=B(\overline{v},\overline{v})+\sum_{i=-1}^{N-1}B(v_{i},\overline{v})+\sum_{i=-1}^{N-1}B(\overline{v},v_{i})$
+$+\sum_{i=-1}^{N-1}\sum_{j=-1}^{i}B(v_{i},v_{j})+\sum_{j=-1}^{N-1}\sum_{i=-1}^{j-1}B(v_{i},v_{j})$
+$=\left[B(\overline{v},\overline{v})+B\left(w+\sum_{j=0}^{N-1}v_{i},\overline{v}\right)+B\left(\overline{v},w+\sum_{j=0}^{N-1}v_{j}\right)\right.$
+$\left.\quad\quad+B\left(v_{N-1},w+\sum_{j=0}^{N-1}v_{j}\right)+B\left(w+\sum_{j=0}^{N-2}v_{j},v_{N-1}\right)\right]$
+$+\sum_{i=-1}^{N-2}\left[B\left(v_{i},w+\sum_{j=0}^{i}v_{j}\right)+B\left(w+\sum_{j=0}^{i-1}v_{j},v_{i}\right)\right].$
+Looking at the last line, we can formally identify the first bracketed term as the one appearing in the equation for $\overline{v}$ and the $i$-th summand as the one appearing in the equation for $v_{i+1},\ i+1\in\{0,1,\dots,N-1\}.$ To show rigorously that $v$ given in (3.12) with $(v_{0},\dots,v_{N-1},\overline{v})$ solving (3.13) is a solution to (3.10) we need to check that $v_{0},\dots,v_{N-1}$ and $\overline{v}$ are sufficiently regular.
+The appropriate regularity class for $v$ in (3.12) to obtain a solution is given in the following definition, see also Remark 3.8 below.
 Definition 3.7.
@@
 Remark 3.8.
-We observe that $q_{i}>1$ for all $i\in\{0,\dots,N-1\}$ and is increasing in $i$. As an immediate consequence of 3.5, Sobolev embedding theorem and interpolation we have that
+We observe that $q_{i}>1$ for all $i\in\{0,\dots,N-1\}$ and is increasing in $i$. As an immediate consequence of Definition 3.7, Sobolev embedding theorem and interpolation we have that
 $v_{i}\in L^{p/2^{i}}(0,T;\mathbb{L}^{r_{i}})\cap C([0,T];H),\quad r_{i}=\frac{2r}{(i+1)(2-r)+2}.$
 In particular, $r_{i}>2$ for all $i\in\{0,\dots,N-1\}$ and is increasing in $i$.
-Therefore one can easily check that all the duality pairings in 3.5 are well defined.
+Therefore one can easily check that all the duality pairings in Definition 3.7 are well defined.
 Moreover, for all $i\in\{1,\dots,N-1\}$,
@@
 While for $\overline{v}$ we can use the standard interpolation inequality $L^{2}(0,T;H^{1})\cap L^{\infty}(0,T;L^{2})\subseteq L^{2/\theta}(0,T;H^{\theta})$ with $\theta=\frac{r-2}{r}\in(0,1)$ and the Sobolev embedding $H^{\theta}(\mathcal{O})\hookrightarrow L^{r}(\mathcal{O})$.
-In particular, if $(v_{0},\dots,v_{N-1},\overline{v})$ is a $(p,r)$ solution of (3.7), then, given $v:=\sum_{i=0}^{N-1}v_{i}+\overline{v}$, $u=v+w$ is a $r/2$-solution of (1.2) in the sense of 1.2.
-The following yields the well-posedness of (3.4) in the sense of 3.5.
+In particular, if $(v_{0},\dots,v_{N-1},\overline{v})$ is a $(p,r)$ solution of (3.13), then, given $v:=\sum_{i=0}^{N-1}v_{i}+\overline{v}$, $u=v+w$ is a $r/2$-solution of (1.2) in the sense of Definition 1.2.
+The following yields the well-posedness of (3.10) in the sense of Definition 3.7.
 Theorem 3.9.
-Let 3.4 be satisfied.
-For each $u_{\mathrm{in}}\in H$, $p\geq 2^{N}\frac{r}{r-2}$ there exists a unique $(p,r)$-solution $(v_{0},\dots,v_{N-1},\overline{v})$ of (3.7) in the sense of 3.5.
+Let 3.6 be satisfied.
+For each $u_{\mathrm{in}}\in H$, $p\geq 2^{N}\frac{r}{r-2}$ there exists a unique $(p,r)$-solution $(v_{0},\dots,v_{N-1},\overline{v})$ of (3.13) in the sense of Definition 3.7.
 Moreover $\overline{v}$ satisfies the energy relation
@@
 (3.19) $+2\int_{0}^{t}\langle(w(s)+\sum_{j=0}^{N-2}v_{j}(s))\cdot\nabla\overline{v}(s),v_{N-1}(s)\rangle\,\mathrm{d}s.$
-If $\left(u_{\mathrm{in}}^{n}\right)_{n\in{\mathbb{N}}}$ is a sequence in $H$ converging to $u_{\mathrm{in}}\in H$ and $(w^{n})_{n\in{\mathbb{N}}}$ is a sequence in $C([0,T];\mathbb{L}^{r})$ converging to $w\in C([0,T];\mathbb{L}^{r})$, then the corresponding unique solutions $((v_{0}^{n},\dots,v_{N-1}^{n},\overline{v}^{n}))_{n\in{\mathbb{N}}}$ converge to the corresponding solution $(v_{0},\dots,v_{N-1},\overline{v})$, each one in the topologies of 3.5.
+If $\left(u_{\mathrm{in}}^{n}\right)_{n\in{\mathbb{N}}}$ is a sequence in $H$ converging to $u_{\mathrm{in}}\in H$ and $(w^{n})_{n\in{\mathbb{N}}}$ is a sequence in $C([0,T];\mathbb{L}^{r})$ converging to $w\in C([0,T];\mathbb{L}^{r})$, then the corresponding unique solutions $((v_{0}^{n},\dots,v_{N-1}^{n},\overline{v}^{n}))_{n\in{\mathbb{N}}}$ converge to the corresponding solution $(v_{0},\dots,v_{N-1},\overline{v})$, each one in the topologies of Definition 3.7.
 Proof.
@@
 Step 1: Linear part of (3.13).
-We argue by induction and exploit maximal $L^{p}$ regularity techniques, see [53, Chapter 3].
-The existence and uniqueness of $v_{0}$ satisfying the corresponding PDE in the sense of 3.5 and the continuous dependence from data, i.e.
+We argue by induction and exploit maximal $L^{p}$ regularity techniques, see [55, Chapter 3].
+The existence and uniqueness of $v_{0}$ satisfying the corresponding PDE in the sense of Definition 3.7 and the continuous dependence from data, i.e.
 $w$ in the topology of $C([0,T];\mathbb{L}^{r})$, follows if
@@
 $\int_{0}^{T}\lVert B(w(s),w(s))\rVert_{X_{-1/2,A_{r/2}}}^{p}\,\mathrm{d}s$ $\leq\int_{0}^{T}\lVert w(s)\rVert_{\mathbb{L}^{r}}^{2p}\,\mathrm{d}s\leq T\lVert w\rVert_{C([0,T];\mathbb{L}^{r})}^{2p}.$
-Now assume we have already proved the existence and uniqueness of $(v_{i})_{i\in\{0,\dots,l-1\}},$ $l\leq N-1$ solving (3.7) in the sense of 3.5 and depending continuously from the data, i.e.
+Now assume we have already proved the existence and uniqueness of $(v_{i})_{i\in\{0,\dots,l-1\}},$ $l\leq N-1$ solving (3.13) in the sense of Definition 3.7 and depending continuously from the data, i.e.
 $w$ in the topology of $C([0,T];\mathbb{L}^{r})$.
-Let us check that there exists a unique $v_{l}$ solving the corresponding PDE in (3.7) in the sense of 3.5 and depending continuously from $w$ in the topology of $C([0,T];\mathbb{L}^{r})$.
+Let us check that there exists a unique $v_{l}$ solving the corresponding PDE in (3.13) in the sense of Definition 3.7 and depending continuously from $w$ in the topology of $C([0,T];\mathbb{L}^{r})$.
 Again, due to maximal $L^{p}$ regularity techniques, it is enough to show that
 $B(v_{l-1},v_{l-1})+B(v_{l-1},w+\sum_{j=0}^{l-2}v_{j})-B(w+\sum_{j=0}^{l-2}v_{j},v_{l-1})\in L^{p/2^{l}}(0,T;X_{-1/2,A_{q_{l}}}).$
-The claim is true, indeed due to 3.6
+The claim is true, indeed due to Remark 3.8
 $w,\ v_{i}\in L^{p/2^{l-1}}(0,T;\mathbb{L}^{r})\quad\text{if }i\in\{0,\dots,l-2\}$
@@
 $\widetilde{f}=-B\left(\overline{v}^{\left(2\right)}+\tilde{v},z\right)-B\left(z,\overline{v}^{\left(2\right)}+\tilde{v}\right).$
-By 2.6 and [29, Lemma 1.14], $\widetilde{f}\in L^{2}\left(0,T;V^{*}\right)$. Then, by Theorem 2.5,
+By Lemma 2.7, $\widetilde{f}\in L^{2}\left(0,T;V^{*}\right)$. Then, by Theorem 2.6,
 $\lVert z\left(t\right)\rVert^{2}+2\int_{0}^{t}\lVert\nabla z\left(s\right)\rVert_{L^{2}}^{2}\,\mathrm{d}s=2\int_{0}^{t}\langle z(s)\cdot\nabla z(s),\overline{v}^{\left(2\right)}(s)+\tilde{v}(s)\rangle\,\mathrm{d}s.$
-Again by 2.6, we have
+Again by Lemma 2.7, we have
 $\int_{0}^{t}\langle z(s)\cdot$ $\nabla z(s),\overline{v}^{\left(2\right)}(s)+\tilde{v}(s)\rangle\,\mathrm{d}s$
@@
 $\lVert z\left(t\right)\rVert^{2}+\int_{0}^{t}\lVert\nabla z\left(s\right)\rVert_{L^{2}}^{2}\,\mathrm{d}s\leq C\int_{0}^{t}\lVert z\left(s\right)\rVert^{2}\left(\lVert\overline{v}^{\left(2\right)}(s)\rVert_{\mathbb{L}^{4}}^{4}+\lVert\tilde{v}(s)\rVert_{\mathbb{L}^{r}}^{\frac{2r}{r-2}}\right)\,\mathrm{d}s.$
-We conclude $z=0$ by the Gronwall lemma, using (3.15) and the integrability properties of $\overline{v}^{\left(2\right)}$.
+We conclude $z=0$ by the Grönwall lemma, using (3.21) and the integrability properties of $\overline{v}^{\left(2\right)}$.
 Step 4: Existence.
@@
 $B\left(\overline{v}^{n},\tilde{v}\right),B\left(\tilde{v},\overline{v}^{n}\right),\overline{f}\in L^{2}\left(0,T;V^{*}\right)$
-by 2.6 and (3.14).
+by Lemma 2.7 and (3.20).
 Then let us investigate the convergence of $\left(\overline{v}^{n}\right)$. First, let us prove a bound. From the previous identity and Theorem 2.6 we get
@@
 $=\lVert u_{\mathrm{in}}\rVert^{2}+2\int_{0}^{t}\left(b\left(\overline{v}^{n},\overline{v}^{n+1},\tilde{v}\right)+b\left(\tilde{v},\overline{v}^{n+1},\overline{v}^{n}\right)+\langle\overline{f},\overline{v}^{n+1}\rangle\right)\left(s\right)\,\mathrm{d}s.$
-It gives us (using 2.6 and (3.14))
+It gives us using Lemma 2.7 and (3.20)
 $\lVert\overline{v}^{n+1}\left(t\right)\rVert^{2}+\int_{0}^{t}\lVert\nabla\overline{v}^{n+1}\left(s\right)\rVert_{L^{2}}^{2}\,\mathrm{d}s$
@@
 Set $z_{n}=\overline{v}^{n}-\overline{v}^{n-1}$, for $n\geq 1$. From the identity above,
-$\left\langle z_{n+1}\left(t\right),\overline{\phi}\right\rangle-\int_{0}^{t}\left(b\left(\overline{v}^{n+1},\overline{\phi},\overline{v}^{n+1}\right)-b\left(\overline{v}^{n},\phi,\overline{v}^{n}\right)\right)\left(s\right)\,\mathrm{d}s$
+$\left\langle z_{n+1}\left(t\right),\overline{\phi}\right\rangle-\int_{0}^{t}\left(b\left(\overline{v}^{n+1},\overline{\phi},\overline{v}^{n+1}\right)-b\left(\overline{v}^{n},\overline{\phi},\overline{v}^{n}\right)\right)\left(s\right)\,\mathrm{d}s$
 $=-\int_{0}^{t}\left\langle z_{n+1}\left(s\right),A\overline{\phi}\right\rangle\,\mathrm{d}s-\int_{0}^{t}\left\langle\left(B\left(\overline{v}^{n},\tilde{v}\right)-B\left(\overline{v}^{n-1},\tilde{v}\right)\right)\left(s\right),\overline{\phi}\right\rangle\,\mathrm{d}s$
@@
 $\left\langle z_{n+1}\left(t\right),\overline{\phi}\right\rangle-\int_{0}^{t}b\left(z_{n+1}\left(s\right),\overline{\phi},z_{n+1}\left(s\right)\right)\,\mathrm{d}s$
-$=-\int_{0}^{t}\left\langle z_{n+1}\left(s\right),A\overline{\phi}\right\rangle\,\mathrm{d}s-\int_{0}^{t}\left\langle\left(B\left(z_{n},\tilde{v}\right)+B\left(\tilde{v},z_{n}\right)\right)\left(s\right),\phi\right\rangle\,\mathrm{d}s$
+$=-\int_{0}^{t}\left\langle z_{n+1}\left(s\right),A\overline{\phi}\right\rangle\,\mathrm{d}s-\int_{0}^{t}\left\langle\left(B\left(z_{n},\tilde{v}\right)+B\left(\tilde{v},z_{n}\right)\right)\left(s\right),\overline{\phi}\right\rangle\,\mathrm{d}s$
 $+\int_{0}^{t}\left(b\left(\overline{v}^{n},\overline{\phi},z_{n+1}\right)+b\left(z_{n+1},\overline{\phi},\overline{v}^{n}\right)\right)\left(s\right)\,\mathrm{d}s.$
@@
 $+2\int_{0}^{t}b\left(z_{n+1},z_{n+1},\overline{v}^{n}\right)\left(s\right)\,\mathrm{d}s.$
-As above, thanks to [29, Lemma 1.14] we deduce that
+As above, thanks to Lemma 2.7 we deduce that
 $\int_{0}^{t}\lvert b\left(z_{n+1},z_{n+1},\overline{v}^{n}\right)(s)\rvert\,\mathrm{d}s\leq\frac{1}{4}\int_{0}^{t}\lVert z_{n+1}(s)\rVert_{V}^{2}\,\mathrm{d}s+C\int_{0}^{t}\lVert z_{n+1}(s)\rVert^{2}\lVert\overline{v}^{n}(s)\rVert_{\mathbb{L}^{4}}^{4}\,\mathrm{d}s.$
@@
 $+\frac{1}{4}\int_{0}^{t}\lVert z_{n}\left(s\right)\rVert_{V}^{2}\,\mathrm{d}s+C\int_{0}^{t}\lVert z_{n}\left(s\right)\rVert^{2}\lVert\tilde{v}\left(s\right)\rVert_{\mathbb{L}^{r}}^{\frac{2r}{r-2}}\,\mathrm{d}s.$
-Now we work under the bounds (3.17) and deduce, using the Gronwall lemma, for $\overline{T}$, depending only from $\lVert u_{\mathrm{in}}\rVert$ and $\lVert\tilde{v}\rVert_{L^{\frac{2r}{r-2}}(0,T;\mathbb{L}^{r})}$, possibly smaller than the previous one,
-$\sup_{t\in\left[0,\overline{T}\right]}\lVert z_{n+1}\left(t\right)\rVert^{2}+\int_{0}^{\overline{T}}\lVert z_{n+1}\left(s\right)\rVert_{V}^{2}\,\mathrm{d}s$
-$\leq\frac{1}{2}\left(\sup_{t\in\left[0,\overline{T}\right]}\lVert z_{n}\left(t\right)\rVert^{2}+\int_{0}^{\overline{T}}\lVert z_{n}\left(s\right)\rVert_{V}^{2}\,\mathrm{d}s\right).$
-It implies that the sequence $\left(\overline{v}^{n}\right)$ is Cauchy in $C\left(\left[0,\overline{T}\right];H\right)\cap L^{2}\left(0,\overline{T};V\right)$.
-The limit $\overline{v}$ has the right regularity to be a weak solution and satisfies the weak formulation; in the identity above for $\overline{v}^{n+1}$ and $\overline{v}^{n}$ we may prove that
-$\int_{0}^{t}b\left(\overline{v}^{n+1}\left(s\right),\overline{\phi},\overline{v}^{n+1}\left(s\right)\right)\,\mathrm{d}s$ $\rightarrow\int_{0}^{t}b\left(\overline{v}\left(s\right),\overline{\phi},\overline{v}\left(s\right)\right)\,\mathrm{d}s$
-$\int_{0}^{t}b\left(\overline{v}^{n}\left(s\right),\overline{\phi},\tilde{v}\left(s\right)\right)\,\mathrm{d}s$ $\rightarrow\int_{0}^{t}b\left(\overline{v}\left(s\right),\overline{\phi},\tilde{v}\left(s\right)\right)\,\mathrm{d}s$
-$\int_{0}^{t}b\left(\tilde{v}\left(s\right),\overline{\phi},\overline{v}^{n}\left(s\right)\right)\,\mathrm{d}s$ $\rightarrow\int_{0}^{t}b\left(\tilde{v}\left(s\right),\overline{\phi},\overline{v}\left(s\right)\right)\,\mathrm{d}s.$
-All these convergences can be proved easily by recalling the definition of $b$.
-Similarly, we can pass to the limit in the energy identity.
+Now we work under the bounds (3.23) and deduce, using the Grönwall lemma, for $\overline{T}$, depending only from $\lVert u_{\mathrm{in}}\rVert$ and $\lVert\tilde{v}\rVert_{L^{\frac{2r}{r-2}}(0,T;\mathbb{L}^{r})}$, possibly smaller than the previous one,
+$\sup_{t\in\left[0,\overline{T}\right]}\lVert z_{n+1}\left(t\right)\rVert^{2}+\int_{0}^{\overline{T}}\lVert z_{n+1}\left(s\right)\rVert_{V}^{2}\,\mathrm{d}s\leq\frac{1}{2}\left(\sup_{t\in\left[0,\overline{T}\right]}\lVert z_{n}\left(t\right)\rVert^{2}+\int_{0}^{\overline{T}}\lVert z_{n}\left(s\right)\rVert_{V}^{2}\,\mathrm{d}s\right).$
+Now we can proceed as in the second step of the proof of [1, Theorem 3.3], showing that the sequence $\left(\overline{v}^{n}\right)$ is Cauchy in $C\left(\left[0,\overline{T}\right];H\right)\cap L^{2}\left(0,\overline{T};V\right)$.
+Its limit $\overline{v}$ is a weak solution of (3.22) on $[0,\overline{T}]$ and , hence, by the previous step, it is the unique solution.
+We refer the reader to [1, Theorem 3.3] for further details.
 After proving existence and uniqueness in $[0,\overline{T}]$ we can reiterate the existence procedure and in a finite number of steps cover the interval $[0,T]$.
@@
 (3.24) $+\int_{0}^{t}\langle\overline{f}^{n}(s)-\overline{f}(s),z^{n}(s)\rangle\,\mathrm{d}s.$
-We can easily bound the right-hand side of relation (3.2) by Young’s inequality and Hölder’s inequality obtaining
+Thanks to 2.7, we can easily bound the right-hand side of relation (3.2) by Young’s inequality and Hölder’s inequality obtaining
 $\frac{1}{2}\lVert z^{n}(t)\rVert^{2}+\frac{1}{2}\int_{0}^{t}\lVert\nabla z^{n}(s)\rVert_{L^{2}}^{2}\,\mathrm{d}s$ $\leq\frac{1}{2}\lVert u_{\mathrm{in}}^{n}-u_{\mathrm{in}}\rVert^{2}+\int_{0}^{t}\lVert\overline{f}^{n}(s)-\overline{f}(s)\rVert_{V^{*}}^{2}\,\mathrm{d}s$
@@
 (3.25) $+C\lVert\widetilde{v}^{n}-\widetilde{v}\rVert_{L^{\frac{2r}{r-2}}(0,T;\mathbb{L}^{r})}^{2}\lVert\overline{v}\rVert_{C([0,T];H)}^{\frac{2(r-2)}{r}}\lVert\overline{v}\rVert_{L^{2}(0,T;V)}^{\frac{4}{r}}.$
-Applying Gronwall’s inequality to relation (3.2) the claimed continuity follows. ∎
+Applying Grönwall’s inequality to relation (3.2) the claimed continuity follows. ∎
 Remark 3.10.
@@
 paths in $C(0,T;H)\cap L^{\frac{2q}{q-1}}(0,T;\mathbb{L}^{2q})$ for each $r\in(1,q_{\mathcal{H}})$, it is progressively measurable with respect to these topologies.
-Combining 3.1, Theorem 3.7 and 3.8 we get immediately the existence of a $q-$solution of equation (1.2) in the sense of 1.2 for each $q\in(1,q_{\mathcal{H}})$.
+Combining Proposition 3.1, Lemma 3.5, Theorem 3.9 and Remark 3.10 we get immediately the existence of a $q$-solution of equation (1.2) in the sense of Definition 1.2 for each $q\in(1,q_{\mathcal{H}})$.
 3.3. Proof of Theorem 1.3(1)
-As discussed above, the results of subsection 3.1, subsection 3.2 provide the existence of a $q$-solution of equation (1.2) in the sense of 1.2 for each $q\in(1,q_{\mathcal{H}})$, moreover such a solution is adapted with paths in $C([0,T];H)$ due to 3.6.
-In order to conclude the proof of Theorem 1.31 it is enough to show the following:
+As discussed above, the results of Section 3.1, Section 3.2 provide the existence of a $q$-solution of equation (1.2) in the sense of Definition 1.2 for each $q\in(1,q_{\mathcal{H}})$, moreover such a solution is adapted with paths in $C([0,T];H)$ due to Remark 3.8.
+Here we are left to discuss the problem of uniqueness.
+In order to reach our goal, we start providing a lemma which shows equivalence of $q$-solutions in the sense of 1.2 and those of the form $u=w_{g}+v$ as described by 3.11 below.
+Then we conclude the proof by providing uniqueness for solutions of the form $u=w_{g}+v$.
+Lemma 3.11.
+Let $T<\infty$, $u_{\mathrm{in}}\in L^{0}_{\mathscr{F}_{0}}(\Omega;\mathbb{L}^{2})$, $q\in(1,q_{\mathcal{H}})$ and assuming 1.1.
+Then $u$ is a $q$-solution to (1.2) in the sense of Definition 1.2 with paths in $C([0,T];\mathbb{L}^{2})$ $\mathbf{P}-a.s.$ if and only if $v:=u-w_{g}$ is progressively measurable with paths in $L^{2q^{\prime}}(0,T;\mathbb{L}^{2q})\cap C([0,T];\mathbb{L}^{2})$ $\mathbf{P}-a.s.$ and it solves, for all divergence-free $\varphi\in C^{\infty}(\mathcal{O};{\mathbb{R}}^{2})$ such that $\varphi=0$ on $\Gamma_{b}\cup\Gamma_{u}$ and a.e.
+$t\in(0,T)$,
+(3.26) $\int_{\mathcal{O}}v(x,t)\varphi(x)\,\mathrm{d}x-\int_{\mathcal{O}}u_{\mathrm{in}}(x)\varphi(x)\,\mathrm{d}x$
+$\qquad\qquad=\int_{0}^{t}\int_{\mathcal{O}}\big{(}v\cdot\Delta\varphi+[(v+w_{g})\otimes(v+w_{g})]:\nabla\varphi\big{)}\,\mathrm{d}x\,\mathrm{d}s.$
+Proof.
+The proof is a trivial consequence of the notion of $q$-solution in 1.2, the regularity of $w_{g}$ in case of 1.1, i.e.
+3.1, and the equivalence between weak and mild solutions for the linear stochastic problem, see 3.5.
+∎
+We call a $v$ progressively measurable with paths in $L^{2q^{\prime}}(0,T;\mathbb{L}^{2q})\cap C([0,T];\mathbb{L}^{2})$ $\mathbf{P}-a.s.$ satisfying (3.26) in the sense of Lemma 3.11 an auxiliary $q$-solution of (3.10).
+The uniqueness of auxiliary $q$-solutions and their independence on $q$ is the content of the following result, which, combined with Lemma 3.11, concludes the proof of the first item in Theorem 1.3.
 Proposition 3.12 (Uniqueness).
-Let $u_{1}$ be a $q_{1}$-solution of (1.2) and $u_{2}$ be a $q_{2}$-solution of (1.2). Then $u_{1}\equiv u_{2}$.
-The uniqueness result in the Ladyzhenskaya–Prodi–Serrin class of 3.9 might be known to experts.
+Let $v_{1}$ be an auxiliary $q_{1}$-solution of (3.10) and $v_{2}$ be an auxiliary $q_{2}$-solution of (3.10). Then $v_{1}\equiv v_{2}$.
+The uniqueness result in the Ladyzhenskaya–Prodi–Serrin class of Proposition 3.12 might be known to experts.
 Here, for completeness, we provide a relatively short proof relying on maximal $L^{p}$-regularity techniques which seem not standard even in the absence of noise.
-Proof of 3.9.
-Recall that $u_{i}=w_{g}+v_{i}$ where $v_{i}$ solves the modified Navier-Stokes equations for $i\in\{1,2\}$ by 1.2. We split the proof into two cases.
+Proof of Proposition 3.12.
+We split the proof into two cases.
 […]
@@
 $=\int_{0}^{t}b(\delta(s),\varphi,v_{1}(s)+w_{g}(s))\,\mathrm{d}s+\int_{0}^{t}b(v_{2}(s)+w_{g}(s),\varphi,\delta(s))\,\mathrm{d}s.$
-As $v_{i}\in L^{2q^{\prime}}(0,T;L^{2q}(\mathcal{O};{\mathbb{R}}^{2}))$ for $i\in\{1,2\}$ by 1.2, we obtain
+As $v_{i}\in L^{2q^{\prime}}(0,T;L^{2q}(\mathcal{O};{\mathbb{R}}^{2}))$ for $i\in\{1,2\}$, we obtain
 $B(\delta,v_{1}+w_{g}),B(v_{2}+w_{g},\delta)\in L^{q^{\prime}}(0,T;X_{-1/2,A_{q}})\ \ \mathbf{P}-a.s.$
@@
 $\subset C([0,T];B^{1-2/q^{\prime}}_{q,q^{\prime}}(\mathcal{O};{\mathbb{R}}^{2}))\ \ \mathbf{P}-a.s.$
-where in the last step we used the trace embedding [53, Theorem 3.4.8] applied with $A=A_{q}$.
+where in the last step we used the trace embedding [55, Theorem 3.4.8] applied with $A=A_{q}$.
 By real interpolation (see e.g. [8, Chapter 6]), we obtain
@@
 for all $f$ for which the right-hand side is finite.
-Hence, by maximal $L^{q}$-regularity of $A_{q}$, again the trace embedding [53, Theorem 3.4.8] as well as the Hölder inequality, there exists a constant $C_{0}>0$ independent of $v_{1},v_{2}$ and $\delta$ such that, for all $t\in[0,T]$ and $\mathbf{P}-a.s.$,
+Hence, by maximal $L^{q}$-regularity of $A_{q}$, again the trace embedding [55, Theorem 3.4.8] as well as the Hölder inequality, there exists a constant $C_{0}>0$ independent of $v_{1},v_{2}$ and $\delta$ such that, for all $t\in[0,T]$ and $\mathbf{P}-a.s.$,
 $\sup_{r\in[0,t]}\|\delta(r)\|_{B^{1-2/q^{\prime}}_{q,q^{\prime}}}^{q^{\prime}}+\int_{0}^{t}\|\delta(r)\|_{H^{1,q}}^{q^{\prime}}\,\mathrm{d}r$
@@
 where in the last step we used the Young inequality and (3.27).
-Now the conclusion follows from the Grownall lemma and the integrability conditions on $v_{1},v_{2}$ due to 1.2 and $w_{g}\in C([0,\infty);L^{2q})$ for all $q<q_{\mathcal{H}}$ by 3.1.
+Now the conclusion follows from the Grönwall lemma and the integrability conditions on $v_{1},v_{2}$ and $w_{g}\in C([0,\infty);L^{2q})$ for all $q<q_{\mathcal{H}}$ by Proposition 3.1.
 […]
 Lemma 3.13 (Compatibility).
-Let $w\in C([0,T];\mathbb{L}^{r})$ for some $r\in(2,4)$ and $2<\widetilde{r}\leq r$. If $(v_{0},\dots,v_{N-1},\overline{v})$ is a solution of (3.7) in the sense of 3.5 with
+Let $w\in C([0,T];\mathbb{L}^{r})$ for some $r\in(2,4)$ and $2<\widetilde{r}\leq r$. If $(v_{0},\dots,v_{N-1},\overline{v})$ is a solution of (3.13) in the sense of Definition 3.7 with
 $p\geq 2^{N}\frac{r}{r-2},\qquad q_{i}=\frac{2r}{r+2+(i+1)(2-r)}$
-and $(\widetilde{v}_{0},\dots,\widetilde{v}_{N-1},\widehat{v})$ is a solution of (3.4) in the sense of 3.5 with
+and $(\widetilde{v}_{0},\dots,\widetilde{v}_{\widetilde{N}-1},\widetilde{\overline{v}})$ is a solution of (3.10) in the sense of Definition 3.7 with
 $\widetilde{p}\geq 2^{\widetilde{N}}\frac{\widetilde{r}}{\widetilde{r}-2},\qquad\widetilde{q}_{i}=\frac{2\widetilde{r}}{\widetilde{r}+2+(i+1)(2-\widetilde{r})}$
@@
 and we are left to show that $\overline{v}=\sum_{i=N}^{\widetilde{N}-1}\widetilde{v}_{i}+\widetilde{\overline{v}}=:\widehat{v}$.
-Due to previous steps we can assume that $v$ is $(\widetilde{p},r)$ solution since $\widetilde{p}\geq 2^{\widetilde{N}}\frac{\widetilde{r}}{\widetilde{r}-2}>2^{{N}}\frac{{r}}{{r}-2}$ We observe that due to 3.5 and 3.6,
+Due to previous steps we can assume that $v$ is $(\widetilde{p},r)$ solution since $\widetilde{p}\geq 2^{\widetilde{N}}\frac{\widetilde{r}}{\widetilde{r}-2}>2^{{N}}\frac{{r}}{{r}-2}$.
+We observe that due to Definition 3.7 and Remark 3.8,
 (3.28) $\overline{v},\widehat{v}\in C([0,T];H)\cap L^{\frac{2\widetilde{r}}{\widetilde{r}-2}}(0,T;\mathbb{L}^{\widetilde{r}}),\ f:=w+\sum_{i=0}^{N-1}v_{i}\in L^{\frac{2\widetilde{r}}{\widetilde{r}-2}}(0,T;\mathbb{L}^{\widetilde{r}}).$
@@
 $B(\delta,\overline{v}+f),\ B(\widehat{v}+f,\delta)\in L^{\widetilde{q}^{\prime}}(0,T;X_{-1/2,A_{\widetilde{q}}}).$
-Now the proof proceeds as in the first case of 3.9 and we omit the details. ∎
+Now the proof proceeds as in the first case of Proposition 3.12 and we omit the details. ∎
 4. Interior regularity
-As announced below subsection 1.1, we prove Theorem 1.3(2).
-To this end, we first prove the interior regularity of $w_{g}$ and afterwards the one of $v$ by exploiting the decomposition introduced in subsection 3.2.
+As announced at the end of Section 1.1, we prove Theorem 1.3(2).
+To this end, we first prove the interior regularity of $w_{g}$ and afterwards the one of $v$ by exploiting the decomposition introduced in Section 3.2.
 4.1. Stokes equations
-We start showing a lemma concerning the relation between the mild and the weak formulation of (1.3) defined below.
-Definition 4.1.
-Let 1.1 be satisfied. A stochastic process $w$ is a weak solution of (1.3) if it is $\mathcal{F}$-progressively measurable with $\mathbf{P}-a.s.$ paths in
-$w_{g}\in C(0,T;\mathbb{L}^{r})$
-for some $r\geq 2$, and $\mathbf{P}-a.s.$ for all $\phi\in\mathsf{D}(A)$ and $t\in[0,T]$,
-(4.1) $\langle w_{g}(t),\phi\rangle=-\int_{0}^{t}\langle w_{g}(s),A\phi\rangle\,\mathrm{d}s-\langle g,\widehat{n}\cdot\nabla\phi\rangle_{H^{-s}(\Gamma_{u}),H^{s}(\Gamma_{u})}W^{\mathcal{H}}_{t}.$
-Since $g$ is time-independent, the last term in (4.1) can be rewritten as a stochastic integral as
-$\langle g,\widehat{n}\cdot\nabla\phi\rangle_{H^{-s}(\Gamma_{u}),H^{s}(\Gamma_{u})}W^{\mathcal{H}}_{t}=\int_{0}^{t}\langle g,\widehat{n}\cdot\nabla\phi\rangle_{H^{-s}(\Gamma_{u}),H^{s}(\Gamma_{u})}\,\mathrm{d}W^{\mathcal{H}}_{s}.$
-Lemma 4.2.
-Let 1.1 be satisfied. There exists a unique weak solution of (1.3) in the sense of 4.1 and it is given by the formula (3.1).
-Proof.
-We split the proof into two steps.
-Step 1: There exists a unique weak solution of (1.3) and it is necessarily given by the mild formula (3.1).
-Let $\psi\in C^{1}([0,T];\mathsf{D}(A))$.
-Arguing as in the first step of the proof of [29, Theorem 1.7], see also [44, Lemma 3], one can readily check that $w_{g}$ satisfies
-$\langle w_{g}(t),\psi(t)\rangle=$ $\int_{0}^{t}\langle w_{g}(s),\partial_{s}\psi(s)\rangle\,\mathrm{d}s-\int_{0}^{t}\langle w_{g}(s),A\psi(s)\rangle\,\mathrm{d}s$
-(4.2) $-\int_{0}^{t}\langle g,\widehat{n}\cdot\nabla\psi(s)\rangle_{H^{-s}(\Gamma_{u}),H^{s}(\Gamma_{u})}\,\mathrm{d}W^{\mathcal{H}}_{s}$
-for each $t\in[0,T],\ \mathbf{P}-a.s.$ The stochastic integral in the relation above is well-defined as a real-valued stochastic integral.
-Recalling that $(W^{\mathcal{H}}_{t})_{t\geq 0}$ is a $U$-cylindrical fractional Brownian motion we observe that $\langle g,\widehat{n}\cdot\nabla\psi(s)\rangle_{H^{-s}(\Gamma_{u}),H^{s}(\Gamma_{u})}$ is given by the linear operator on $U$
-$h^{\prime}\mapsto\langle gh^{\prime},\widehat{n}\cdot\nabla\psi(s)\rangle_{H^{-s}(\Gamma_{u}),H^{s}(\Gamma_{u})}=L_{\psi}(gh^{\prime}),$
-where $L_{\psi}:=\langle\cdot,\widehat{n}\cdot\nabla\psi(\cdot)\rangle_{H^{-s}(\Gamma_{u}),H^{s}(\Gamma_{u})}.$ By the ideal property of the Hilbert-Schmidt operators we have $\mathscr{L}_{2}(U,H^{-s}(\Gamma_{u}))=H^{-s}(\Gamma_{u};U)$ and obtain that
-$\|\langle g,\widehat{n}\cdot\nabla\psi(s)\rangle_{H^{-s}(\Gamma_{u}),H^{s}(\Gamma_{u})}\|_{U^{*}}\lesssim\|g\|_{H^{-s}(\Gamma_{u})}\|\nabla\psi(s)\|_{H^{s}(\Gamma_{u})}~{}~{}\text{a.e.
-on }\Omega\times(0,T).$
-In conclusion, the stochastic integral in (4.1) is well-defined as a real-valued one, since
-$\mathbf{E}\Big{|}\int_{0}^{t}\langle g,\widehat{n}\cdot\nabla\psi(s)\rangle_{H^{-s}(\Gamma_{u}),H^{s}(\Gamma_{u})}~{}\mathrm{d}W^{\mathcal{H}}_{s}\Big{|}^{2}$
-$\leq\mathcal{H}(2\mathcal{H}-1)\int_{0}^{t}\int_{0}^{t}\|\langle g,\widehat{n}\cdot\nabla\psi(s)\rangle\|_{U^{*}}\|\langle g,\widehat{n}\cdot\nabla\psi(v)\rangle\|_{U^{*}}|s-v|^{2\mathcal{H}-2}~{}\mathrm{d}s~{}\mathrm{d}v$
-$\leq\mathcal{H}(2\mathcal{H}-1)\|g\|^{2}_{H^{-s}(\Gamma_{u};U)}\int_{0}^{t}\int_{0}^{t}\|\nabla\psi(s)\|_{H^{s}(\Gamma_{u})}\|\nabla\psi(v)\|_{H^{s}(\Gamma_{u})}|s-v|^{2\mathcal{H}-2}~{}\mathrm{d}s~{}\mathrm{d}v,$
-which is finite since $\psi\in C^{1}([0,T];\mathsf{D}(A))$ and $\mathcal{H}>1/2$.
-Now consider $\phi\in\mathsf{D}(A^{2})$ and use $\psi_{t}(s)=S(t-s)\phi,\ s\in[0,t]$ as test function in (4.1) obtaining
-(4.3) $\langle w_{g}(t),\phi\rangle=$ $-\int_{0}^{t}\langle g,\widehat{n}\cdot\nabla S(t-s)\phi\rangle_{H^{-s}(\Gamma_{u}),H^{s}(\Gamma_{u})}\,\mathrm{d}W^{\mathcal{H}}_{s}.$
-Recalling the definition of the Dirichlet map $\mathcal{D}$, (4.3) can be rewritten as
-(4.4) $\langle w(t),\phi\rangle=$ $\int_{0}^{t}\langle\mathcal{D}[g],AS(t-s)\phi\rangle\,\mathrm{d}W^{\mathcal{H}}_{s}.$
-Then, exploiting the self-adjointness property of $S$ and $A$ we have that weak solutions of (1.3) satisfy the mild formulation. Therefore they are unique.
-Step 2: The mild formula (3.1) is a weak solution of (1.3) in the sense of 4.1.
-We begin by noticing that $w_{g}$ has the required regularity due to 3.1.
-Let us test our mild formulation (3.1) against functions $\phi\in\mathsf{D}(A^{2})$.
-It holds, exploiting self-adjointness property of $S$ and $A$
-$\langle w(t),\phi\rangle$ $=\int_{0}^{t}\langle\mathcal{D}[g],AS(t-s)\phi\rangle\,\mathrm{d}W^{\mathcal{H}}_{s}$
-$=\int_{0}^{t}\langle g,\widehat{n}\cdot\nabla S(t-s)\phi\rangle_{H^{-s}(\Gamma_{u}),H^{s}(\Gamma_{u})}\,\mathrm{d}W^{\mathcal{H}}_{s}\quad\mathbf{P}-a.s.,$
-where in the last step we used the definition of Dirichlet map. To complete the proof of this step it is enough to show that
-$\int_{0}^{t}\langle g,\widehat{n}\cdot\nabla S(t-s)\phi\rangle_{H^{-s}(\Gamma_{u}),H^{s}(\Gamma_{u})}\,\mathrm{d}W^{\mathcal{H}}_{s}=-\int_{0}^{t}\langle w_{g}(s),A\phi\rangle\,\mathrm{d}s$
-(4.5) $+\langle g,\widehat{n}\cdot\nabla\phi\rangle_{H^{-s}(\Gamma_{u}),H^{s}(\Gamma_{u})}\,W^{\mathcal{H}}_{t}\quad\mathbf{P}-a.s.$
-The relation (4.1) is true. Indeed,
-(4.6) $\int_{0}^{t}\langle w_{g}(s),A\phi\rangle\,\mathrm{d}s$ $=\int_{0}^{t}\,\mathrm{d}s\int_{0}^{s}\langle\mathcal{D}[g],S(s-\tau)A^{2}\phi\rangle\,\mathrm{d}W^{\mathcal{H}}(\tau)\quad\mathbf{P}-a.s.$
-The double integrals in (4.6) can be exchanged via stochastic Fubini’s Theorem, see [48, 3]. Therefore the double integral in the right-hand side of (4.6) can be rewritten as
-$\int_{0}^{t}\,\mathrm{d}s\int_{0}^{s}\langle\mathcal{D}[g],S(s-\tau)A^{2}\phi\rangle\,\mathrm{d}W^{\mathcal{H}}_{\tau}$
-$\qquad=\int_{0}^{t}\,\mathrm{d}W^{\mathcal{H}}(\tau)\int_{\tau}^{t}\langle\mathcal{D}[g],S(s-\tau)A^{2}\phi\rangle\,\mathrm{d}s$
-$\qquad=\langle\mathcal{D}[g],A\phi\rangle W^{\mathcal{H}}_{t}-\int_{0}^{t}\langle\mathcal{D}[g],AS(t-\tau)\phi\rangle\,\mathrm{d}W^{\mathcal{H}}_{\tau}$
-$\qquad=\langle g,\widehat{n}\cdot\nabla\phi\rangle_{H^{-s}(\Gamma_{u}),H^{s}(\Gamma_{u})}\,W^{\mathcal{H}}_{t}$
-$\qquad-\int_{0}^{t}\langle g,\widehat{n}\cdot\nabla S(t-\tau)\phi\rangle_{H^{-s}(\Gamma_{u}),H^{s}(\Gamma_{u})}\,\mathrm{d}W^{\mathcal{H}}_{\tau}\quad\mathbf{P}-a.s.$
-Inserting this expression in (4.6), (4.1) holds and the proof is complete. ∎
-Let $(v_{1},\dots,v_{N-1},\overline{v})$ be the $(p,r)$-solution to (3.7) as defined in 3.5 given by Theorem 3.7. Let $N_{0}$ be the $\mathbf{P}$ null measure set where at least one between
+Let $(v_{0},v_{1},\dots,v_{N-1},\overline{v})$ be the $(p,r)$-solution to (3.13) as defined in Definition 3.7 given by Theorem 3.9.
+Let $N_{0}$ be the $\mathbf{P}$ null measure set where at least one between
 $w_{g}$ $\notin C([0,T];\mathbb{L}^{r}),\quad v_{i}\notin W^{1,p/2^{i}}(0,T;X_{-1/2,A_{q_{i}}})\cap L^{p/2^{i}}(0,T;X_{1/2,A_{q_{i}}}),$
@@
 In the following, we will work pathwise in $\Omega\setminus N_{0}$ even if not specified.
-Thanks to the weak formulation guaranteed by 4.2 we can easily obtain the interior regularity of the linear stochastic problem (1.3).
+Thanks to the weak formulation guaranteed by Lemma 3.5 we can easily obtain the interior regularity of the linear stochastic problem (1.4).
 Indeed, we are exactly in the same position of [1, Corollary 4.4] and the following holds.
 We omit the proof as it follows verbatim the one of [1, Corollary 4.4].
-Corollary 4.3.
+Lemma 4.1.
 Let 1.1 be satisfied.
-Let $w_{g}$ be the unique weak solution of (1.3) in the sense of 4.1.
+Let $w_{g}$ be the unique weak solution of (1.4) in the sense of Definition 3.4.
 Then, for all $0<t_{1}\leq t_{2}<T,$ $x_{0}\in\mathcal{O}$, $\rho>0$ such that $\mathrm{dist}({B(x_{0},\rho)},\partial\mathcal{O})>0$,
@@
 In contrast to [1], as $w_{g}\notin C([0,T];\mathbb{L}^{4})$, we cannot work directly on $v$.
-However, recalling that the solution $v$ to (1.2) proven in subsection 3.3 satisfies
+However, recalling that the solution $v$ to (1.2) proven in Section 3.3 satisfies
 (4.1) $v=\sum_{i=0}^{N-1}v_{i}+\overline{v}$
-where, again, $(v_{1},\dots,v_{N-1},\overline{v})$ is the $(p,r)$-solution to (3.7), cf., subsection 3.2.
-The advantage of having the splitting (4.7) at our disposal is that $v_{i}$ satisfies a linear problem where the forcing terms only depend on $v_{1},\dots,v_{i-1}$.
-Thus, by 4.3 and an induction argument, we can prove that $v_{i}$ is smooth inside $(0,T)\times\mathcal{O}$.
-While to prove the corresponding statement for $\overline{v}$, we can exploit that $\overline{v}$ is a Leray solution (i.e., it has finite energy) and therefore the Serrin regularization can be adjusted to our situation.
-We begin with analyzing the interior regularity of $v_{i}$ for $i\in\{1,\dots,N-1\}$.
+where, again, $(v_{0},\dots,v_{N-1},\overline{v})$ is the $(p,r)$-solution to (3.13), cf.
+Section 3.2.
+The advantage of having the splitting (4.1) at our disposal is that $v_{i}$ satisfies a linear problem where the forcing terms only depend on $v_{0},\dots,v_{i-1}$.
+Thus, by Lemma 4.1 and an induction argument, we can prove that $v_{i}$ is smooth inside $(0,T)\times\mathcal{O}$.
+While to prove the corresponding statement for $\overline{v}$, we can exploit that $\overline{v}$ is a Leray solution (i.e.
+it has finite energy) and therefore the Serrin regularization can be adjusted to our situation.
+We begin with analyzing the interior regularity of $v_{i}$ for $i\in\{0,\dots,N-1\}$.
 Lemma 4.2.
 Let 1.1, $r\in(2,4)$ and $p\geq 2^{N}\frac{r}{r-2}$.
-Let $(v_{1},\dots,v_{N-1},\overline{v})$ be the $(p,r)$-solution of (3.7) in the sense of 3.5.
+Let $(v_{0},\dots,v_{N-1},\overline{v})$ be the $(p,r)$-solution of (3.13) in the sense of Definition 3.7.
 Then for all $i\in\{0,\dots,N-1\},$ $0<t_{1}\leq t_{2}<T,$ $x_{0}\in\mathcal{O}$, $\rho>0$ such that $\mathrm{dist}({B(x_{0},\rho)},\partial\mathcal{O})>0$,
@@
 As in the first step of the proof of Theorem 3.9 we argue by induction exploiting strongly the linear and triangle structure of (3.13).
-Before starting we observe that, by [53, Theorem 3.4.8], it follows that
+Before starting we observe that, by [55, Theorem 3.4.8], it follows that
 (4.2) $v_{i}\in C([0,T];\mathbb{L}^{q_{i}})\cap L^{p/{2^{i}}}(0,T;X_{1/2,A_{q_{i}}}).$
@@
 $h_{0}$ $=\partial_{t}\psi_{0}\omega_{0}-2\nabla\psi_{0}\cdot\nabla\omega_{0}-\Delta\psi_{0}\omega_{0}-\psi_{0}w_{g}\cdot\nabla\omega_{w}.$
-Due to 4.3
+Due to Lemma 4.1
 $h_{0}\in L^{p}(0,T;H^{-1,q_{0}}(\mathbb{R}^{2}))\quad\mathbf{P}-a.s.$
-Then, again by maximal $L^{p}$-regularity techniques for the heat equation (see e.g. [53, Theorem 4.4.4] [37, Theorems 10.2.25 and 10.3.4]) and the trace embedding of [53, Theorem 3.4.8],
+Then, again by maximal $L^{p}$-regularity techniques for the heat equation (see e.g. [55, Theorem 4.4.4] [39, Theorems 10.2.25 and 10.3.4]) and the trace embedding of [55, Theorem 3.4.8],
 ${\omega}^{*}_{0}\in C([0,T];L^{q_{0}}(\mathbb{R}^{2}))\cap L^{p}(0,T;H^{1,q_{0}}(\mathbb{R}^{2})).$
@@
 (4.4) $\Delta(\phi_{0}v_{0})=\nabla^{\perp}\omega_{0}\phi_{0}+\Delta\phi_{0}v_{0}+2\nabla\phi_{0}\cdot\nabla v_{0},\quad(\phi_{0}v)|_{\partial B(x_{0},\rho+\varepsilon/4)}=0.$
-From the regularity of $\omega_{0}$, by standard elliptic regularity theory (see for example [61, Chapter 4]), it follows that $\phi_{0}v_{0}\in C([t_{1}-\varepsilon/4,t_{2}+\varepsilon/4];H^{1,q_{0}}(B(x_{0},\rho+\varepsilon/4);\mathbb{R}^{2}))\cap L^{p}(t_{1}-\varepsilon/4,t_{2}+\varepsilon/4;H^{2,q_{0}}(B(x_{0},\rho+\varepsilon/4);\mathbb{R}^{2}))\ \mathbf{P}-a.s$.
+From the regularity of $\omega_{0}$, by standard elliptic regularity theory (see for example [64, Chapter 4]), it follows that $\phi_{0}v_{0}\in C([t_{1}-\varepsilon/4,t_{2}+\varepsilon/4];H^{1,q_{0}}(B(x_{0},\rho+\varepsilon/4);\mathbb{R}^{2}))\cap L^{p}(t_{1}-\varepsilon/4,t_{2}+\varepsilon/4;H^{2,q_{0}}(B(x_{0},\rho+\varepsilon/4);\mathbb{R}^{2}))\ \mathbf{P}-a.s$.
 Therefore, since $\phi_{0}\equiv 1$ on $B(x_{0},\rho+\varepsilon/8)$
@@
 $-\psi_{0}w_{l-1}\cdot\nabla\left(\omega_{w}+\sum_{j=0}^{l-2}\omega_{j}\right)-\psi_{0}\left(w_{j}+\sum_{j=0}^{l-2}v_{j}\right)\cdot\nabla\omega_{l-1}.$
-Due to 4.3 and the inductive hypothesis
+Due to Lemma 4.1 and the inductive hypothesis
 $h_{l}\in L^{p}(0,T;H^{-1,q_{l}}(\mathbb{R}^{2}))\quad\mathbf{P}-a.s.$
@@
 Let 1.1, $r\in(2,4)$ and $p\geq 2^{N}\frac{r}{r-2}$.
-Let $(v_{1},\dots,v_{N-1},\overline{v})$ be the $(p,r)$-solution of (3.7) in the sense of 3.5.
+Let $(v_{0},\dots,v_{N-1},\overline{v})$ be the $(p,r)$-solution of (3.13) in the sense of Definition 3.7.
 Then, for all $0<t_{1}\leq t_{2}<T,$ $x_{0}\in\mathcal{O}$, $\rho>0$ such that $\mathrm{dist}({B(x_{0},\rho)},\partial\mathcal{O})>0$,
@@
 $\tilde{v}$ $=w+\sum_{j=0}^{N-1}v_{j},\ \widetilde{\omega}=\operatorname{curl}\widetilde{v}.$
-As described in 4.2, arguing as in the proof of [29, Theorem 7], we can extend the weak formulation satisfied by $\overline{v}$ to time-dependent test functions $\phi\in C^{1}([0,T];H)\cap C([0,T];\mathsf{D}(A))$ obtaining that for each $t\in[0,T]$
+As described in Lemma 3.5, arguing as in the proof of [31, Theorem 7], we can extend the weak formulation satisfied by $\overline{v}$ to time-dependent test functions $\phi\in C^{1}([0,T];H)\cap C([0,T];\mathsf{D}(A))$ obtaining that for each $t\in[0,T]$
 $\langle\overline{v}(t),\phi(t)\rangle-\langle u_{\mathrm{in}},\phi(0)\rangle$ $=\int_{0}^{t}\langle\overline{v}(s),\partial_{s}\phi(s)\rangle\,\mathrm{d}s-\int_{0}^{t}\left\langle\overline{v}\left(s\right),A\phi(s)\right\rangle\,\mathrm{d}s$
@@
 $-\psi\overline{v}\cdot\nabla\widetilde{\omega}-\psi v_{N-1}\cdot\nabla\widetilde{\omega}.$
-Due to 4.3 and 4.4 the terms
+Due to Lemma 4.1 and Lemma 4.2 the terms
 $\widetilde{v}\cdot\nabla\psi\omega-\psi\left(\widetilde{v}-v_{N-1}\right)\cdot\nabla\omega_{N-1}-\psi\overline{v}\cdot\nabla\widetilde{\omega}$
 $-\psi v_{N-1}\cdot\nabla\widetilde{\omega}\in L^{2}((0,T)\times\mathbb{R}^{2})\quad\mathbf{P}-a.s.$
-Therefore $h\in L^{2}(0,T;H^{-1}(\mathbb{R}^{2}))+L^{1}(0,T;L^{2}(\mathbb{R}^{2}))\ \mathbf{P}-a.s.$ Then, arguing as in the first step of the proof of [41, Theorem 13.2], the fact that ${\omega}^{*}$ is a distributional solution of (4.13) implies that ${\omega}^{*}\in C([0,T];L^{2}(\mathbb{R}^{2}))\cap L^{2}(0,T;H^{1}(\mathbb{R}^{2})).$ Therefore
+Therefore $h\in L^{2}(0,T;H^{-1}(\mathbb{R}^{2}))+L^{1}(0,T;L^{2}(\mathbb{R}^{2}))\ \mathbf{P}-a.s.$ Then, arguing as in the first step of the proof of [43, Theorem 13.2], the fact that ${\omega}^{*}$ is a distributional solution of (4.7) implies that ${\omega}^{*}\in C([0,T];L^{2}(\mathbb{R}^{2}))\cap L^{2}(0,T;H^{1}(\mathbb{R}^{2})).$ Therefore
 $\omega\in$ $C([t_{1}-\varepsilon/4,t_{2}+\varepsilon/4],L^{2}(B(x_{0},\rho+\varepsilon/4)))$
@@
 $-\widehat{\psi}\left(\widetilde{v}-v_{N-1}\right)\cdot\nabla\omega_{N-1}-\widehat{\psi}\overline{v}\cdot\nabla\widetilde{\omega}-\widehat{\psi}v_{N-1}\cdot\nabla\widetilde{\omega}.$
-By 4.3, 4.4 and relation (4.2) it follows that
+By Lemma 4.1, Lemma 4.2 and relation (4.2) it follows that
 $\widehat{h}\in L^{2}(0,T;H^{-1/2}(\mathbb{R}^{2}))\ \mathbf{P}-a.s.$
@@
 $\cap L^{2}({t}_{1}-{\varepsilon}/{64},{t}_{2}+{\varepsilon}/{64},H^{5/2}({B(x_{0},{\rho}+{\varepsilon}/{64})};\mathbb{R}^{2}))\quad\mathbf{P}-a.s.$
-This concludes the proof of 4.5. ∎
+This concludes the proof of Lemma 4.3. ∎
 Corollary 4.4.
 Let 1.1, $r\in(2,4)$ and $p\geq 2^{N}\frac{r}{r-2}$.
-Let $(v_{1},\dots,v_{N-1},\overline{v})$ be the $(p,r)$-solution of (3.7) in the sense of 3.5.
+Let $(v_{0},\dots,v_{N-1},\overline{v})$ be the $(p,r)$-solution of (3.13) in the sense of Definition 3.7.
 Then, for all $0<t_{1}\leq t_{2}<T,$ $x_{0}\in\mathcal{O}$, $\rho>0$ such that $\mathrm{dist}({B(x_{0},\rho)},\partial\mathcal{O})>0$,
@@
 $\omega_{w}=\operatorname{curl}w$ $\in C([t_{1}-2\varepsilon,t_{2}+2\varepsilon],C^{\infty}({B(x_{0},\rho+2\varepsilon)}))\quad\mathbf{P}-a.s.$
-and $\omega^{*}=\omega\psi\in L^{2}((0,T)\times\mathbb{R}^{2})$ supported in $[t_{1}-\varepsilon,t_{2}+\varepsilon]\times B(x_{0},\rho+\varepsilon)$, then, arguing as in the proof of 4.5, it follows that $\omega^{*}$ is a distributional solution in $(0,T)\times B(x_{0},\rho+\varepsilon)$ of
+and $\omega^{*}=\omega\psi\in L^{2}((0,T)\times\mathbb{R}^{2})$ supported in $[t_{1}-\varepsilon,t_{2}+\varepsilon]\times B(x_{0},\rho+\varepsilon)$, then, arguing as in the proof of Lemma 4.3, it follows that $\omega^{*}$ is a distributional solution in $(0,T)\times B(x_{0},\rho+\varepsilon)$ of
 (4.11) $\partial_{t}\omega^{*}$ $=\Delta\omega^{*}+\widetilde{h}$
@@
 From the regularity of $\omega,\ \overline{v},\ \widetilde{\omega},\ \widetilde{v},\ \omega_{N-1},v_{N-1}$, then $\widetilde{h}\in L^{2}(t_{1}-\varepsilon,t_{2}+\varepsilon;H^{-1}(B(x_{0},\rho+\varepsilon)))\ \mathbf{P}-a.s$.
-By standard regularity theory for the heat equation, see for example Step 2 in [41, Theorem 13.1], a solution of (4.17) with $\widetilde{h}\in L^{2}(t_{1}-\varepsilon,t_{2}+\varepsilon;H^{k-1}(B(x_{0},\rho+\varepsilon))),\ k\in\mathbb{N}$, belongs to $C([{t}_{1}-\varepsilon/2,{t}_{2}+\varepsilon/2];H^{k}(B(x_{0},{\rho}+\varepsilon/2)))\cap L^{2}({t}_{1}-\varepsilon/2,{t}_{2}+\varepsilon/2;H^{k+1}(B(x_{0},{\rho}+\varepsilon/2)))$.
+By standard regularity theory for the heat equation, see for example Step 2 in [43, Theorem 13.1], a solution of (4.11) with $\widetilde{h}\in L^{2}(t_{1}-\varepsilon,t_{2}+\varepsilon;H^{k-1}(B(x_{0},\rho+\varepsilon))),\ k\in\mathbb{N}$, belongs to $C([{t}_{1}-\varepsilon/2,{t}_{2}+\varepsilon/2];H^{k}(B(x_{0},{\rho}+\varepsilon/2)))\cap L^{2}({t}_{1}-\varepsilon/2,{t}_{2}+\varepsilon/2;H^{k+1}(B(x_{0},{\rho}+\varepsilon/2)))$.
 Therefore
@@
 $\cap L^{2}({t}_{1}-{\varepsilon}/{16},{t}_{2}+{\varepsilon}/{16};H^{2}(B(x_{0},{\rho}+{\varepsilon}/{16})))\quad\mathbf{P}-a.s.$
-Reiterating the argument as in Step 3 in [41, Theorem 13.1] the claim follows. ∎
+Reiterating the argument as in Step 3 in [43, Theorem 13.1] the claim follows. ∎
 Proof of Theorem 1.3(2).
-The claim follows by 4.3, 4.4 and 4.6 and a localization argument.
-To begin, recall from the proof of Theorem 1.3(1) in subsection 3.3 that there exists a solution (1.2) on the time interval $[0,T+1]$ and it is given by $\widetilde{u}=w_{g}+\sum_{i=0}^{N-1}v_{i}+\overline{v}$ where $(v_{0},\dots,v_{N-1},\overline{v})$ is the $(p,r)$-solution to (3.7) on $[0,T+1]$ for $r<2q_{\mathcal{H}}$, $N$ as in (3.5) and $p\geq 2^{N}\frac{r}{r-2}$.
-Then, by 4.3, 4.4, 4.6 and a standard covering argument, for all $t_{0}\in(0,T)$, $\mathcal{O}_{0}\subset\mathcal{O}$ such that $\mathrm{dist}({\mathcal{O}_{0}},\partial\mathcal{O})>0$,
+The claim follows by Lemma 4.1, Lemma 4.2 and Corollary 4.4 and a localization argument.
+To begin, recall from the proof of Theorem 1.3(1) in Section 3.3 that there exists a solution (1.2) on the time interval $[0,T+1]$ and it is given by $\widetilde{u}=w_{g}+\sum_{i=0}^{N-1}v_{i}+\overline{v}$ where $(v_{0},\dots,v_{N-1},\overline{v})$ is the $(p,r)$-solution to (3.13) on $[0,T+1]$ for $r<2q_{\mathcal{H}}$, $N$ as in (3.11) and $p\geq 2^{N}\frac{r}{r-2}$.
+Then, by Lemma 4.1, Lemma 4.2, Corollary 4.4 and a standard covering argument, for all $t_{0}\in(0,T)$, $\mathcal{O}_{0}\subset\mathcal{O}$ such that $\mathrm{dist}({\mathcal{O}_{0}},\partial\mathcal{O})>0$,
 (4.13) $\widetilde{u}\in C([t_{0},T];C^{\infty}(\mathcal{O}_{0};\mathbb{R}^{2}))\quad\mathbf{P}-a.s.$